Hall C 12 GeV A1n/d2n Meeting Minutes

Do Not Forget List
Study Material
Analysis Milestones
(Collab mtg)
(Collab mtg)
(Collab mtg)
05/29/2018 05/22/2018 05/15/2018 05/08/2018 04/10/2018 03/19/2018
02/15/2018 02/06/2018 01/30/2018 01/18/2018
12/12/2017 12/08/2017 10/17/2017 08/03/2017

  • List of things we should not forget
  • Add NJL model (A. Thomas) to plots, see Cloet et al paper.
  • Enlist theorists' help to extract neutron from 3He data; Enlist theorists' help to extract delta q/q from A1n (JAM, J. Qiu, ).
  • Some discussions on theory, see theory/
  • Beam scraping on glass wall: (discussion with Kent P. on 3/8/2020): "beam position difference at the injector without tuning is at the 1micron level. With relativistic boost the size will be smaller at the target, but beam spot may get squashed in one direction vs. the other which would increase the size in the wider direction. Overall, 1micron difference at target is a safe estimate. Also, noise itself is 6-8 micron for each Gaussian peak (helicity plus vs minus), so the h+ peak is only slightly shifted from the h- peak. And no, IHWP flip does not flip this shift since some effects are flipped and some are not." -- Conclusion: we can estimate helicity dependence in the glass scraping by taking 1micron over 4mm (size of raster, since the scraping "peak" is no higher than the main peak), which hopefully is small: 1/4000=0.00025 vs. stat error of measured asymmetry 0.003-0.004 in the most precise x bin.
  • Use Bosted code for positron background calculation, see Simona's email thread from 8/17/2020. Also useful are F2 and D. Flay's 6 GeV data on the positron background. (Note: Xiaochao has a fix for Peter's pi0 code.)
  • Moller: divide all measurements into three periods (before/after Wien flip and after spot move), fit all measurements in each period with a straight line and quote +/-3% uncertainty on the average value will be good enough (Dave G. 3/9/20)
  • PVDIS asymmetry: too late to flip target spin direction so will need to calculate the effect and correct the data (hey any expert in the collaboration?).
  • Non-cycled SHMS Q1,2,3, see SHMS Q1 study technote, and measured hysterisis loops; Dave M. said Q2 and Q3 are essentially EM magnets so not affected. Q1 has the biggest effect from non-cycling but the field difference is not visible on the Hall probe so at most +/-1G out of 5kG, or +/-0.02% based on field study.
  • Misfunctioning scaler module on HMS hclog3841229. Not sure when the problem started, could be for all A1n/d2n running up to the hclog entry date. For offline analysis, we will need to be careful when using HMS scalers. Fortunately the same scalers are copied to SHMS DAQ and can be checked using SHMS runs' data.
  • QE asymmetry data, see PLB 788, 10 January 2019, Pages 117-121 and Phys. Rev. Lett. 113, 232505; QE asymmetry calculation, need to email the 3 groups that did SAGDH calculations: Bochum group, Golak et al. nucl-th/0202064, nucl-th/0508017, Email: ufgolak@cyf-kr.edu.pl; Hanover group, Deltuva et al., nucl-th/0308045, https://arxiv.org/pdf/nucl-th/0406065.pdf, https://arxiv.org/pdf/nucl-th/0503012.pdf, Email: arnoldas.deltuva@tfai.vu.lt; INFN Roma Glvanni Salme PWIA, Email: giovanni.salme@roma1.infn.it
  • Wien flip LEFT vs. RIGHT happened twice: once during A1n and 2nd time durint the Covid shutdown. See hclog3786542 for some nice explanations by Joe Grames
  • Propose A1n SoLID rungroup proposal for a future PAC.

  • A1n/d2n Analysis and Monthly progress (Month of Jan 2022)
    1. Melanie working on Radcor and POLRAD2.0, and need to re-bin data in E' (not x) so each bin of HMS and SHMS shares the same kinematics (important for radiative corrections).
    2. Mingyu: N2 dilution is done, doing asymmetry double check, waiting for E' binning from Melanie, will also look into nuclear corrections.
    3. Mike making progress on polarized yield data/sim comparison, need to correct QE dilution. Note that one can either figure out QE fraction and use asymmetry to correct, or subtract QE spectrum and then form asymmetry. The two methods should lead to the same results.
    4. Now it's Mike and Mingyu's turn to understand the helicity decoder.
    5. Junhao will give a summary/update of d2n asymmetry analysis on Feb. 15th. Also need to make sure the binning can be combined with Murchhana's xsection work.
    6. Murchhana running RC external for d2n cross section radiative corrections.
    7. Bill pointed out there is a missing line of setting istop to nonzero value near HMS mc_single_arm line 466. This causes some events not passing the aperture cut to have istop=0.

  • A1n/d2n Analysis and Monthly progress (Month of December)
    1. Melanie started working on POLRAD2.0. Got old A1n code from 2001 and updating SF and asymmetry fits in the code now.
    2. Mingyu continued stability checks. Suggested defining "z value" to pick errorneous runs (outliers).
    3. Mike need to post similar stability check outputs for A1n/d2n based on his script. Need to make sure that the summary page includes the replay version used.
    4. Brad reminded us that we need to have a clear-defined runlist for all production and all "special" runs, like this https://hallaweb.jlab.org/wiki/index.php/List_of_special_runs_(d2n) for 6 GeV d2n.
    5. No meeting on 12/14 (Xiaochao final exam), and 12/28 (X'mas break)

  • A1n/d2n Analysis and Monthly progress (Month of November)
    1. We spent most of the month to understand helicity synchronization again. Looks like we are okay.
    2. Melanie spent quite some time to study the "bad/good" raster cuts and statistical analysis of whether the change in the asymmetry is significant. Also started looking into POLRAD running.
    3. Mingyu
      1. making progress on A1n N2 dilution, with x-binned results (but low-E' setting needs a lower x range where asymmetry data exist); JP commented that x-dependence may be present, especially near resonances. Brad commented the 2010 d2n analysis showed no kinematic dependence of N2 dilution.
      2. In N2 analysis, made yield stability plot. Suggest making this for all runs and to rule out "bad" runs.
      3. After N2 dilution is done, will revamp asymmetry code for an analysis cross check.
    4. Eventually, goal on A1n also include our own nuclear corrections.
    5. Mike making some progress on collimator punch-through simulation. Need to focus on getting the uncertainty caused by QE.
    6. Regarding target polarimetry, Junhao finished this for BigBrother, Dutch still not done. Mingyu is almost done with pulsed NMR analysis, still need to extract Ptarg for each run.

  • A1n/d2n Analysis and Monthly progress (Month of October)
    1. Melanie's scraping analysis: found PID cuts can remove most of scraping events. On the other hand, asymmetry of events with "bad raster cut" (cutting on the scraping side) is large, though not always negative. Do we have problem with helicity synchronization? -- still investigating.
    2. Mingyu working on an analysis notes on the window dilution, posted notes in elog108. Need everyone to read and comment.
    3. Next step for A1n asymmetries would be (with tentative name assigned) radiative corrections (Melanie), N2 dilution and positron background analysis (Mingyu) -- scratch positrons, we didn't take data for A1n, so just N2 dilution;
    4. Murchhana looking into characterization of acceptance uncertainties (continuing). Need someone to do COSY study (and who)?
    5. Mike got more QE calculations, starting to look into collimator "punch through events" to make the elastic simulation to agree better with simulation
    6. Junhao working on N2 dilution, latest is that he is looking into detailed helcity beam/software structure and synchrnozation, cutting beam trips, etc. While it is not necessary to use helicity scalers for cross section analysis, it's best to understand (in prinple the xsec results can be extracted from non-helicity-gated scalers, but the same results should also be obtained from gated scalers). Also, Junhao is in charge of d2 asymmetries, so better understand this fully.
    7. Jixie finished target field calculation and analysis for d2. He reported this on Oct. 19th and passed on many information to Murchhana for followup calculations.

  • A1n/d2n Analysis and Monthly progress (Month of September)
    1. A newer round of replay (see last minute) is now ready to use.
    2. Melanie is working on scraping. The idea is to treat it like an unpolarized glass window dilution. This has to be done run-by-run. Already did some cutting study but there is typically no significant shift in the extracted asymmetry given that our statistical uncertainty is so large (relatively to any systematic effect). Once window scraping is done, will move on to asymmetry radiative corrections.
    3. After we shifted method for the window dilution (see last minute), Mingyu has made some progress on this.
    4. Junhao still working on N2 dilution.
    5. Murchhana on d2n cross section: we focused on discussing how to get the uncertainty of the acceptance. Bill showed that for F2, one varies the field of all magnets, regenerate COSY field (both forward and backward) and rerun mc-single-arm. However, we think this method only gives us the relative stability of the acceptance (with is sub-%) but not absolute. Suggested contacting Abishek who showed DIS data analysis at the last Hall C meeting.
    6. Mike N. showed progress on elastic asymmetry, see elog112. The spectrum on page 4 looks odd, but otherwise good progress. Need to get polarized xsections.
    7. We received QE calculation from Scopetta group, see elog114.
    8. Bill posted preliminary beam polarization in elog21.

  • A1n/d2n Analysis and Monthly progress (Month of August)
    1. Melanie reached a good agreement on elastic asymmetry sign, see elog24. Also studied Delta transverse. Suggested to also get the sign of transverse asymmetry for elastic and QE (since the cuts are well developed). If the uncertainties are decent we might make use of these asymmetry data
    2. Melanie on asymmetry analysis: still need to cut out scraping and finalize uncertainties. Latest (no scraping cut) plots can be found in A1n elog21.
    3. Most recent window dilution by Mingyu was hitting a bump (see elog111: the simulated window yields do not agree with data for both production and reference empty cells. This disagreement seems to change with angle: data much below sim for 30 deg, about the same for 20, and data above sim for 13 deg. Mark suggested comparing focal plane distributions but after struggling with this comparison for elastic analysis, we decided to not use simulation to correct for empty target window spectra. Instead we will use RC external directly (take ratio of cross sections for empty ref / production) to correct for the different radiation length of the cells.
    4. Junhao had some results on d2n asymmetries (see d2n elog15, but now is focusing on getting N2 dilution. Last status was some weird yield fluctuation on pol3He cell yield, see d2n elog 16. Suggested to look at different factors (in particular LT) and also get a rough estimate of the dilution factor to see how the fluctuation affects the systematic uncertainties.
    5. Murchhana: latest had code working for xsection extraction (input 1.0, extracted 1.0, for example). Continue cross section extractions.
    6. Mike N. still doing elastic comparison. Latest status is cann't get good agreement on both target and fp variables (if a cut works for target, it makes fp comparison of sim vs. data worse, and vice versa). Suggest to move ahead to extract asymmetries. It's likely that whatever cuts we use the final results are quite similar.
    7. Bill posted preliminary beam polarization in elog21.
    8. We found some latest patch was not implemented in our 1c replay (this affects optics and window width in z). Brad will start another round of replay soon.

  • 07/13/2021 and 7/20/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie and Junhao continued working on respective asymmetry analysis
    2. Mingyu's window dilution analysis is going well. There is a gap in the simulated window spectrum that needs looked into. Suggest developing a method to determine systematic uncertainties, and then extend to x-binned window dilution analysis (is it needed though?). It appears Will is the only cell that does not have window thickness data but measuring it now (EEL or UVA) has radcon complications.
    3. Mike showed elastic/QE plot, it appears previous deficit in simulation was due to not identifying the correct paddles in the data vs. sim comparison.
    4. Murchhana's "mock" cross section extraction is still off by 7%. This is by setting the xsection to be 1.0. Suggested looking into the details. We really expect to see the extracted "xsection" to be 1.0 in this case.
    5. Jianping mentioned a lattice calculation talk on d2n at the EIC workshop, see here

  • 06/22/2021 through 7/6/2021 A1n/d2n Analysis and Weekly progress
    1. We had meetings every week but my notes are scarce. Here is a summary of what happened based on my memory:
    2. Mingyu is still working on RC_external caclulations. The previous problem with heavy element calculation was a problem in Fortran input file format which was fixed 3 weeks ago. In his latest post elog101, there is a nice summary table of what needs to be done (kinematic settings, target material settings) but the problem is that the table is less than 20\% done. Moving forward, I suggested:
      1. Choose one kinematics that had all RC_external grids completed, do window dilution analysis.
      2. Ask Junhao for help in setting up d2n input files (kine, target thickness).
      3. Ask Murchhana to determine window position from d2n data. These will go into input file of the simulation. She posted elog103 on the "target cell z position" result.
      4. Jixie will help to speed up the farm jobs/streamlining scripts.
    3. Melanie continued analyzing livetime and charge corrections. Last reports are posted at A1n elog14, A1n elog15, and LT correction in A1n elog16... A1n DIS LT corrections are now Okay, but not sure about the status of 1-pass data (EL-clean).
    4. Junhao presented d2n asymmetry analysis, suggested sorting out beam IHWP states and apply sign flip. Last status was one of the IHWP IN/OUT state seems to have coding problem (asymmetry value and uncertainty change visibly with Q and LT corrections) .
    5. Murchhana started extracting xsections from simulated files. We spotted some problems during our Friday discussions before July 4th. Latest status was for d2n 18-deg, extracted xsections are higher than calculation but are self-consistent (more or less, among the 4 choices of acceptance cuts). For A1n 30-deg the agreement is actually very good. This should give us a hint why the 18-deg setting does not agree well.
    6. Mike proceeded to polarized 3He elastic runs (8.5 deg SHMS). Elastic and QE agree with data fairly well but we had to add inelastic (just above pion threshold) from F1F209 of Christy et al. And even so the simulation is still below data on low E' side. Eric also warned that 3He fit is not very well constrained at low Q2. Suggested:
      1. Plot W and W(3He) to see if can achieve a better separation of inelastic vs. QE;
      2. Contact Chao Peng to see how SAGDH data were analyzed.
      3. Also showed 20-deg SHMS data (or 18 deg??) but do not seem to see elastic peak at all. Suggested adding EL and IN contributions.
      4. Start asymmetry analysis for el/QE.
    7. Golak's group calculation for more QE point is finished, see elog98 (this replaces the previous elog 96).

  • 06/15/2021 A1n/d2n Analysis and Weekly progress
    1. Mike reported (1) Looked more into E/p drift (data quality monitoring). Probably will need to plot calo raw ADC spectra to diagnose what's going on; (2) started comparing 3He ref cell data (all scintillator bar on) with Golak's QE and elastic simulations.
    2. Mingyu showed that (1) RC_external's cross section for heavier element (anything bigger than 3He) does not agree with xstree code. After the meeting we traced the problem to input file formatting -- the Z and A values are not read in correctly. Mingyu will rerun everything for the window element, Jixie and Bill will help him to automize things. (2) F1F2IN21 is ready in xstree but not int RC_external code. This has to wait for later.

  • 06/08/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie reported on livetime (elastic) and helicity-gated livetime (DIS) analysis. Something is not quite right: livetime for elastic or Delta 1-pass runs are higher than 100%. DIS helicity-gated livetime (A1n) fluctuate too much. Will ask Carlos to analyze a few runs and cross check.
    2. Junhao started working on asymmetry extraction, suggested following proposal binning.
    3. Mike posted run stability checks, see elog95. E/p value shows interesting drifts. Suggested plotting all runs in time order (use color to separate kinematic settings).
    4. Xiaochao is in communication with Golak group on calculating finer grids around the ppn breakup threshold, see elog98.
    5. We discussed upcoming UGM and Hall A/C meetings.

  • 06/01/2021 A1n/d2n Analysis and Weekly progress
    1. HUGS is ongoing for this and the next two weeks. We may need to adjust meeting time to see if it fits better students' schedule.
    2. Murchhana reported that the reason for the delta-dependence of the acceptance calculation was due to the "3-sigma" "good recon" cut. The sigma value of xp(recon)-xp(ini) for example has strong z dependence. If fitting z-depenent sigma then this is gone. However, given that there are less than 0.1% of events with recon outside of the 3-sigma range, it was decided that repeating the same exercise for the recon quality of other variables is not necessary and we can extract the acceptance without the 3-sigma cuts.
    3. Junhao reported summary of pressure curve analysis for all runs, see elog94. Nothing in too glorious agreement with the fill density expectation and results vary from cell to cell and from kinematics to kinematics. Will need Mingyu's RC xsec grid for radiative corrections.

  • 05/25/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie presented livetime analysis. Suggestions: (1) check eta definition (it should be 100% for low rate runs), and prescale used; (2) plot livetime asymmetries in addition to livetime; (3) for asymmetry plots need to plot error bars (to see how much the cut reduces the statistics), and better binned in x (or W for 1-pass runs); (4) BCM reading shows slight different charge from Chao's online replay (used the same params), need to check.
    2. Murchhana showed updated acceptance study, see elog90. The changes since last time are: now using recon not thrown variables, and added "good recon" cuts - for each variable, fit the residue (thrown-recon) with an Gaussian and apply a +/- 3sigma cut. The problem is the cut seems to have a high delta dependence. After the meeting we found out this is because the "sigma" value vary strongly with delta for two of the 4 variables (and weakly for the other 2).
    3. Mingyu continued working on generating RC_external tables for the GE180 glass, see elog89. The discontinuity in RC_external output was traced (on slac after the meeting) to how RESMODD in F1F2IN09 is used for nuclear target and it cuts off at Q2=11 GeV2. We asked Eric Christy but was told the F1F2IN21 isn't optimized for 3He yet. Since Mingyu already generated tables for all other material, we will continue using F1F2IN09 and will contact Eric later (end of summer) for an update.
    4. Mike reported checking Golak table's smoothness and test out the intepolation algorithm for "in between" kinematics (using both his own and David F.'s code), see elog91. The interpolation looks very good. For "smoothness", it may be necessary to request for 0.1-deg steps (currently 0.25 deg) near the QE edge but we will wait another week before going back to Golak. Suggested: add the "below ppn" points which aren't so many. Then will proceed with last week's to-do list.
    5. In-between meeting notes: Xiaochao posted from XEM elog a bump in xpfp reconstruction, see elog84. Though looks like if we apply an abs(delta).lt.8 cut then it should be automatically rejected.

  • 05/18/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie presented livetime calculation updates, see elog86. Looks good (for A1n DIS very low rate) but will check the calculation with the helicity cuts and runs with higher rates. Cautioned that both EDTM pulser and beam ramp rate changed during the run.
    2. Mike showed plots for elastic 3He with Golak's QE xsections implemented. Comparison with data looks good for runs where all scintillator bars were on. Will: (1) make plots of xsec vs. theta for fixed Eprime and see if finer theta binning is necessary (finer than 0.25-deg calculation will have to go back as a request to the Golak group); (2) proceed to runs with custom configurations of scintillator bars but will start with ref cell runs with known pressure, then to runs of elastic polarized cell. (3) Also need to code the asymmetries. There were some nice chats and followup code snipits about how to do "4-point interpolation" (colinear? coplanar? Binomial? what's that jargon again for the double linear intepolation?)
    3. Junhao showed more pressure curve analysis for Brianna and Austin. Will (1) check N2 subraction in detail as DIS runs do not produce as good results as 1-pass runs. N2 correction plays a much bigger roles for DIS runs; (2) We noticed Dec'19 Brianna HMS runs show nice elastic peak. Might want to consider using those for physics analysis (later, for Mike).
    4. Mingyu showed more RC external xsec grid outputs. One of the settings show plotting glitch that needs checked. Still running RC-external for the glass windows because we need 7 grids and then take the mass-average (that's for the 7 elements of the glass composition).

  • 05/11/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie finished pion PID and contamination analysis for d2n, summarized all (both A1n and d2n) in a nice slide, see elog78 (the details are on slide 26 of this PDF or 2nd attachment of the elog entry). Previously we talked about dividing the data (where spectrometer momentum is very close to NGC threshold) below and above NGC threshold but given the pion contamination is so low, Brad said that's not necessary. Same is probably true for the NGC efficiency "dip" other experiments are seeing. Next will focus on livetime calculation with and w/o beam cuts. So far w/o cut results agree with report file.
    2. Murchhana showed acceptance study for SHMS 18-deg and 30-deg and HMS 20-deg and 30-deg, see elog79. Look all reasonable. SHMS delta is asymmetric but that could be just how the magnets bend the particles. For A1n 30 deg settings may want to reproduce theta/phi plots for |vz|.lt.17 or 18cm just so to make sure the cuts are still fine for the smaller vz range. Mingyu or Melanie will use these plots to determine acceptance cuts in the asymmetry analysis.
    3. Mingyu showed window thickness measurements with more points, see elog82. Because the previous 3-point measurement and the new one do not place on the same curve (the 2nd point is 10% thicker than target center) though they are consistent within error bars, the new average up-stream window thickness (now 1.014) is less than previous value (1.08). Suggest using the curve for now and use the variation for uncertainty study. Target length variation has been found negligible.

  • 05/04/2021 A1n/d2n Analysis and Weekly progress
    1. Replay is done but we need to check if any variables are missing.
    2. Mingyu showed RC external calculations for 6 GeV d2n kinematics, see elog75. Look reasonable. Also showed target cell window thickness measurement using a cut-out from Todd. Measured only two points, will measure more before returning it to Todd as it can be useful for the upcoming GEn-2 running.
    3. Melanie continued working on livetime corrections. A relevant elog is at elog74 but we did not discuss as it's not final.

  • 04/27/2021 A1n/d2n Analysis and Weekly progress
    1. Replay is almost done!!!
    2. Melanie: posted skim script at elog73. Now everyone can develop own analysis script!
    3. Melanie also pointed out the hodo ADCTDCtimediff on SHMS are different among channels. Mark said previously this was not important but with hcana's recent update it's safer to align them to zero. Fortunately it was already done prior to the latest replay. Phew!
    4. Mingyu:
      1. showed a few plots of RC_external's output cross sections. Elastic and QE tails are negligible. Questions remain why xsec_RC is lower than xsec_Born. We decided Mingyu should double-check the code with established data (hint hint: David Flay's thesis).
      2. showed calculation of window thickness with raster on. The difference is at 1E-4 level when the beam is centered. Even for 1.5mm of off-centered beam the effect is tiny.
      3. Regarding window non-uniformity, Chris does not have this info. Todd commented he has a window for GEn now and can do a multi-point measurement. (Similar questions were raised for GEn as well.)
    5. Murchhaha
      1. showed: for SHMS with rectangular (xptar,yptar) cuts, wide vz cut and plotted acceptance vs. delta. Results look promising because the 3 tightest cuts show nearly flat delta acceptance. Jianping suggested using forward-MC (with some model like F1F2) and extract the cross section using these acceptance grids. Any dependence on the (xptar,yptar) cut used can be treated as the relative uncertainty of the extracted acceptance as well as the method to extract xsections. Mark reminded that when doing so, we need to apply bin-centering corretions.
      2. Still working on HMS and 30-deg settings (rootfiles disappeared over the weekend and have to be regenerated)
    6. Mike: concluded the carbon excited states are quite important. We made some suggestions for how to improve the plot. One thing we noted is the MC elastic peak seems to be offset from data. Need to plot W(12C) or equivalently excitation energy (not delta, which has the angle in it) and see if the offset is consistent with the expected knowledge of Ebeam and Eprime (both are at 5E-4 level??).
    7. Jixie reported on making the acceptance code working with multi-thread method.
    8. Post meeting chat: we exchanged "2nd shot experiences"...

  • 04/20/2021 A1n/d2n Analysis and Weekly progress
    1. Murchhana reported on acceptance study (round 1, no cut), see elog70. Suggest:
      1. For round 2, apply 2D theta vs. phi cut (probably multiple), delta between -10% and 10% (for HMS) and an equally wide range for SHMS, and make all acceptance 1D delta plots.
      2. Generate round 1 plots for 30 deg MC files. Need to pay attention that the z cut should be very wide (perhaps +/-24cm is good).
      3. Need to automize and run all these on the farm.
    2. Melanie showed the Calo E/p plot for d2n settings. Observed a 3-4% shift from A1n. May need to rerun the calibration, see elog72.
    3. Mingyu showed the first round of RC_external xsec calculation, see elog71. There were questions on whether xsec_RC is lower than xsec_Born, so will check/plot the elastic or QE tails. Also need to consider that with the raster and/or the beam offset from (0,0), the upstream glass window is thicker than the window thickness itself.

  • 04/13/2021 A1n/d2n Analysis and Weekly progress
    1. Melanie reported checking ECal calibration for d2n data. Some settings not consistent with A1n observation and need re-calibrated.
    2. Melanie also reported PID analysis for d2n. The problem is that at least two of the d2n setting had the SHMS central momentum close to the Cherenkov threshold pion momentum, and it's best to divide the data into "above" and "below" threshold. The exact value of the threshold cut (supposedly using delta) should be determined from data. For -7.5 GeV/c runs, no need to apply NGC cut because the momentum is too much higher than the threshold the NGC does not contribute to PID at all.
    3. Junhao reported on pressure study using 3He reference cells, see elog65. General suggestions are that in addition to look into pressure gauge accuracy and calibration, should look into the low-prescale (high deadtime) ref N2 run 11474 as it will be a main source of the systematic uncertainty of this study.
    4. Friday 4/16 discussion: focused on dry runs for APS April meeting. Plus some discussions on the acceptance and RC_external usage.

  • 04/06/2021 A1n/d2n Analysis and Weekly progress
    1. Mike on elastic/QE: raised the possiblity that 12C excited states may be the reason for simulation missing data in the el/QE gap region. Will ask CSR for excited state information.
    2. General discussions: kinematic binning for window dilution will be "naturally applied" when analyzing data.
    3. Friday 4/9 discussion: Junhao went over asymmetry analysis roadmap; We discussed RC_external, and Jixie showed how to submit (old system) farm jobs as we will need to generate xsec tables.

  • 03/30/2021 A1n/d2n Analysis and Weekly progress
    1. Junhao presented pressure curve of 3He elastic using September SHMS data, see . Plenty of discussions and to-do's:
      1. Overall looks like already reached 5%. Need to improve to reach 2% (this will go into xsec uncertainties)
      2. can try cutting on W(14N) to reject N2 elastic, but there will still be N2 QE contamination. A better way may be to use N2 ref cell data and scale down using N2 fill pressure of the production cell and see how well it works.
      3. Can't ignore the vacuum run and the 382Torr ("vacuum") run. These should locate directly on the line if the analysis is done correctly.
      4. Currently the line does not cross zero (or the empty target point). It could be that there is window contamination even with the tight z cut. Can try to use the vacucum run to subtract this "vacuum window contamination".
      5. z cut seems to be too much on the positive side where the acceptance is dropping. SHould shift z but to more symmetry and avoid the sharply-dropping region (use 3He pol cell run to judge).
      6. need to add error bar to horizontal axis: about +/- 1 psi for pressure for each run.
      7. If everything is done correctly, the "3He pressure" vs. cut should be a flat line rather than continously decreasing. -> this should be the goal.
      8. Instead of using elastic peak, another method is to integrate the full elastic+QE spectra. This should work provided all pressure curve runs were taken under exact same conditions. Can plot yield from all runs and compare.
      9. Someone needs to help checking the TC pressure equation -> will leave this to target meeting.
    2. Murchhana showed plots on the simulated acceptance using 30-deg MC files. We discussed the following:
      1. MC used 6mm-dia raster, will cut on 4.5mm when extracting the acceptance;
      2. Do we expect 100% acceptance for some region? -- maybe not. These spectrometers are not like HRS where accp~100% for a wide range.
      3. Currently we only have 30-deg MC files made by Jixie. Jixie will show Murchhana how to submit jobs and she will run high-stat MC for all d2n angles.
      4. plan to add the acceptance event-by-event using the skim script -- nobody seems to object it so we will do this.
      5. After the meeting, Mark and Jixie discussed whether we need to add xtar (or vertical position of raster) to the acceptance and make it 5D. The answer currently is no, because acceptance for a certain (xp,yp,ztar,delta) bin should be quite fixed and not vary much within that bin (for different raster position). Murchhana could run a setting of MC with the raster off and see how big the difference is in the acceptance output.
      6. Discussed how to "verify" the acceptance using data: (JP) normalize data by sigma_Mott and the result should be F1,F2 which should be smooth functions of (x,Q2). But we will need a skim script for doing this so the script can be used on pass1 replayed data.
    3. General discussions: To add radiated xsec to MC, we decided to use rc_external code, see https://github.com/JeffersonLab/rc-externals. Dave Gaskell commented (last Friday) that this is the state-of-art code for radiated xsections. David Flay looked at the code briefly and confirmed the code does what it's supposed to do and is a good one to use. We may need to run this code and generate huge tables like F2 did (since running this event-by-event will make the MC too slow).
    4. Discussions delayed to next week: (1) Do we need window dilution analysis for d2n? (2) focal plane mismatch -> later; (3) PID progress for d2n?
    5. Friday 4/2 discussions: Bill showed us how to run RC_external, see recorded video here. We will need to add pol3He target info first (Mingyu).

  • 03/23/2021 A1n/d2n Analysis and Weekly progress
    1. Murchhaha showed plots on the focal plane comparison of SHMS MC vs. data, see elog61. There is a corner missing for the downstream window in the xfp vs. yfp plot and is reflected in other variables too. This is not visible in A1n data (A1n doesn't see downstream window at all). This is more obvious at the SHMS -7.5GeV/c setting.
    2. Mingyu showed preliminary window dilution analysis using simulated windows only, see elog58. Overall good results. Next step is to move onto using empty target data as the first-order correction and use MC to correct at the second order. To do so, MC will need to incorporate radiated cross sections.
    3. We spent the rest of the meeting to discuss preparation for the collaboration meeting.

  • 03/16/2021 A1n/d2n Analysis and Weekly progress
    1. Our collaboration meeting is next Thursday. Please register here and the event page with agenda and Bluejeans connection information is at event/434. We discussed theory groups to contact for future talks/work.
    2. Presentation on radiative corrections by David Flay, see elog55 for slides and bluejean recording. One possible project remains to do is to convert POLRAD to C++. Will check with Sylvester if a C++ package already exists. If not, this can be a possible task for students.
    3. Mark looked into MC over the week and found a few things:
      1. Related to the suppression of HMS MC in phi angle, see elog51. This is a one-word fix in mc-single-arm. The shms side was fixed long time ago but hms was only fixed last week.
      2. Related to the HMS optics problem in Junhao's report two weeks ago, see elog52, basically zreact needs to be calculated twice. This needs modification in hcana (or replay) and thus followup from Sylvester+Brad. THis effect is larger at small angles. As for short target (4-10cm), depends on whether a cut is applied to z.
      3. When reading through (our version of) the code, had the following comments: It is probably a minor point but it looks like you are still including the scattering chamber window when calculating the multiple scattering. You should probably remove the scattering chamber window and put more air (N2?) between the cell and the HMS or SHMS entrance. Also it looks like you are using the cyrotarget design for multiple scattering instead of the actual target with the glass side windows. ANyway these are minor points, but probably worth fixing. -> need followup from the sim crew.
      4. (this was by email): suggest to not cut too tight on theta or phi. Suggest abs(theta).le.0.08 and abs(phi).le.0.06 for SHMS, as an example. Wider cuts are needed for extended targets.
      5. Emphasized our knowledge of P0 is at 0.2% level for HMS and SHMS. Knowledge on SHMS acceptance is still not well understood especially for our long target.
      6. In addition to the two updates from last week, talked about the importance of correcting "DC offsets", see elog60. This correction seems to have improved Murchhana's d2n simulation/data comparison greatly for the windows. Also suggested looking into focal plane checks (x,y,xp,yp). Our data are hitting the DC edge. The survey of DC chamber edge should be checked with data.
    4. Murchhana presented the d2n simulation update. The 3 fixes improved a lot the agreement between MC and data on the windows, see elog57. Next step is to compare focal plane distribution between MC and data, develop acceptance cuts, and to generate acceptance files using Jixie's high-stat simulation outputs.
    5. Mingyu is working in the target lab today so missed the meeting. He will report on the window dilution study next week.
    6. Junhao presented progress on N2 pressure curve, see elog54. Details below:
      1. N2 elastic peak is very small for HMS 1-pass runs (again, SHMS runs without the target collimator are useless and we did not take N2 pressure curve runs during d2n 1-pass running when the collimator was in). Will try to get the uncertainty in the N2 elastic peak and then wrap up this topic.
      2. Will move on to N2 ref. runs for DIS. For 3He pressure runs, will check both DIS and 1-pass d2n runs on the SHMS with the target collimator.
    7. Other updates:
      1. Survey results for book-keeping see elog53
      2. Todd sent us the Reference cell window info, see elog59
      3. Bill pointed us to Aruni's talk on focal plane study: f2_optics.pdf

  • 03/09/2021 A1n/d2n Analysis and Weekly progress
    1. Today's meeting will be short because of the SoLID DOE Science Review.
    2. Brad: replay is still ongoing, mostly checking all updates and working with Sylvester on including the helcity scalers.
    3. Melanie: also still checking on the skim script.
    4. Mike on elastic/QE:
      1. used DC residue to add smearing: hdc_sigma increased from 0.03 to 0.06 (in cm) to match z width in data for the central foil. Junhao mentioned he did not update the residue params after DC calibration -> Mike will work with Junhao on this.
      2. need to check what gas is in NGC in mc_single_arm, as the 600um seems large.
      3. So far did not use radiated cross sections. - will need to radiate the cross section - consult David Flay?
    5. Jixie showed last friday how to use XSTree code to extract acceptance from simulated rootfiles. We need very high statistics. jixie will hand over a script to Brad (or vice versa) for running the simulation on farm. But need to wait until some mc-single-arm fixes are done, see Mark's item below.
    6. Murchhana showed the first simulation for d2n runs, see elog. Discussions below:
      1. Data windows are much lower than simulation. Not sure why. Mark mentioned we are still understanding SHMS (but the windows for d2n are at smaller y so should be better understood than A1n??)
      2. We need the window thickness data for Tommy. Mingyu will write to Chris.
      3. Please check if the cuts on theta and phi are switched (there appear to be a cut on theta at 0.05 on slide 20, while the texts above the plot shows 0.06). Then, make sure the cuts do not cut into data or MC so we have a fair comparison.
    7. Mingyu found a mistake in calculating the number density for GE180. After correcting this the agreement of data vs. MC is much better for the windows for A1n, see 2nd PDF on elog49. Otherwise still working on the window dilution analysis.
    8. Junhao worked on compiling all special runs, otherwise still working on the pressure curve analysis.

  • 03/02/2021 A1n/d2n Analysis and Weekly progress
    1. Brad: replay is done for about 1/3 of runs, but encounters some farm problem. Still sorting it out. Suggested posting location of completed jobs on wiki or elog.
    2. Melanie presented a first draft of the skim script. see elog47. Details:
      1. Beam trip cuts can be customized using 3 variables (examples): min current (4uA) and time before (10s) and time after (5s - 10s in elog). For the given example values, the events between 10s before the current drops below 4uA, and 5s after the current rises above 4uA, will be cut out and assigned a "false" flag.
      2. The helicity-gated scaler group (Carlos' work) is not implemented yet (same is true for pass1 replay). Currently using "standard scalers". Brad asked if standard scalers are gated and if so, should be able to cut out the beam trip too. Melanie will look into this.
    3. Junhao presented first round of work for the N2 pressure curve, see elog46. Details:
      1. Surprisingly, the N2 elastic peak is clearly visible in the HMS run (no target collimator) from Dec 2019. Even got the first N2 pressure curve! Some suggestions below
      2. check optics: z_react should has no dependence on any of the other variables;
      3. after checking optics, see if the dp vs. th correlation remain. Note that th is the "out of plane" or "vertical" angle should have no physical correlation with dp for elastic or QE.
      4. tighten z and see if can get rid of that N2 elastic peak for the empty ref cell run;
      5. if can get rid of the N2 elastic peak from empty cell data, choose that z cut and repeat for all N2 runs; redo the curve;
      6. before we give up on getting a "precision" result from HMS runs, suggest give it a try to separate the N2 elastic peak from 3He peak for the 3He run. We might get a large uncertainty due to overlapping 3He and N2 elastic peaks, as expected.
      7. After all above is done, repeat the analysis for September d2n 1-pass N2 pressure curve runs and a 3He run from the same period. The z cut may not be necessary because of the collimators, but needs check.
    4. Brad reminded us the special run list should be posted at https://hallcweb.jlab.org/wiki/index.php/Pol_He-3_Analysis_Resources#Special_Runs. Xiaochao posted the A1n list there, only thing missing might be the positron runs, can't recall if we did the positron run for d2n or A1n.
    5. Mingyu presented latest simulation for A1n, see elog49. Comments and observations:
      1. The non-physical events within the empty vacuum ref cell run were rejected by PID cuts, not ibcm1 cut, so likely due to the good track requirement in etracknorm;
      2. changing beam position in MC does not cause a visible effect. The empty target HMS phi MC is still narrower than data. Looking at this carefully, could be the acceptance effect: the windows might get cut in theta or phi as they are far away from z=0. Perhaps applying acceptance cuts will help.
      3. Updated MC to BigBrother window thickness: Upstream is 138um which is thicker than the previous 125um. Also moved the cell by 1cm upstream (survey of TOmmy shows it was 1cm downstream, so 1cm off is not atypical, still waiting for BB survey info). Now the window height of data still above MC (both SHMS upstread and HMS both windows). Could be nonuniform glass thickness? Wall scraping makes the beam scraping the (thicker) edge of the window -- not sure.
      4. Discussed the next step on A1n MC: Since we have not decided if cross section analysis is necessary, should wrap up MC work with the goal of extracting and finalizing the glass window cut and the window dilution factor.
      5. Another possible output of MC is the acceptance. But for asymmetry extraction, some "eyeballing" might be good enough, i.e. just choose region where acceptance is reasonable (not dropping sharply).
    6. Mike presented an update on 12C foil analysis: added QFS and found it working better than F1F2QE; data is wider than MC; and data elastic peak is lower and wider than MC. Will: (1) add smearing; (2) check if all radiation effects are counted for in MC for elastic; then (3) move on to 3He.
    7. SoLID Science Review is next Mon-Tue. We will have a short meeting on Tuesday 3/9.
    8. We had planned to hold the first collaboration meeting on 3/17 or 3/18. Show time, everyone!

  • 02/23/2021 A1n/d2n Analysis and Weekly progress
    1. Mingyu presented latest simulation for A1n. (Note that slides shown last week had a scaling factor applied to match data to simulation at z=0, and we spent the week to figure out the agreement after removing scaling, including realizing the sim spectra were missing the cut istop==0). See elog45. They still match (without scaling)!! Comments and observations:
      1. The abs(z_react).lt.30 cut removed the hour-glass shape in the raster replay plot (these are non-physical events);
      2. checked and EDTM events are not plotted (as desired). Also adding ibcm1.gt.1uA cut to clean up data following Bill's suggestion last week.
      3. The BigBrother data windows still look quite upstream by a few mm (both HMS and SHMS data), though the deficit seen last week in the SHMS z=(10,20)cm region is gone. No likely no problem in optics and sim matches data well there now!
      4. HMS data shows downstream window but simulation do not. -- not sure how to explain this but we can just cut it out?
      5. For empty target, last week we saw "grassy backgroun" for abs(z).lt.20cm which can't exist for vaccum. After applying abs(z).lt.30cm, ibcm1.gt.1uA, and PID cuts, now these events are gone. It's not clear which cut cleaned up these events though -- likely ibcm1, but PID cuts also possible if it requires a clean track (as expected for cutting on etottracknorm).
      6. For empty target data, the windows match well simulation (indicating ref cell is mounted at exactly where it should be), though the winow height is a bit off. Also the sim phi distribution is narrower than data. We thought maybe we didn't illuminate the simulation enough but this difference remain after Mingyu widened the initial phi range this week. Thus we need to look further whether it is coming from, for example, beam positions.
      7. At this precision, Mingyu will do: (1) study which cut removed the vaccum events, PID or ibcm1? (2) implement beam position to see if it improves phi agreement for empty target; (3) implement exact window thickness for BigBrother and ref cell -- though keep in mind if we see scraping then the beam is not hitting the center or thinnest part of the window, so we have some uncertainty here; (4) Could consider now mispointing and shifting the cell position, but we are not at this stage yet.
    2. With simulation at this precision, we started to argue whether it's worthwhile to also extract large-x cross section data for A1n -- stay tuned!
    3. Junhao will discuss with Xiaochao on how to start the pressure curve analysis after the meeting. -- done, we will see how N2 peak looks next week!
    4. Melanie emailed an update that she has requested merging of the helicity-gated scalers.
    5. Murchhaha emailed an update that she is returning to the US this Friday.
    6. Brad emailed an update that he is going to miss the meeting and will be back next week.

  • 02/16/2021 A1n/d2n Analysis and Weekly progress
    1. Brad: still working on the kinematic file for replay launch (HMS is completed but SHMS still needs work, some varialbes such as Ebeam and target mass are totally incorrect). Will work with students to update this using run lists.
    2. Melanie: working on the "skim" script. Beam trips are done and the cut is reliable, still need to add helicities.
    3. Mike on SHMS simulation: added cherenkov flag
      1. the carbon foil ztar is still wider than simulation.(We thought the foil is thick at the meeting but then I recall this is small angle scattering so it's possible the foil looks thick in z);
      2. Data shows events beyond z=20 and rising, could be secondary events from the collimators get reconstructed at random places. This is very hard to simulate (so maybe find a way to get rid of the effect);
      3. fp-y distribution: MC seems to have a "flat top", should check what fp-y=0 events show up in other variables and maybe they get cut off some where?
      4. More suggestions: for elastic spectra, plot W(12C) to check if the elastic peak is at the right place. Also always plot the two foils' simulation in two colors; Do the same comparison study for HMS (to have more tools to help with spotting problems); Then, if we don't see any major problem, next step is to figure out extra smearing (hard to do - DC smearing? optics recon smearing? unknown smearing??).
    4. Mingyu on DIS simulation comparison with data, see elog 42. Some comments and suggestions:
      1. For replayed data need to reject non-physics events (such as EDTM) and but a cut on current above zero (Bill mentioned the replay may put zero-current events at some fixed/or random places). This should (we hope) explain the raster y vx. x "fan shape" events, and the events in the middle of the empty cell run (vacuum? dark matter?).
      2. Note: all pol3he data are from cell BigBrother, some tiny bit of scraping is seen.
      3. Initial theta and phi ranges are a bit too narrow (simulation is narrower than data for SHMS theta, HMS theta, and a bit on SHMS phi).
      4. Need to add electron cut. Something like (E/p.gt.0.8).and.(E/p.lt.2.0).and.(Npe of Cherenkov.gt.1) would be good.
      5. For raster x and y, as well as focal plane x,y,xp,yp, plot 1D histogram in addition to 2D.
      6. Target windows are not at the same ztar as simulation, could be a combination of target mount (couple of mm?), spectrometer mispointing (not yet corrected in replay).
      7. Suspect the acceptance is not well understood for positive phi side (z=+10 to +20cm at 30 deg), if this remains we will ask Mark for advice.
      8. After the meeting we realized some normalization problems that need corrected.
      9. Some fine-tuning for the long term: raster shape, window position fine-tuning, resolution smearing...
      10. Ugh! This is a set-back: a scaling factor was applied in all slides to match data at z=0 to MC. Need to remove this factor and attempt absolution comparisons.
    5. Junhao showed some slides to discuss pressure curve analysis. May be hard to do because of the poorer resolution of Hall C spectrometers. Simulation to correct for radiative effects can be for later.

  • 02/09/2021 A1n/d2n Analysis and Weekly progress
    1. Brad: worked on the kinematic file for replay launch;
    2. Xiaochao: Received setting 3 calculation from the Golak group. This concludes the calculations we expect from this group (their model is non relativistic and won't work for setting 4 the largest angle). Also received an initialization email from the second group that they are dusting off their code now.
    3. Mike: added the target collimator and is using all available data to "fine tune" the collimator position.
    4. Bill: Recall the SHMS resolution problem shown in Hall C meeting (Abishek's talk)? Turns out that was solved by turning the flag of the 2nd GC "on" in the simulation, see MC_SHMS_Cher_flag_slide.pdf.
    5. Mingyu: made some progress on simulation. May need to shift the window position of the pol3He cell a bit along z, and add N2 simulation for the gas outside the cell.
    6. separate meeting on 2/11 am: Radcor - maybe ask David Flay for the tool, for both asymmetry and xsections; d2n simulation: okay for Murchhana to learn how to calculate yield since it helps to gauge if the simulation is correct (even if we eventually use the acceptance method to extract xsec for d2n).
    7. separate meeting on 2/11 pm: We will send pseudo data to Wally, and have invited him for our March collaboration meeting. Need to know theoretical uncertainty and its dependence on "how long hte theory extraction will take" (varies from 5 minutes of a crude extraction using Pn,Pp model to "never" for the ideal extraction -- "ideal" meaning "exact exact exact" in the eye of theorists).

  • 02/02/2021 A1n/d2n Analysis
    1. Drift chamber calibration: Junhao completed the 2nd round of DC calibration following the hodoscope calib completed last week, see elog41. Some channels has wider residue than others but the consensus is that's normal (Mark, Brad) . Checkings for some settings are still ongoing but the params are ready to go.
    2. Replay (Brad): will do a "sanity" check of DC params and then launch the full replay.
    3. Data analysis script: Sylvester will work with Melanie on this. The output of the code should be N+ and N-'s binned in (x,Q2) -- similar to online, with PID cuts for electrons and ytar (or ztar) cuts for windows; beam trip cuts; "decent" cuts on the solid angle and delta should also be applied (can have two sets: loser acceptance cuts for A1n, tigher ones for d2n, etc); and the W greater than 2GeV cut. All cuts can be exported as a flag (i.e. 0=not passing, 1=passing). Once we understand better the beam scraping, some beam position cuts should also be applied.
    4. Mike's elastic analysis: MC/data normalization looks pretty good now. There are a few comments and suggestions:
      1. XSTree code used Bosted's fit. From Jianping: For QE, Bosted's works well for carbon and other heavy targets, but not so well for 3He. Eric is still working on improving the 3He fit.
      2. theta (OOP) angle looks symmetric in MC but not in data. Suggest cutting tightly on delta: (+/-8%) for HMS, (-10%,22%) for SHMS.
      3. check MC error bars, they look underestimated;
      4. for HMS, add N2 gas simulation for the background;
      5. for SHMS, (1) remove the "collimator" flag in mc_single_arm input (that's the spectrometer collimator, not target collimator); (2) Add target collimator to mc_single_arm (Jianping will provide a correspondant); (3) Add smearing to MC to mimic data (this smearing is due to tracking algorithm and data indicates it is not a Gaussian, but can start with Gaussian for simplicity); -- Shujie can help with both (2) and (3); (4) add 3He elastic; (5) modify XSTree to add spin asymmetries.
    5. Scratch notes of useful points from various Hall C User's meeting talks last week:
      1. F2 talk (Bill): e+ background, Q2 had 1.8% tune to match MC with data on the focal plane (note this was for early SHMS experiments before the magnet tunes were fully understood)
      2. NPD talk: used ESR (known as VM2000 reflector) as reflective material.
      3. EMC talk (Abishek): (1) 12C data ztar width MC is wider than data (mostly due to optics tune, but Dave Mack argues some of the non-Gaussian spread is due to tracking algorithm and can be improved); (2) Some normalization problem that can be checked by re-measuring 12C foil thickness (now, since these foils don't explode).

  • 01/26/2021 A1n/d2n Analysis
    1. Hodoscope: Finally found the bug in HMS Vp code: the Vp fitting histogram range is too small and the data are out of range. This caused the low stat problem (see last two months' minutes) and the failed fitting repeatedly. After extending the upper range from 50 to 100 (thanks to Mark for finding it out), now Vp calib code works perfectly fine. Mingyu redid the calibration. Also checked SHMS to make sure the histogram range is fine. Final hodoscope parameters are posted on elog38. Suggestion to Carlos: add a 2D plot and ask the user to check the histogram range prior to fitting.
    2. Calorimeter: Melanie emailed this morning that the calibration is final. She is now updating the PID results using the updated Calo calib in preparation for the Hall C talk this Thursday.
    3. Since hodo calib is now final, we asked Junhao to rerun the DC calibration. Three settings: HMS, SHMS A1/d2n, and SHMS d2n*. If 3-4 hours per setting the DC calib may be completed within 24 hours.
    4. Replay goal: launch the replay jobs to farm before the Hall C meeting -- Brad.
    5. Monitoring scripts: may need fine tuning but probably more efficient if we test the code with the replay; Physics script: Melanie will look into this after PID work and the Hall C meeting; helicity code: Carlos will ask Steve to accept the git version, maybe during the Hall C meeting.
    6. Mike's carbon QE work: MC/data seems to off by factor 10, will continue working on it;
      1. Shujie offer that Hall A 3H/3He experiment had new results/fits on 3He QE -- Jianping commented it's highly Q2 dependent so the Q2 min of 0.6 of the 3H experiment may not be good enough for us;
      2. The model used is F(y)-scaling (or y-scaling?), again Jianping commented that it may work okay for heavy targets (like 12C) but not for 3He where few-body is important, also the Hall A work won't provide us the asymmetry.
      3. Nevertheless, since we start with carbon QE comparison as a benchmark, suggest Mike contact Shujie to get 12C QE calcuations. For kinematic grid see QE elog25 (bottom of file Golak_comm).
    7. Xiaochao received setting 2 calcuations from Golak group. See same entry as before: elog25.
    8. Discussed upcoming Hall C meeting talks:
      1. Went through Junhao's d2n slides;
      2. Mingyu's talk slides aren't ready (put all time into hodo calibration this week), suggest using some cool pictures from this this google drive directory;
      3. Melanie's slides will be DNP slides with updated PID results so probably okay;
      4. Bill's Moller talk was informative -- take away message is we can do day-by-day corrections to Moller results based on Q.E. and beam energy (electron spin precession varies by 1.75 deg for every 1MeV change in energy, and depends also on where the energy shift comes from - NL, SL, or both;
      5. Mike: no slide yet(?)
      6. Suggest all presenters to cross-reference each other's talk and acknowledge all RCs on the slide (since this is the first a1n/d2n talks at the Hall C meeting).
    9. Other misc items: Xiaochao talked through how to calculate yield from mc_single_arm with Mingyu and Murchhana.

  • 01/19/2021 A1n/d2n Analysis
    1. Hodoscope: still not done, Vp calibration not working due to low stat. Could tune the beta to 1.0 by adjusting the LCoeff in Vp param but we should avoid such tweaking method.
    2. Calorimeter: Melanie is doing final check. Goal is to find outlier blocks (param in principle should all be constant for all blocks after online gain match) and tweak by hand.
    3. Some analyzer/replay readiness discussion
    4. Mike showed some preliminary comparison of carbon foil data. Suggestion: (1) Run the QE option of mc_single_arm + xsections (not the elastic mode because the data is mostly QE at this kinematics); (2) In addition to comparing target variables, also compare focal plane ones.

  • 01/12/2021 A1n/d2n Analysis
    1. Focus on replay readiness (notes by Brad)
    2. Hall C analyzer (hcana) repo status -- Sylvester
      1. hcana code readiness for pass1 replay
        • Sylvester is working on integrating upstream hcana changes
        • moved hel_scalers_*.json files to REPORT_OUTPUT/
      2. discussed how an upstream change in how FADC hit selection is handled will impact us
        • decided to merge that patch (best to stay close to upstream) and make needed updates to our config files
          • we will need to establish 'default pedestals' (run new script)
          • change introduces a timing shift in the adc_tdc-tdc hit selection formula
            • need to add offset(s) in tdc hit times for the hodoscope
          • Sylvester + Mark will work together to work out optimal changes
      3. Sylvester will run some tests and push out a new singularity container
    3. Parameter file status:
      • Hodoscope calibrations -- Mingyu
        • We now understand the shift in SHMS beta noted earlier
          • due to shift in s2y plane that correlates to replacement of a failed scaler module (https://logbooks.jlab.org/entry/3846102) that shares the same ribbon cables with the associated TDCs between runs 11372 and 11376. Best guess is that the termination in the replacement scaler wasn't matched and this resulted in signal distortion and the observed ~2ns s2y timing shift
          • Brad will look at signals and follow up for future experiments
          • Solution for us is to go with Mingyu's updated calibration for runs
        • Issue with the shoulder on HMS beta in -7.5 GeV/c kinematic remains
          • Recalibration using a different run reduces the shoulder somewhat, but HMS beta peak is still non-Gaussian. Indicates an underlying problem with the HMS calibration.
          • The beta peak with the 'old' hodoscope calibration is narrower and more symmetric, but is only slightly offset from 1. Probably a better calibration than the new one, but this is a solvable issue.
          • Mingyu will plot beta vs. paddle number with the new calibration, and take a look at how the calibration constants changed between the 'old' config and the new one. We'll see a report on this next week.
      • Calorimeter calibrations -- Melanie/Sylvester
        • Melanie reports that the single 'defocussed' calorimeter calibration is looking optimal for all runs (as hoped). She is doing a few final checks, but it is looking good
        • Pion cross calibration results look consistent with the defocussed calibration results as well. Also a good sign.
      • NOTE: We'll need to double check the PARAM settings with the updated hcana and ensure all is well. (Also verify/address the expected timing shift noted earlier.)
    4. Confirm what replay script will be used for 'pass1' replay
      • SCRIPTS/hcreplay “-m all” ... is probably a bit too verbose on the output
      • Brad and Sylvester will pick/make a more balanced output.def list
        • Mark suggested we pack as many “QA” check/monitoring histograms into the replay output as possible.
        • Brad, Sylvester will followup with Mike N. to see what can be added
    5. Per-run Monitoring script readiness/update -- Mike N.
      • Couldn't make today's meeting, will follow-up with Brad, Sylvester offline.
    6. Simulation comparison with data, adding raster, formula checks? -- Murchhana, Mingyu
      • found an angle sign inconsistency in the cross section code: it used Hall A's code to transform from HCS to Spectrometer CS, which treats Left/Right HRS as positive/negative angles, but when calling this code, Jixie's code gave positive angles for both SHMS and HMS (basically it used the wrong angle for the HMS). This can be easily corrected by always input negative angle setting for the HMS when calling Jixie's x-section code.
      • There was a follow up discussion on how 'our' code differed from the Hall C maintained upstream code. It sounds like Murchhana and Mingyu are working off a fork of the upstream repo that Jixie had modified but were not sure what or how significant the changes were.
        • Murchhana and Mingyu will push documented git repos up to Github or ANL's GitLab (ie. with clear commit descriptions AND a clear ReadMe describing how to run the code and generate the output we need.
        • Murchanna and Mingyu will email the repo links to the analysis mailing list
        • Mike N and Sylvester will see if we can get the important changes merged upstream
    7. Elastic/QE simulation update? -- Mike N.
      • No report/update this week.
    8. Miscellaneous items:
      • New a1n_d2n_analysis@jlab.org analysis-focussed mailing list created: Subscribe if you aren't already on the list.
      • Junhao emailed that he is working on AFP loss analysis and will report next week.
      • Students sent the requested param files vs run-list to Brad. In most cases we have a common run-set that will work with all runs
        • Hodoscope calibrations still have some issues (plans to address noted above)
        • Need to confirm that Cerenkov cuts are consistent/optimal with Melanie's work
        • Once the hodoscope calibration issues are understood AND the upstream hcana changes are merged we will need to do one (hopefully) last sanity check and we are ready to go for the pass1 replay

  • 01/05/2021 A1n/d2n Analysis
    1. We first focused on the full-replay readiness:
    2. Hodoscope calibration status:
      1. Mingyu modified the calib code so if there are not enough events (less than 1000 for SHMS or 200 for HMS) for a good fit of the drift speed, then set to 15.0 cm/ns for that paddle/channel
      2. A "shoulder" in beta remains for HMS, could plot beta vs. focal-plane x to check if that's related to a specific paddel - to do for Mingyu. Bill mentioned there is also a "good sc paddle" branch that can be used to choose hits at middle of the paddle only.
      3. The summer 7.5-GeV setting problem (shift in beta) remain. Comparison of calib params shows c1 varied quite a bit from A1n to the d2n* (7.5GeV summer) calibration but c1 is not even used in hcana (only c2 is). Cearly we need a focus group on this problem now -- Mingyu, Carlos, Xiaochao, Mark
    3. Calorimeter calibration status - Brad, Sylvester: some blocks stil have zero or even negative params that are not reasonable. To recap:
      1. Melanie has done several iteration of calibration, first using the defocused run and then tried "merging" with params from DIS runs. The change in dE/E is negligible, see 07/28/2020; Brad argued that the merging method may not be reliable and so we should use the defocused run results for all runs.
      2. But the defocused run params may still have the "too strong correlation" problem. Sylvester/Brad will check the params and see if any block is signfiicantly different from others (in principle, all blocks are gain-matched which means all params should be the same, see minutes from 2020/11/10.)
      3. Sylvester will check with Melanie on the status of the pion peak check. This will be useful as supplemental info for the param check above.
    4. Brad asked all students to send him the params and the run number ranges for each set of param.
    5. Murchhana showed formula checking between mc_single_arm, hcana, and her own derivation using the optics notes (or HCS, TCS definitions), see elog33. So far, mc_single_arm and hcana, and the optics notes are all consistent though Murchhana's own derivation is off by a minus sign. In a followup discussion after the meeting, we found if we use Jixie's code to calculate the x-section and the rates, the input angle for HMS side should be negative (same as for hcana and the optics note). We will re-run HMS simulation and update next week. (60 years after the era of Enrico Fermi and the separation of "physicists" into experimentalists and theorists, we can say this is where both groups still share the same struggle: spend hours or days to chase a minus sign, or multiple minus sign errors here and there.)
    6. Mike N. provided a short update on elastic/QE simulations. Continued working on the energy loss. Also tried integrating the QE results from elog25 and tried a spline fit for continuous coverage in E'. In fact Xiaochao has just updated this entry to include a finer-bin version from Golak group dated 1/1/2021, for 9.25-deg and 4-MeV E' bins. The E' values are different from the previous "0.5-deg theta and 10-MeV E' bins", so technically one could combine both coarse-bin and the new results for a better input to the spline.

  • 12/22/2020 A1n/d2n Analysis
    1. Last meeting of 2020! Does anyone recall when we had our last meeting of 2019? We had it on 12/30 (I think) in the counting house!
    2. APS/Hall C talk plan: Hall C: A1n-Melanie, d2n-Junhao, target-Mingyu, elastic-Mike; APS: A1n-Melanie, d2n-Murchhana, target-Junhao, elastic-Mike. We will review abstracts on Jan. 5th.
    3. Carlos' notes on helicity scaler decoding can be found at doc1105, including a sample script.
    4. Replay script (hcswif): see minutes from 12/18 (last Friday). Bill found an error in the script and will fix it after consulting Steve. Both Melanie and Junhao tried the script and managed to set the parameter using run number ranges, see Junhao's brief instructions under folder 20201222/
    5. Replay plan -- Brad will check all calibration/params and launch the first round over X'mas. Will focus on cooking partial set of data per each setting and check output.
    6. Replay scripts: for cutting beam trips, it appears there are now 3 ways:
      1. First is Sanghwa's script that is already in the Hall C replay calib (Bill uses it for F2 currently), see instructions;
      2. Second is to use Carlos' new helicity scaler code. The beam trip cut can be done within the same pass (with a 1/120-sec delay, Carlos already have this cut in the sample script);
      3. Melanie was also working on something with Sylvester, but only half-finished.
      4. Melanie will look into all 3 and come up with a plan.
    7. Replay script regarding physics extraction -- later.
    8. Hodoscope mystery: Brad checked the hodoscope calibration and some paddles on the edges have incorrect drift speed results due to poor fitting (or low stat). Mark suggested Mingyu to determine certain criteria and to set those paddles ot 15.0 if fitting fails. Then plot beta again to see if it fixes the 7.5-GeV run problem. - We may meet tomorrow or next week to follow up on this.
    9. Melanie on pion peaks in calorimeter blocks: now added the "single block" cut. Most channels are fine but some are not. For SHMS shower blocks, Mark suggested to require "MIP peak in corresponding preshower block in front of that shower block" which may help to cleanup the spectra. For HMS, suggest requring "MIP peaks in all 4 blocks that line up" to clean up. Will try to fit the peak, fine tune the param, and plot dE/E to see if it improves or stays the same. (We will only need to revisit this if the tuned params give worse dE/E).
    10. Melanie on pion contamination: Looking okay/good enoughf or now. Will document so will not forget the details when writing thesis.
    11. Simulation formula checking: Mingyu is still working on it (comparing optics notes, hallc_replay, and mc_single_arm). Bill added that we should also check Jixie's code where the physics xsections are calculated.

  • 12/18/2020 Discussion Session
    1. Bluejean recording link: part 1 is on farm replay, part 2 on how to define different parameters for specific run number range - unfinished, part 3 on simulation;
    2. Bill's slides, and a couple of other stuff can be found under 20201218/.
    3. This should supercede the previous recording on 11/20.
    4. Example input files on A1n setting: shms_B.inp, shms_C.inp, test3.inp, test4.inp
    5. To do on farm replay: 1) learn how to use hcswif to submit jobs; 2) find out how to set params for specific run number ranges (on ifarm); 3) submit list of short (50k) replay jobs to farm, 1-2 runs per kinematic setting and verify outputs, verify params used.
    6. To do on simulation: 1) continue formula check; 2) learn how to write an input file; 3) run mc_single_arm then getXsec and plot rate vs. ztar.
    7. Followup on setting runnumber ranges: looks like we were on the right track on this screen. To change cut params and calib params, replace pointers g_ctp_parm_filename and g_ctp_calib_filename by the appropirate files. Also see brief instructions under 20201218/.

  • 12/15/2020 A1n/d2n Analysis
    1. Sylvester sent an email that the latest HMS optics have been added to hallc replay. Those doing replay should download the latest version.
    2. Carlos: working on the BCM asymmetry error calculation in the scaler code. Brad will check Carlos' notes on the helicity scaler. We would also like to see a detailed description of where the uncertainty of BCM comes from and how it's propagated.
    3. Melanie on pion peaks in calorimeter (see elog31): wrote code based on Vardan's but we are not sure if the peaks we see are real physics (pion) peaks, or a cut-out of the low-energy background or even pedestals. Need to add the parts from Vardan's code that select events that have no energy deposit in all other blocks, and remake all plots. This may take some time, and we decided to "do as much as we can this week, but can proceed to the first round of replay without it."
    4. Melanie on pion contamination (see elog30): Added "method3".
      1. Now the values for HMS are 0.1% (method 1), 0.44% (method 2), and 0.03% (method 3 by scaling down the pion spectrum). Method 3 assumes the pion sample in the final analysis has the same shape as the "clean pion" sample and this adds to uncertainty. Note: previously we thought the 0.1% is an underestimation because the pions in our analysis sample may be subject to a different pion rejection from the "clean" pions used in the rejection analysis. But now looking at these numbers (that 0.03 is way below 0.1), it is equally possible that the "real" pion rejection is higher than the value we are using (due to "clean" pion samples "not really clean").
      2. Overall we think the 0.1% from method 1 is good to a factor of 2, i.e. the real value may be between 0.05% and 0.2% which is good enough.
      3. For SHMS though, the 3 numbers are 0.2%, 0.4%, and 0.03%. The 0.2% is on the high side and we asked Melanie to refit the pion contribution (by tuning the functional form), since the red fit has a clear overestimate of the actual data near E/p=0.8 and also above 1.2, for both HMS and SHMS.
    5. Mingyu on hodoscope:
      1. Bill pointed out hclog 3818947 in which Simona noted that "starting with runs 10815 in SHMS and 3633 in HMS", "the timing changed so you will have to revisit your timing cuts". This solves the mystery why all summer runs have different diffTime for hodoscope paddle for both HMS/SHMS, and now we have the exact run numbers!
      2. The beta shift for -7.5GeV momentum runs is still now mystery: the -7.5GeV d2n runs from March do not show the shift, only -7.5GeV runs from the summer do. This means the shift is not caused by physics (and not by the difftime shift since other momentum settings do not show it). While we do not use the beta spectrum for analysis, the shift is worrisome as it may indicate deeper problems in the replay. Mingyu will send a couple of rootfiles to Brad for checking.
    6. We discussed replay script readiness. Bill will work with Carlos on getting the scripts ready and hand them over to students by Friday.
    7. Murchahana and Mingyu showed some mc_single simulation comparison with data. For A1n 30 deg, the SHMS downstream 3He simulation clearly shows a faster drop than data, see mc_shms_30deg_screenshot.png. For HMS the ztar seems reversed. To do: (1) add circular raster; (2) check ztar calculation. Here is a breakdown:
      1. check ztar calculation in hallc_replay in e.g. analyze.C, pay attention to whether the spectrometer angle is defined with a sign.
      2. Derive how ztar should be calculated following the definitions of various coordinate systems. A good reference is Carlo's thesis, for example on page1.png. You should derive equations that look like page3.png (note: I am not saying these Eqs.() are incorrect, but we do need to check everything with the sign convention of our choice).
      3. Similar to Carlo's thesis that shows how SHMS coordinates are defined, this HMS optics note should tell you how HMS coordinate systems are defined. Also see Holly's slides on HMS optics
      4. Another reference is the Hall A technote 02-012 on HRS optics.
      5. check ztar calculation in the latest mc_single code e.g. mc_single_arm.f, note the mc_single input defines spectrometer angle to be positive for both SHMS and HMS.
      6. Put everything together and make sure we have a consistent (i.e. correct) picture.
    8. Mike N. showed slides on mc_single, see elog32, with a focus on the formula used for multiple scattering and ionization loss. Jianping noted need to work on Bremsstrahlung too.

  • 12/08/2020 A1n/d2n Analysis
    1. Holly: merged SHMS optics to hallc replay as "default". (Sylvester confirmed the merge by email). Note: this does not affect simulation since mc_single is based on COSY for both forward and backward transports. Still not clear of HMS optics have been merged (Mark wasn't at the meeting). With this progress in optics (SHMS only) we can now start doing a couple of things:
      1. Mike N. can start comparing simulation with data for elastic scattering, with a focus on focal plane comparison
      2. Murchhana and Mingyu, both completed simulation of 3He gas and glass windows separately, can now replay some data runs and try a first comparison of simulation vs. data on the pol3He cell production data. (Murchhana will learn how to use the DIS cross section code first).
    2. Carlos: finished helicity decoding, Steve is checking the code before merging. His section of the presentation was recorded in bluejean, see recording link (must use jlab login to access). Highlights:
      1. Outputs of his code is the same as report files (Steve's version)
      2. We discussed how to cut on physics events.
      3. Working on a document to describe what the code does.
      4. Bill mentioned there is an existing script to add varaibles (such as BCM or current) to replay output event tree "T", see bcm_check_module.pdf. Can adopt this for our helicity and current cut variables.
    3. Melanie: Still trying to make Vardan's script work for plotting pions in the calo blocks -- will move this to higher priority for the coming week. Meanwhile, making good progress on pion contamination. We reviewed the slides (see elog29 and made a few suggestions:
      1. For method 1 (applying PS cuts to Etot/P spectrum, fitting pion and electrons separately, divide pion fit/electron fit by NGC rejection factor): when dividing NGC rejection factor, should use the low-momentum-setting rejection factor results for both low- and high-momentum settings. This is because NGC performance should be independent of the kinematics. Low-P setting has more pions so probably cleaner pion samples (and more accurate results on the rejection factor).
      2. The value given by method 1 is probably an underestimation. This is because the actual pion event causing the contamination is different from the pion clean sample used in NGC PID analysis.
      3. For method 2 (applying PS cuts and NGC electron cut to Etot/P spectrum, try fitting pion and electron separately, divide pion fit/electron fit which gives directly pion contamination): the pion fit looks too high (either constant or rising at high Etot/p). If possible, use the same functional form as method #1 and scale it down to the green spectrum here. We can call this "method 3".
      4. Overall good agreements between the two methods and the range between method 1 and "mod method 2" should give us a good idea of the uncertainty on the final pion contamination
      5. All above is for HMS. Still working on HMS.
    4. Mingyu on hodoscope check, see elog28
      1. d2n runs from March 2020 have the same difftime range as A1n runs. So we should use two sets of difftime cuts: one for "spring" and one for "summer".
      2. A1n hodo calib works for all d2n settings except the 7.5 GeV setting. For 7.5 GeV, did another calibration that seems to work fine for all runs at this setting. So we could use two calibrations: one for "low-medium momentum" setting and one for 7.5-GeV setting. But consulted Mark, and he said the beta spectrum should not change with momentum. So we may still have some problem here.
      3. Will continue checking all runs for difftime and calibrated beta spectrum.
    5. Xiaochao will send her SAMC code to Mike N. (It's fortran!)
    6. Looks like we are on-track for launching the reply before the winter break. Will need some farm/batch scripts. Bill will look into it (looks like John Matter is the expert here) and tell us (MMMJ) how to use them. We could split up the runs to 4 sets so everyone can practice using farm.
    7. There are a couple of conferences with deadline in early Jan. We should discuss this by the end of year (if and when we get invitations) and get us organized on a full coverage of talks - APS April 2021 and DIS 2021.

  • 12/01/2020 A1n/d2n Analysis
    1. Burcu warned us that we have been using the same momentum shifts in our replay as J/Psi and CT. About momentum offset: this can only be determined from precision ep data. We will continue using what we have been using but keep this in mind when we do precision e-3He elastic analysis. (Not important for DIS). Mark commented this shift is typically at the 0.2-0.3% level for SHMS and ~0.1% of HMS if P0 is below 5 GeV.
    2. Optics: Holly commented she has been replaying data without the momentum offset. She is not optimizing delta so this offset should have little effect on the optics of SHMS she has been working. Tried xtar correction but so far not working (yet).
    3. Both Holly and Mark will send the optics files to Sylvester for merging.
    4. Carlos on helicity scaler coding: showed some screen shots of current work. One problem is the clock does nto seem to move forward. Another problem is the 1MHz clock rate does nto look exactly at 1MHz. Suggestions: (1) Is it possible the scaler event clocks are read differently and what is shown is actually the new reading every (1/120) seconds? In this case, one would see values different from 8243 or 8244 at the start and end of the run where a helicity cycle is cut short. (2) There is a "undefined" helicity time window every time the helicity is flipped so one should expect to see not 8333 (1E6/120) but ratehr some values smaller. That's maybe why we always see 8243 or 8244.
    5. Mike on elastic QE: looked into details of all Eloss and MSC needed and will implement these into mc_single. Will ask Chao and Mike Paolone for the laste SAMC code.
    6. Mingyu on hodoscope checking:
      1. Tried another d2n run but all timing shifts persists. But applying A1n calibration seems to work okay for these runs if applying very wide timing cuts. (Both HMS and SHMS). So maybe the only thing that needs changing is the timing cuts?
      2. When trying to update timing cuts for the newly fitted d2n peaks, ran into (possibly) formatting problems that result in some histograms being empty. This may be the reason why previouse attempt to rerun calibration using d2n runs failed (reported last week but did not write down in minutes - XZ). Will look into this formatting problem next.
      3. Junhao double checked the script and got the same results (shifting in all channels).
      4. Next, Mingyu will check d2n runs from March and see if we need to split d2n runs into Spring vs. Summer and using different difftime cuts.
    7. Murchhana showed some simulated spectra for the glass windows and 3He (no xs weighting). Will be good if we can compare these with real data with the latest optics implemented.
    8. Xiaochao received the first QE calculation from the Golak group for the 8.5 deg setting only, see elog25. These are for 0.5-deg theta and 10-MeV E' bins, 10 times coarser than our original request.
    9. Next week we will meet at 3pm.

  • 11/24/2020 A1n/d2n Analysis
    1. Optics: Mark HMS optics is ready, will email Sylvester for merging. Holly: will work on xtar correction for SHMS (expect 4 times smaller effect than HMS though).
    2. Melanie on calorimeter calib:
      1. showed pion peaks in SHMS preshower blocks, some blocks are shifted from the expected value of 80 MeV. Next will make plots for the shower blocks. Need to use cuts for single block only (so the pion event goes through the whole block and does not share energy with other blocks). (XZ) emailed Vardan for his scripts and received them within the week.
      2. But what do we do if we do see shifts? Discussion: 1) first we may want to calibrate all blocks using pion peaks (using Vardan's scripts); 2) then compare the two params for the pion and electron peaks in Etot to see if one is visibly better than the other. 3) In principle, the pion method should be better since it removes the strong correlation of params of adjacent blocks.
    3. Mingyu on hodoscope check of d2n:
      1. For unknown reasons, all d2n hodoscope panels show different timing (by 5 to 10ns) from A1n data. All difftime cuts need to be refitted for d2n. What is causing the shifts??
      2. To double-check, MIngyu will use a different d2n run. If the shifts remain then will study if we need a different hodo calib for d2n.
      3. Also to double check, Junhao will try to reproduce the plots using a d2n run.
    4. Pion peak check in calorimeter -- next time.
    5. Junhao: d2n DC calibration seems fine.
    6. Melanie on PID:
      1. After applying low calorimeter cuts on the pion samples, now the pion rejection for NGC is much higher. (where is the elog?). Also found that the "peacock feather" structure in the PSH vs. Etot plot is from binning effects.
      2. Realized the previous pion contamination wasn't calculated correctly (these were shown on page 39 and 47 in elog 24 slides). What needs to be done is to: Method 1: make one Etot/p spectrum with preshower electron analysis cut (blue), one spectrum with preshower electron cut and NGC pion sample cut (red). After that, apply Etot electron cut and calculate the ratio of integrated red to integrated blue (integrate above the Etot cut). THis would be the pion contamination without NGC rejection. Divide this ratio by NGC's rejection factor that corresponds to the NGC electron cut of our choice should give us the final contamination. Note: if the red spectrum clearly shows an electron peak then will need to use fits (sum of a quadratic function for pion and a Gaussian peak for electron) to separate electrons from pions and take only the pion fitted spectrum to integrate when calculating the pion contamination.
      3. Or using Method 2: make the same blue spectrum as in method 1 with preshower electron analysis cut, then make one spectrum with preshower analysis electron cut and NGC electron analysis cut (green). One should see the green spectrum is dominated by the electron peak but some pion background may remain. After that, do a two function fit (e+pi) to the green and see if can separate electrons from the pion background. If the green spectrum is too small at low Etot/p, may use the value at low Etot/p as an upper limit of pion level. Now, apply Etot electron cut and calculate the ratio of the two functions (pi over e) in the two-function fit of the green spectrum. This would be the pion contamination with NGC rejection (since NGC cut is already in the green spectrum).
      4. Compare method 1 vs. 2 to make sure the results are consistent with each other.
      5. Note: applying the NGC electron analysis cut will change the pion contamination slightly since the cut efficiency is not 100%. But this should be a small effect.
      6. For a graphical explanation see PID_Black_Friday.pdf
    7. Carlos reported having technical difficulties to create the helicity scaler tree. -> see more details the week of 12/1. Also Hall A does not have a class for writing helicity scalers.
    8. Mike Nycz will be doing the elastic/QE analysis! We discussed the difficulties of modifying mc_single for the elastic setting: 1) focal-plane matching (important for scintillator bar matching); 2) multiple scattering in the spectrometer: to do this precisely need cosy for magnetic transport of the MSC spread. THe best we can do without such COSY model is to use approximations.

  • 11/17/2020 A1n/d2n Analysis
    1. Optics: Based on last week's progress, tried 5th order and also more iterations but did not see improvement for the negative z foils. Will summarize next time.
    2. Calibration check following last week's merge:
      1. Murchhana on NGC check: all SHMS are good; all HMS ref time plots have less background than previoiusly, while all difftime and calibration histograms are the same as before. (Note: this was solved the week after. discovered a wrong rootfile was used.)
      2. Melanie on Calorimeter check: all difftime and calibration histograms look good. Pion peaks check -> need to wait for next time.
      3. Mingyu showed hodoscope compairson of old replay and old calib vs. new replay with new calib. Slight improvement for both HMS and SHMS. Hodo diff time is also good for A1n data set. Will email this to Brad for merging and proceed to d2n data check.
      4. Junhao: previous DC calib already used Mingyu's hodo calibration so should be good.
      5. Jianping is wondering if we can use beta from hodoscope for PID. However the 2m drift time of the hodoscope itself won't be long enough so we will need the 22m from the target. In this case we must use RF (beam bunch) time as the starting time and so far this is not in the hall c replay code. Resolution of scintillators is in the order of 300ps (Mark). WE may revisit this later for positron run analysis (where background is huge and we need to reject protons too).
    3. Melanie showed PID work, see elog24
      1. Showed calorimeter pion rejection which is around 4000+/-1500 (on page 5, to be more specific, the rejection is 27641/7 and error bar is in the order of 1/sqrt(7)). To do: Will look into higher statistics and maybe a better way to evaluate error bar. (Evaluation of the error in N1/N0 where N1 is a subset of N0 has always been a topic of discussion!); Other comments:
      2. Showed NGC pion rejection, still has problem because only applied "high cuts" on the calorimeter for the pion samples. There appears to be large background in the calorimeter spectrum that is near zero in both directions. These must be removed by using "low cuts" on the calorimeter for the pion samples.
      3. There is no need to apply window (z) cuts when doing PID.
      4. For evaluating pion contamination (for example in the Etot spectrum on page 39 and 47), got about 1% pion contaminationfor the low mementum setting with high PID cuts, electron eff is about 90% (low); We noted that need to separate real pions from the knock-on electrons (the spectrum shows small electron peak after the pion cut is applied). -- see update on 11/24 though, the 1% result is not correct.(in a good way!)
    4. We discussed briefly how to use simulation to extract cross sections (mostly for d2n). Bill mentioned F2 is using the simulation code to fully simulate the yield (with both internal and external radiation losses) and then compare with yield from data. Then, the experimental xs is (Data/MC)*(xs)_rad. We may not use the same method. (Jianping promotes the method of using simulation to get the acceptance only), correct the data yield by the acceptance to get the cross section.
    5. Along the same simulation topic, we noted for elastic or QE analysis, we need to fully rely on simulation for the itnegration of xs model and acceptance. To do so we also need to cut out region where the simulation does nto agree with data on the focal plane.
    6. Next step: Melanie will work on analysis script in preparation of the first cooking; Junhao will look into pressure curve, Murchhana into simulation (for learning), Bill will get the Moller reanalysis organized (now 3%, will be 2% after re-analysis, has a plan from Dav G. and may take 4 person * weeks); looking for a plan for elastic analysis.

  • 11/10/2020 A1n/d2n Analysis
    1. Optics: Mark reported trying more iteration on xtar-xptar fitting. For negative z, more iteration helped the optics at 18 deg. At 30 deg, it is also much better but the z=-20 foil is still not that perfect. Will try more with iteration and then the odd-order terms suggsted last week. Holly: will try the xtar fit this week for SHMS (meanwhile, latest SHMS optics can be found at doc1101).
    2. Murchhana on NGC calibration: HMS methods 1 vs. 2 and the difference is 1/8.4 (method 1) vs 1/6.2 (method 2), see HMS_CER_Calib_method1_d2n.pdf, HMS_CER_Calib_method2_d2n.pdf and SHMS_NGCER_calib_method2_d2n.pdf. Previously we thought the difference could have come from extra electronic noise that widens the peak, but a short check shows the change in gain (param) is the opposite of what we would expect from a noisy widening of the peak. Please note that the NGC spectra in the method 1 slides show very clear 2- and 3-pe peaks and their position seem to confirm the results from method1.
    3. Calibration organization and a forward flow plan - Brad: checked all history of 4 detector calibrations. All but hodoscope look fine and have been pushed to github. Hodoscope calibration looks was done based on a very old replay code and should be repeated. To do:
      1. All will download the latest "merge" repo from the directory Brad provided;
      2. Mingyu will repeat hodoscope calib;
      3. all other 3 detector timging and calibration will produce new sets of checking plots to make sure things are okay;
      4. DC will need to be re-calibrated after hodo is done.
    4. We discussed calorimeter calibration. Brad noticed a lot of params are different between different kinematics. This should not be the case. We noted this is probably because when normalizing E/p for electrons to be 1.0, the params of adjacent blocks are highly correlated (or, there are multiple minima for the same fitting). In principle, if the blocks had been gain-matched (they were, someone mentioned Simona did that using pion or MIP peaks), then there should be little variation in the params from block to block and from kine to kine. We decided to use pion peak to check that all blocks are aligned. (We have a lot of pions!) Melanie will make the plots, basically, use NGC to select pions and plot E/p for each block.
    5. Brad requested to have the latest optics in the repo. (Mark said 1-2 weeks, Brad: that's okay).
    6. Scaler readiness - Carlos: Took over from Melanie and working on adding the helicity scaler tree to hcana. All other scalers are already in the tree but the helicity scaler is different. (We noted if Hall A can do this in Podd then C should be able to do this in hcana).
    7. PID - Melanie reported on PID (see elog21:
      1. PID of DIS low momentum setting. Pion contamination is higher (as expected for the low momentum). Can achieve 1% pion contamination with electron efficiency 90%.
      2. Pion sample is still not clean - really need to add "low cuts" to both preshower and Etot to reject the background. WE noted there is a strutured background in the ps vs. Etot 2D histogram that looks artificial. Maybe some integer problem? should we look into this?.
      3. Note: checking the calorimeter params and make sure all pions line up in all blocks may help to narrow the pion peak and reduce pion contamination.
    8. Analysis script (lower level) readiness: extracting electron and pion counts with all PID, acceptance, and good track cuts. Cut beam scraping. Cut beam trips(?), kinematic binning. Melanie volunteered on this because she had worked on this previously.
    9. Simulation: We have the latest mc_single_arm and can now make a plan:
      1. DIS: Bill mentioned for F2, they take Tree from mc_single_arm, and combine it with DIS cross section xs(E',theta) that has both internal and external radiation included. (So, external radiation is a simplified version taking the zreact=0, since there is no Z dependence in the model xs). Then, take (Data)/(MC)*(xs)_rad to give "measured xsec (from data)".
      2. For QE, we will have Born xsec and Asymmetry, not radiated. But we can do the same since mc_single output Tree has both vertex and radiated kinematics.
      3. For elastic we will need to modify mc_single_arm. (XC, Mingyu?) -- Thank goodness it is still in Fortran!
      4. What do we do if simulated FP distribution does not look the same as data? This is important since we will need to simulate scintillator plane on/off to match data. Answer: cut out region where the disagreement is large (that's nearly all negative z region).
    10. General discussion on analysis plan: 1) We would like to see if the 121A group can contribute manpower to elastic/QE analysis. 2) We will need to proceed to simulation work. Murchhana will pick up this for d2n; 3) Junhao will start look into pressure curve analysis (no simulation needed, all relative analysis), principles and script preparation.
    11. Moller reanalysis: Bill has a list from Dave a while ago. It will take about 4 person*weeks. Now we have 3% Moller. With reanalysis can get to 2%. Will see if can do this with Murchhana in the next 3 months (complete by early spring).
    12. Other: Xiaochao sent QE calculation need to 3 theory groups last week. For details see elog19

  • 11/03/2020 A1n/d2n Analysis
    1. All 3 DNP talks went well!
    2. Mark on HMS optics:
      1. Found a problem: xtar had been fixed by cosy (cosy-based recon from data) in the optics fitting code. This can be a problem because xtar can be at cm level for long targets and if cosy model is different from real situation. Tried fitting xtar in the code and it seems to have made positive z foils much better (no more curved pattern for xsieve vs. delta in the negative delta region!). On the other hand, neg z foils still have the same problem as before.
      2. Plan: will consider using 3rd order coefficients for xtar fitting (since only odd power will cause a difference between pos and neg ztar).
      3. Overall, looks like the problem is half-solved and we are looking forward to further progress on this next week1.
      4. Note that this effect is large for HMS because (dxtar/dztar)=(1mard/1mm)*xtar. The slope is only about 1/4 of this value for SHMS but Mark will contact Holly (not at the meeting) to see if SHMS can be improved too.
    3. Brad: still need the foundation of what was used in everyone's calibrations. In progress...
    4. Xiaochao proposed we set a goal of having everything ready to go (repeat calibrations, iterate caibrations, and finalize optics) by Thanksgiving. This way we can get the scripts ready in December and launch pass1 replay by X'mas, take two week of vacation, and come back ready to analyze! Everyone agreed. (IN other words: if we are not ready to replay all data, then no X'mas...)
    5. Melanie presented SHMS PID that now includes NGC efficiency and pion contamination using A1n's high momentum DIS data, see A1nd2n-ana-log18.
      1. NGC pion rejection is very low, around 10. Typically should be 100-500. Could be that the pion sample cuts are not tight enough (there are only the high cut each on prsh and sh, not lower cut. Suggest plotting the pion region larger (and with NGC "less than something" cut to focus on the pions if possible). Some typical Hall A plots for pion and electron samples can be found from the 6 GeV Hall A test run page: LHRS and RHRS. To be exact, the links are LHRS samples, and RHRS samples.
      2. Brad also suggested cutting out the mirror edges to see if that caused the low pion rejection. This is for understanding the problem at hand only, since for analysis we will probably include all regions.
      3. By adjusting preshower, Etot/p and NGC cuts, now have reached 0.43% pion contamination while electron efficiency is still at the 95% level or above. This is very good. Will proceed to low momentum next (pion contamination may be higher and difficult to suppress to the 0.4% specified in the proposal), and then HMS.
      4. We are reminded that we do have pion asymmetry data from 6 GeV A1n, g2n, and d2n data. With the high pion background for our DIS kinematics, applying the correction should not be a big problem.
    6. Since we finished all discussions relatively early, we spent a few minutes to discuss the meaning of the subscripts in sigma_1/2 and sigma_3/2. Here are a few references in which these are defined: In summary, the subscripts refers to the helicity of the photon-nucleon system or the total spin projection of the photon-nucleon system.

  • 10/27/2020 A1n/d2n Analysis
    1. Mark: tried single-foil fit but it didn't work any better. (Terms are very different but do not agree with data). Xseve vs delta has the same problem as before. We discussed why the optics does not work. Could it be that too many high-order terms are included and we do not have data to fit them all? Need to continue investigating.
    2. Mingyu presented hodoscope timing checks, but all of them have shifts. We do not understand why. But while discussing this, we realized that: see next item
    3. The calibrations of detectors may not have done in sequence. For example, Junhao's DC calibration was done using the old hodoscope params, not Mingyu's latest. In principle, hodoscope and DC (or any other correlated detector pairs) should be calibrated in iteration (hodo->DC->hodo->DC etc). Brad will talk to the students, figure out the condition under which each detector was calibrated, and figure out a master plan for going forward. (Will need to repeat some calib work). Meanwhile Carlos will update his document to reflect the flow of the calibration work.
    4. We spent about 1.5 hours doing dry run for the upcoming DNP talks.

  • 10/20/2020 A1n/d2n Analysis
    1. DNP talk will be in 10 days. WE have 3 talks -- Melanie, Mingyu and Murchhana. Plan to have a first draft of slides within the week and will discuss/circulate by email.
    2. Holly on SHMS optics: After moving the beamline window to -32.4cm, the 30-deg -20cm foil looks much better but there is still offsets in the foil Ztar position between different angles, see hztar_global_shift.pdf. We asked to document the different versions of optics (1 global, 1 for 30-deg only and 1 for all other smaller angles).
    3. Mark on HMS optics:
      1. the Ztar all look super good (13.5, 16.4 and 30-degrees) with max offset at 2mm. However the angles are not good for foils aways from zero, in particular the xsieve (oop angle) for negative delta. The focal-plane distribution of data different wildly from simulation - definitely something in the magnets that was never understood. (Mark suggests that when we do reconstruction may as well set oop angle to be zero.) We suspect missing or insufficient cross terms in delta and/or xsfp.
      2. Previous optics were good to (+/-10cm in z)*sin(20-deg) which gives us +/-3.42cm only.
      3. li>
      4. Next step: Mark will try fitting a single foil and see which terms stands out, and may give us some ideas for how to fix this.
      5. Mark also pointed out it is hard to ID holes for negative delta runs though he thinks they are still somewhat okay (i.e. can still see hint of the missing hole on the sieve, which helps to orient the picture).
      6. We should try to fix the problem, for the sake of understanding the spectrometers.
    4. Junhao: checked DC timing cuts for d2n sample runs, all A1n cuts are good to use (see d2nlog#10; Found a bug in DC calib code but it does not affect the current calibration work.
    5. Melanie presented Calorimeter PID analysis, see ...
      1. On pion contamination: presented 3 sets of results: 1 without NGC cut, 1 with NGC (Npe above 3) cut, and 1 with both NGC (Npe above 3) and preshower (above 0.04) cuts. The lowest pion contamination is at 1-2% level. This is high, the 2006 proposal stated we need pi/e below 0.4% which roughly requires pion rej >1E3.
      2. Will move on to NGC PID analysis.
      3. BCM: done!
      4. Calorimeter timing done (for both A1n and d2n) -- note: Calorimeter DiffTime check was never concluded for A1n, last entry on this was on 6/30.
      5. Checked calorimeter calibration params using d2n runs and presented dE/E for d2n. Both consistent with the NIM (or future NIM) plot though both are on the higher side of established data. We can move on.
    6. Mingyu presented slides on simulation of windows. These used Jixie's latest mc_single code. Still some small problems to fix on ztar for HMS.
    7. Murchhana presented cherenkov timing cuts (done!). Also NGC calibration using 4 d2n runs. These can be compared with the NGC parameters for A1n presented on 8/11/2020. We discussed whether we should use cosmic or production setting results, and if we use cosmic, which one to use (there was two cosmic runs used for NGC calibration, one during A1n and one during d2n.)

  • 10/13/2020 A1n/d2n Analysis
    1. Junhao showed progress on DC calibration using defocused runs (see A1n+d2n combined elog-2). Modified t0 fitting procedure to use integrated t0 histogram to smooth out the fluctuations. This seems to produce good and reliable t0 results. Drift distance plots are plotted in liny now and seem to have peaks on both left and right hand sides but both are due to how the program bin data and do not reflect problems. Residual plots seem to look good. Residue plots look okay but SHMS has long tails (disappear after electron cuts). A1n runs show pretty good residues. Given that the residue plots are the main criteria, the DC calibration may be considered done (MJ). Should make these DC plots for a few sample runs throughout A1n+d2n run period to check stability.
    2. Murchhana checked all reference timing cuts for d2n, see d2n log8. Brad thinks these cuts are good enough so we should move on.
    3. Next step for d2n, detector timing cuts: We will do the same as for A1n: DC-Junhao, hodoscope-Mingyu, NGC-Murchhana and Calo-Melanie (confirmed by email). Melanie already did HMS calorimeter calibration for d2n, see Melanie_HMS_d2n_calib_short_20201014.pptx and will move onto SHMS calorimeter.
    4. Update on survey results: Jianping presented the spreadsheet on the survey: D2N_Compass_Summary.xlsx from 10/6/2020, D2N_Compass_Summary_Update.xlsx from 3/19/2020 and corrected. These have been entered as Combined elog3. The data shown still need "translation" into actual HCS. Some main take-away messages are:
      1. Cell window position of ref cell and pol3He cell do not match and differ by as much as 1cm. We will have some fun analyzing these data.
      2. The flange front of the upstream beamline window is at -32.1cm so the window should be even more upstream (Brad or JP will fish the flange info out of 400+ CAD drawings).
    5. Both Holly and Mark will have to refit the optics with the new beamline window position.
    6. Michael Nycz submitted an entry on beam energy summary for both A1n and d2n, see Combined elog1

  • 10/06/2020 A1n/d2n Analysis
    1. Junhao showed progress on the DC calibrattion (on d2n elog). Drift distance plots seems to be flat but need linear scale. HMS plots look "flatter" than SHMS's. Mark suggested (1) using defocused runs to get all t0's; (2) check some t0 wires "by hand" and refit if necessary. (3) We discussed what to do for the wires on the edge and Mark + Carlos suggested using a combination of "by card" and "by wire/fit by hand". (Note recall Mingyu's hodoscope calibration changed more on SHMS than HMS which explains why SHMS DC needs more work too).
    2. Optics progress: not much, global fit still does not reproduce all ztar well. At the minimum we can use two optic matrices, one for 30 deg and one for all other (smaller) angles. But survey results for the upstream beamline window position may help - waiting for now.
    3. Melanie talked about pion contamination analysis but needs more work. A one-slide suggestion for rhow to get the pion contamination factor can be found at PID_EC_1page.pddf.

  • 09/22/2020 A1n/d2n Analysis
    1. Discussion on optics readiness: Following last week's plots, Holly thinks there is problem in SHMS z/ytar recon for the upstream foils. Each indiviudal setting optimization is fine but global matrix does not recon -20cm and -30cm beamline windows well. 1D ztar all angles from global matrix. Mark suggested to not use the beamline windows for optimization, reasoning that these are on the ytar edges and may be affected by the momentum setting of the magnet. This also means survying the beamline window (which we discussed at the meeting) will not help.
    2. Also discussed why center foil seems to be at -3mm in z. Seems this is "on the edge of survey pointing result". On the other hand Mark showed one slide on SHMS mispointing, showing a consistent difference between survey (red points) and what data indicate (green) which indicates there may be a real 2.6mm z offset of the foils. - need confirmation from HMS side and more survey of both SHMS+HMS mispointing around 18-20 deg will help. - Brad+JP will work this into the survey plan.
    3. We discussed whether Junhao should complete DC calibration. The consensus is that Junhao should run the calibration code and produce quality-check plots for the DC. If they look reasonable then we will not need to redo the calibration.
    4. Melanie presented SHMS calorimeter PID analysis, see PID_9.21.20.pdf
      1. Overall very reasonable results. we discussed the large background that triggers NGC but has etracknorm below 0.2. JP suggested these are knock-on electrons from pions hitting the entrance window of NGC. The cut of etracknorm.gt.0.2 thus should be applied to the clean electron samples but not clean pion samples (since these below 0.2 events are pions).
      2. Next to do is to find out pion contamination fraction vs. cut.
      3. Then will move on to NGC PID and HMS PID.
      4. Other to dos: wrapping up BCM calibration, replaying an empty cell run for window check -> later.
    5. Mingyu started on using mc_single code for the windows.
      1. First try changing the input file works but the recon z does not look right. Mark reminded us that the recon optics of mc_single was changed for both HMS+SHMS in early summer.
      2. Mark's answer: For SHMS you need to down load this recon file and make a symbolic link from this file to src/shms/shms_recon.dat; For HMS you need to download this recon file and make a symbolic link from this file to src/hms/recon_cosy.dat
      3. Once recon matrices are updated, next step afterwards will be to add DIS cross sectionsf for the windows (so far 3He is hardwired in our version of the code).
    6. Ddiscussed QE asymmetry calculation. Xiaochao will contact the theory groups after collecting all needed kinematics from Brad and the elastic people. Other theory work includes DIS d2n (Todd being the contact) and A1n extraction.

  • 09/15/2020 A1n/d2n Analysis
    1. Melanie discussed her SHMS PID analysis but the current status is that she will check the SHMS acceptance first before proceeding to more detailed PID. When checking the acceptance cuts (for A1n 30 deg data) we noticed that the windows are still quite wide even though the latest optics from Holly was used - the upstream window is little narrower than previous online analysis but not much better, and the downstream window completely disappeared while it was sort of visible (like a bump) in previous replay. The phi (horizontal angle) distribution is quite off-centered though it could be physics (JP) but XZ thinks it's too slanted for DIS. We will need to first make sure the optics are done correctly before doing acceptance. Some ideas for how on proceed:
      1. Melanie will replay some recent d2n runs and compare to the online replay to make sure we see similar window distributions -- this is done, see screenshot Screenshot%20from%202020-09-16%2012-22-57.png, where the left plot is A1n 30 deg using old (online) optics, middle plots are A1n 30 deg using new (latest) optics, and the right plots are d2n 18 deg data using the new optics. It appears that (1) The upstream ztar shows better resolution in the new than the old optics; (2) A1n doesn't see as narrow window peaks as d2n due to the larger angle; and (3) there seems to be a small bump for the downstream window using old optics but it completely disappeared using the new optics -- emailed Holly+Mark and still investigating. (see next line item for the full picture).
      2. Melanie will replay some empty cell runs and attempt to do subtraction of windows using data. The results can be compared with Mingyu's simulation (3He gas only).
      3. Mingyu will work with Xiaochao on adding windows to the simulation. We will need this for offline analysis eventually.
    2. Followup on optics:
      1. Holly's reply: the optics data used were: @30 deg, zFoils at 13.34,0,-20 and @18 deg, zFoils at 20,13.34,0,-20,-30. So, in general, we are optimized to cover a yTar range from -6.7 to 10 cm. -> will look into Dec optics data where the +20cm foil was still present to see if can learn anything.
      2. Two new plots from Dec optics runs 9611-9621 (these didn't have enough statistics for the sieve optimization but did have 7 foils evenly spaced at 0, +/-6.7, +/-13.4 and +/-20cm: 2D delta vs ytar, 1D ztar.
      3. Here is a plot for 3-foil ztar using the new optics data that shows the -20cm foil is reconstructed to -19.2cm. One more colorful plot on ztar: 1D ztar all angles. Holly will look into what is causing the shift (for the -20cm foil and the central foil too that shows slight shift - survey??).
      4. Jianping commented that we should see if there is a shift in ztar between the empty cell and production cell since we need to subtract one from the other for the windows.-> back to Melanie's to do: need replay an empty cell run.
    3. Melanie showed results for BCM calibrations, see 9.15.20_BCM.pdf. Discussions below:
      1. The main observation (see slide 12) is that BCM1 and BCM2 have larger uncertainty in the gain than BCM4A,4B,4C. There is also a slight difference between HMS and SHMS calibration results. We suggested comparing the HMS vs. SHMS and take the difference as an uncertainty, and compare it to the gain uncertainty from the calibration itself. Also suggested comparing to Dave Mack's online BCM calibration results in docid=1040, which was linked to the 2/20/2020 BCM calibration hclog entry hclog3789823.
      2. We discussed which BCM to use and the asnwer is to use BCM1 or BCM2 and not BCM4ABC. Dave Mack's answer is "BCM1,2 have better RF cables, so when the temperature changes in the hall and cable run, the gain changes less."
      3. Dave Mack pointed out the nonlinearity at the very small current values, see ResidualPlots.png
      4. We are still missing Unser calibration. Dave M: I'll do the wire calibration when the run is over. If the Unser gain changes 0.1% +- 0.1% I won't update the bcm calibrations of course.
      5. While discussing BCM scalers, Carlos reminded (or told) us that the G0 scalers had overflow/bit-flip problems. These were discoveredd in April 2020 so all A1n run period had the problematic scalers. Steve Wood's summary of the G0 scaler is at Scaler_Problem.pdf. Discussions with Carlos (after the meeting) led to the following to do items (Melanie?):
        1. Add diagnosis lines to hallc replay (red texts on slide 4), run replay on a few A1n runs (choose some representative runs such as one elastic, and one run per DIS setting) and see how big the problem is for A1n data (ie. see how many warnings the replay spits out. Carlos suggested that "One would have to look at ~line 499 of hcana/src/THcScalerEvtHandler.cxx and add a "cout" statement as it is done in slide 4 to identify when this occurs. Maybe, something more obvious than a "cout" might be needed to clearly identify the problem."
        2. Check the scaler map and figure out what channels are affected (the g0 scaler modules were used for HMS slots 11+12 and SHMS slots 12+13).
        3. Carlos mentioned that most users don't actually look at the scaler reads individually, but rather use a code written by Sanghwua Park (https://hallcweb.jlab.org/doc-private/ShowDocument?docid=949) to determine the averaged beam current cuts. - documenting this so we can study it later.
        4. Brad confirme that most scalers were copied on both spectrometers so that's another way to fix the problem.
      6. Murhana showed online asymmetry results for SHMS kin X. Overall the transverse asymmetries showed consistency and nice sign flip. The longitudinal asymmetries seem to have the same sign between two settings. This is possibly due to the longitudinal runs were a mix of March and Summer running, and some beam settings may have changed. Upon following up with Dave Gaskell who instructed us that the only variable needs tracking is the VWienAngle (which correlates with FLIP LEFT/RIGHT), and comparison of two halog checklist entries (halog3802375 and halog3849189) confirmed that the summer runs went back to the pre-Feb.15 setting of VWienAngle=+88, while March runs had VWienAngle=-90. According to Dave Gaskell, we have all the information needed if we have these 4 variables: "HWienAngle, VWienAngle, Phi_FG, and FlipState". In practice, VWienAngle correlates with FlipState and Phi_FG was used for Qweak but looks like not anymore, so just VWienAngle alone should be good enough for A1n/d2n..
      7. Carlos reported working on adding helicity-gated scalers to the root tree in the Hall C replay. We will need this to cut out beam trips.
      8. We discussed starting offline cooking soon. DC calibration is not done (Junhao) but it may not be that important and online version may be good enough. We will need to make sure that optics is done correctly (recall we are also waiting for more optics data at 18 deg) and that Carlos' helicity decoding of scalers is in place.
      9. Jixie mentioned the python script for generating fields for all A1n/d2n kinematic settings is completed (see link in last week's minutes). Note: recall this may not be final because we are expecting post-run 18-deg HB field mapping. Will plan for post-run field mapping work to start next Monday.

  • 09/08/2020 A1n/d2n Analysis
    1. Holly reported that she had checked the delta-scan runs taken on SHMS but found the data to be not optimal for improving the optics due to multiple-scattering. MS from target scattering chamber and the NGC (in front of DC) made the carbon peaks - elastic and the 44 (or 4.4?) MeV peaks -- to not be distinguishable from each other. Because the current delta optics of SHMS were based on data taken with the NGC removed, they are superior to our latest optics delta-scan data. Conclusion: no need to do further optics delta calibration using A1n data. We are still waiting to see if more data can be taken during d2n to improve statistics of ytar and angle optimizations (mostly in the very negative delta region). Brad has made a note of this for the runplan.
    2. On Optics, latest SHMS matrices from Holly have been added to the d2n online replay: hclog 3831644.
    3. Brad found a misfunctioning scaler module on the HMS side, see hclog3841229. Not sure when the problem started, could be for all A1n/d2n running up to the hclog entry date. For offline analysis, we will need to be careful when using HMS scalers. Fortunately the same scalers are copied to SHMS DAQ and can be checked using SHMS runs' data.
    4. Jixie wrote a python code which can read the field simulation from TOSCA and combine them to field maps, see TOSCA_simulation_for_Pol_He3_Field. The current maps are for SHMS at 18 deg d2n coil setting, and more TOSCA files will be added. The goal is for everyone to be able to run the code and produce all field maps needed for both target and data analysis offline.
    5. Melanie sent in slides (she couldn't attend synchronously) on SHMS Calorimeter PID, see Melanie_PID_9.8.2020.pdf. Xiaochao sent back video comments after the meeting.

  • 08/18/2020 A1n/d2n Analysis
    1. Xiaochao asked Donal for how to calculate Npe for NGC, and received a code pmtfixed.c. References can be found here: shms-cerv6.pdf. He commented that we should not see a big change in Npe for the range of A1n kinematics because the only momentum-dependent factor that goes into Npe calculation is the beta value of the electron.
    2. Holly reported workring on the optics, see shms_optics_ext_18aug20.pdf with the full document at docdb-1075.
      1. Focused on achieveing a global fit for SHMS optics, used all 5 pass optics runs from 11-deg and 30-deg (A1n) and 14.5 and 18 deg (d2n). Results look very good. The "tree" structure is almost not seen at 30-deg.
      2. Still checking how offsets are implemented - are they added in the optimization or after?
      3. Would be nice to know if earlier expereiments need to rerun their replay using the new optics - will look into this
      4. Will start analyzing the delta-scan runs taken with 8.5 deg in Jan. This should not interfere with angle and ytar optimization (which are just completed) but will double check.
      5. Optics run list is from Jixie -> Jixie will need to double-check all optics runs are in the list.
      6. Do we need more optics data? - if more data are taken on HMS, can use more 18-deg optics data on SHMS.

  • 08/11/2020 A1n/d2n Analysis
    1. Detector analysis
    2. Murchhana did the HMS NGC calibration using both Methods 1 and 2 but the results are not consistent, see HMS_CER_Calib_method1_murchhana.pdf and hcer_calib_method2_murchhana.pdf.
      1. For Method 1 alone, different runs give slightly different SPE values. Jianping recommended using SPE obtained from a production run for the final calibration.
      2. Mark suggested calculating expected Npe using known gas pressure, particle momentum, and size of the NGC.
      3. The difference is 60% (5 vs. 8) between the two methods. But Mark pointed out the Gaussian fit won't work for Npe as low as 5 and one must use proper Poisson fit (see NGC note from last week for how to generalize Poisson to continuous function).
      4. As comparison, the online calibration was done in hclog3754419 with Method 1 only. Another entry hclog3753676 shows the full PMT Int spectrum (no fit). This entry hclog3753926talks about pedestal shifts.
      5. Murchhana will try using Poisson fit, and will also try other runs like 5-pass DIS (higher E'), but these are of lower priority than checking online yield for d2n.
    3. Mingyu reported the hodoscope calibration is completed. THe previous problem (with run 10602) was low statistics on two bars not enough to determine light velocity, see slides Hodo_Calibration_SHMS_10345.pdf. THe parameter files are copied here: hhodo_TWcalib_3408.param, hhodo_Vpcalib_3408.param, phodo_TWcalib_10345.param and phodo_Vpcalib_10345.param.
    4. Cameron started doing NGC intrinsic efficiency analysis. Obtained good Npe map. But during presentation we realized he was using a Delta run which was triggered on ELclean. Need to switch to using a 3/4 trigger run instead (we had this only for DIS settings. 1-pass had a couple of 3/4 runs but not with all paddles on for SHMS).
    5. Jixie reported received 8.5-deg TOSCA map for HB field and completed the calculation. Overall the calculated values are in good agreement with his model (within uncertainty). However since he found the scaling factor to match HB current density in TOSCA with the actual field/current data varies with the HB angle (it shouldn't, but the observation is it does), need field mapping for HB@8.5 deg (single coil on) at the end of d2n. Brad will ensure this happens.

  • 08/04/2020 A1n/d2n Analysis
    1. Isaias passed by last night.
    2. Holly reported obtained different matrices for the 4 angles. Calibration for delta less than -13% is really difficult so we will probably cut this out in the data analysis. Also looked at 2-pass optics data but found them not very helpful.
    3. Melanie checked DIS runs to make sure the calorimeter resolution stays the same and that we can use the same calibration params for all runs.
    4. Xiaochao's NGC notes: Notes_on_NGC_PID_allpages.pdf, sent to Melanie, Murchhana and Cameron to follow.

  • 07/28/2020 A1n/d2n Analysis
    1. Optics: Holly and Mark will report on this next week;
    2. Melanie presented her version of Calorimeter calibration results, see Melanie_ECal_Calib_7.28.20.pdf. Overall conclusion is the same as last week and will move on to PID analysis. For the record, here are the final calibration param files: HMS_final_cal_param.zip and SHMS_final_cal_param.zip
    3. Maria presented further study on how NGC SPE calibration depends on mirror cuts, see maria_hallc_7_24_2020.pdf.
    4. Xiaochao has an explanation for this: when the cut is loose, the Npe distribution within the cuts can vary and the PMT spectrum is not a single Poisson distribution. This widens the peak since the peak center is porportional to Npe but the width is proportional to sqrt(Npe), and cause the fitted Npe value (mean/sigma)2 to be smaller than the real value and the SPE calibration is larger. Will write a document on suggested PID analysis plan.
    5. Brad: beam is on track to start this weekend.

  • 07/21/2020 A1n/d2n Analysis
    1. Progress on HMS optics by Mark, see docid=1073 where the last file Poltar%20HMS%20optimization.pdf is a presentation for today.
      1. First round of optimization done for HMS optics taken at 13.5, 16.4 and 20 deg. Still need to include 30 deg A1n data
      2. Negative delta still has curvature and angle optimization need more iterations. We need to make a plan on this.
      3. Would like more data with foil at +13 moved to -13cm and at 20 deg. Brad will accommodate this into the runplan.
    2. Jay said something about the accelerator and then darted out to another meeting.
    3. Maria reported on SHMS NGC calibration, see hallc_7_21_2020_ngcer_calib_satnik.pdf. This should be compared with the online calibration by Burcu posted at hclog3749987. It looks like the param values are all smaller than the online value. Need to study if this is important? Next Maria will move on to using defocused runs.
    4. Murchhana: will travel to JLab this week.
    5. Cameron: showed a fit to SHMS calo resolution, see SHMS_Cal_Calib_Res_Fit.pdf. We think the quality of the calibration is good enough. The data from F2 and D2, and the simulated curve are all from Vadan Tadevasyan, see resolution.dat and README. Did not find data that can be used to make a similar plot for HMS.
    6. Jixie/Brad:
      1. need correction coil settings for 1-pass SHMS@8.5 deg and -2.1 GeV/c setting. We will try two approaches in parallel: Brad will ask Steve for a TOSCA geometry file; and Jixie will run his model to come up with a best estimate.
      2. Because we never had field mapping at 8.5 deg, will be good to do some measurement with HB field only, perhaps after the run is completed. It does not look like we can do this before the run restart.
      3. Jianping emphasized the field requirement for 1-pass calibration is not as stringent as d2n production.
      4. need correction coil settings for all d2n kinematics. Jixie will double check.
    7. Brad: checked a few hardware things noticed in the analysis:
      1. both HMS and SHMS FADC reshaper fixed, see hclog3815712 (not sure what difference we expect to see in the data?)
      2. SHMS NGC DiffTime in PMT4 is now fixed (now similar to other PMTs), see hclog3815555 and hclog3815556. The time delay of PMT4 (3 in C++) noted on 2020/5/26 minutes should not be there anymore. For d2n analysis need to recheck DiffTime cut.
      3. investigated SHMS shower timing (see hclog3815557but turns out analysis (see 6/2/2020 minutes below) showed different DiffTime for two HMS calo PMTs. Will investigate those next.

  • 07/14/2020 A1n/d2n Analysis
    1. Mingyu presented progress on hodoscope calibration:
      1. Followed last week's suggestions to use cosmic run, but the fitted time-walk didn't change much. Plotting data vs. fit for larger range didn't show problem (though cosmic data had low statistics and hard to compare), see Hodo_Calibration_HMS_3410.pdf and Hodo_Calibration_SHMS_10604.pdf.
      2. The fitted c1,c2 parameters are all different from the old parameters, in particular c1 changed by quite a bit for both HMS and SHMS. This could be due to time shift (cable length change??). See Hodo_Calibration_Time_Walk_Correction.pdf
      3. HMS is fine (done, see last week's report). But SHMS still has problem with the calibrated beta being below 1 and wide. Use old time-walk parameters and the subsequent calibration steps did not cause widened beta. See This means the problem is in the time-walk step. Hodo_Calibration_Time_Walk_Correction.pdf. Will follow up by email.
    2. Calorimeter: Cameron did 3 iterations for each of SHMS and HMS calorimeters. In each step, merged new calibration parameters with defocused ones (for blocks not illuminated by DIS runs). See HMS_Calorimeter_Calibration.pdf and SHMS_Calorimeter_Calibration.pdf. It looked like iteration beyond #0 are not needed though.
    3. Melanie presented results on calorimeter calibrations, see Calo_calib_Melanie_20200714.pdf. Results are quite similar. We suggested checking a few runs on or just before March 13th to make sure the same calibration constants will work for all runs of A1n. Then we will be done with calo calibration!
    4. For comparison, here is talk by Vaden on the calorimeters that contains plots for both HMS and SHMS: HallC_calorimeters.pdf. A (possibly older) HMS plot can found in Fig.15 of the HMS Calorimeter NIM paper (or arxiv1204.6413). The SHMS plot used for the draft NIM paper is shms_calo_res.pdf
    5. Holly on optics: Looked into other SHMS kinematics and found the optics data to be okay. So we do not need to take more optics data. We hope Mark or Holly will fine-tune the optics calibration as the next step (and maybe complete it before d2n starts).
    6. We talked about analysis plan: Cameron will start looking into Cherenkov calibrations. Once it is done, Melanie can look into full-scale PID analysis. We will wait for Junhao to complete DC calibration, and hodoscope is nearly done (hopefully). We will continue weekly analysis meeting through d2n run.
    7. Target polarimetry: Junhao may complete the NMR analysis couple of months after d2n is done. Mingyu is making good progress (on beam depolarization effect).
    8. Also talked about SoLID rungroup proposal on longitudinal spin for summer 2021 PAC. Maybe Melanie and Mingyu can look into expected results using SoLID SIDIS simulation results, once d2n is completed.

  • 07/07/2020 A1n/d2n Analysis
    1. Holly presented first round of optics calibration for 14.5 deg: optics_run10786_7July20.pptx. Some comments:
      1. the lowest delta slice is -11.5. Can we optimize for more negative delta such as -13 and even -15? - Yes (great!)
      2. Overall the delta vs. ytar looks better (straighter) than 11 deg results shown last week, but it could be due to more spread at larger angle.
      3. Brad asked if we can go through all optics (the other two angles) to see if there is any problem (missing or problematic data) there. If there are we can take more data at upcoming run.
    2. Brad sent out shift signup page yesterday. Emphasized on following new lab MEDCON-5 procedure, reserve ResFac early if needed.
    3. Jixie reported on TOSCA target field calculation: 18-deg is ready, 11-deg model is ready, and 14.5-deg model is being worked on.
    4. Melanie reported on SHMS calorimeter calibration: SHMS_Cal_Calib_7.7.20.pdf. Comments:
      1. the sigma value reached is comparable to F2 results: 4% for higher and 4.4% for lower momentum setting (good!)
      2. After a 2nd iteration, the problematic preshower block looks better, but no major change in sigma value.
      3. Next will try more iterations and chain one more run.
      4. Will also move on to HMS next.
    5. Cameron updated on SHMS calorimeter calibration: got some preliminary results: 5.5% for the higher momentum and 4.9% for the lower momentum settings. We suggested following Melanie's iteration procedure and see if can reach the same sigma values. Will also move on to HMS.
    6. Maria presented SHMS NGC PulseInt plots for method2, with cuts on "no signal in other PMTs" applied. THe peaks are cleaner and can fit with a Gaussian. Some suggestions:
      1. Add x/y cuts and timing cuts, and following the procedure to calculate the calibration constants. Compare with online values and see if they make sense (i.e. they should be close in value).
      2. Mark commented elastic runs may not be great (in addition we had only 1-2 paddles on per layer). XZ suggested using Delta runs instead. Here is the run list on 1-pass Delta transverse runs from Jan 2020: SHMS: 9817 through 9823; HMS 2640 through 2646. We will probably not use Dec runs because they all did not have much polarization.
      3. Repeat the process above on delta runs and compare calibration results (with elastic runs). Delta runs probably work better.
      4. Repeat the process on some DIS runs and compare results. DIS runs do not have lot of electrons (not ideal for calibration) but this serve as a double check.

  • 06/30/2020 A1n/d2n Analysis
    1. Mingyu reported on hodoscope time-walk calibration using DIS runs, see Hodo_Calibration_HMS_3408.pdf and Hodo_Calibration_SHMS_10602.pdf. Q and A below:
      1. HMS: peak is closer to 1.0 but with slightly wider sigma. In principle sigma should be smaller. Mark mentioned it's better to use cosmic runs (hits cover the full bar rather than center of the bar
      2. SHMS: didn't work. new beta peak is way off.
      3. Suggestion 1: plot old time-walk calibration values on top of the new ones. In principle time-walk should not change unless some bars/PMTs are dying. This way we can spot if there is any outlier.
      4. Suggestion 2: optimize the fit range of the TDC vs. amplitude 2D scattered plot (for SHMS the fitted amplitude range is clearly too small), and adjust the vertical plotting range so one can check the fit quality more clearly. Better also plot the fit from the previous calibration on top.
      5. It doesn't look like any single bar can be outlier (since the new beta peak is a nice Gaussian, just not at the right place). But could plot beta for each single bar to see if one of the bars is causing problems.
      6. Once this DIS run's calibration is fixed, can repeat the procedure on a Delta(1232) transverse run, and a low momentum DIS run.
      7. minor point: should add more significant figures to the printing of the fit parameter.
    2. On Cherenkov calibration (Maria):
      1. after not finding SPE peak on SHMS for a while, we asked Mark and it turns out method 1 will not work for SHMS NGC because the PMTs are operating at very low gain and SPE peaks are technically below the threshold. For SHMS NGC we can only use method 2. HMS NGC can be calibrated with both methods 1 and 2
      2. For method 2, we suggested Maria to follow all cuts listed in Simona's hclog entry hclog 3650663
      3. We discussed the purpose of the Cherenkoc calibration. The way it works now, the goal is to have the SPE peak of all mirrors/PMTs aligned. Then one can check if the average Npe for an electron is the same for all 4 (2) mirrors. Of course, if one mirror/PMT has low light yield for electrons then there will be a (x,y) dependence on the PID performance, but that can be done in the analysis at a later stage (dividing x,y into small bins, extract average Npe and efficiency for that bin. One can then include the resulting x,y-dependent efficiency in the MC.)
    3. Mark finished a first round of the A1n 11 deg optics calibration, see slides SHMS%20optics%20Pol%20Helium.pdf. Overall the optics is much improved, but we had some questions:
      1. The data are fit only in delta =(-10,15)% range. Data below -10% are very hard to fit. We asked if it's possible to fit to that range. Answer: not easy to do.
      2. Holly is looking into other angles (30 deg, 14 deg, 18deg) for SHMS, and Mark is moving to HMS optics.
      3. On the 4 angles' optics data, in principle they should all be optimized to a single matrix, but Mark isn't sure if some settings have the magnet saturated. Will optimize the 4 angles separately and compare. If they are similar/consistent, can combine the results to produce a single optics matrix.
    4. On Calorimeter calibration: Cameron got the code working and is trying to change the code such that one can merge the new calibration constants (new but may not cover all blocks) with the previous one (from the defocused run that contains all blocks).
    5. Jixie is working on getting the gradient predictions for d2n. But TOSCA results for the fields from HL, HS coils after coil rotation differ from TOSCA results for HL, HS before rotation (i.e. it's not a simple rotation of the coils. Some iron structure must be at play). Thus it will be useful to obtain field mapping data for single HL and single HS coils at the current orientation to calibrate TOSCA. Of course, these coils don't produce large gradients so one can still use Jixie's model to produce "some" gradient prediction.

  • 06/16/2020 A1n/d2n Analysis
    1. All detector timing cuts are done. We should commit (is that the right word for github?) all final RefTime and Detector Timing cuts to github. Which github site should we use?
    2. David Winters (UVA) presented pull plot analysis for A1n. The idea is to produce pull plots for each xbj bin and for as many run condition "filters" as needed, such as IHWP state, cell name, momentum setting, and spectrometer. See his asymmetry website (slides will be posted later). One can use the plots to examine consistency between different runs (spot problematic ones), and can repeat this after as the offline analysis is carried out.
    3. Melanie noticed there are double peaks in Preshower DiffTime plots for ELreal runs taken in mid March (before the shutdown). They appear to be good electrons. Nobody knew what could cause the double peak. Suggested Melanie to look into projected x vs. y distributions for the second peak. If no problem is spotted then we can probably move on.
    4. Maria presented a few NPE plots from elastic runs. Next step is to use calorimeter cuts to single-out electrons and use method 2 of the calibration. We talked about using cosmic runs (SHMS on 12/10, HMS run 2501? on 12/15, and possibly many other runs) to find single PE peak and use method 1. Brad suggested a script "hclog_search [cosmics]" (with [] the key word to be searched) that runs on cdaq. (Maria will need the two-step verication to access hallgw, to access cdaq from offsite).
    5. hclog entries on Cherenkov calibration can be found by searching "bduran AND calibration". Some important ones are here: hclog3753676 hclog3754419 hclog3753926 hclog3749987
    6. We discussed in general where to find a "master list" of all hcana variables. Mark's method is to use "gHaVars->PrintFull()" in hcana or add this to the end of the replay script. A partial list can be found at Pol_He-3_Analysis_Resources#Calibrations (scroll down to see varialbes for calorimeter, cherenkov). After the meeting, Brad posted another partial list at Pol_He-3_Analysis_Resources#hcana_Analyzer_Variable_Names. As for whether a variable is an array Carlos suggested we can "inspect" on a leaf variable (though we won't know what the index is for. For that we need to read the hcana code).
    7. Will be good if we can compile a "dictionary" for the full list of variables, with their physical meaning, (index meaning), etc. Perhaps Maria can get this started from Cherenkov and calorimeter variables, and each of us can fill in as we carry on the analysis.
    8. Junhao wasn't at the meeting and Mingyu got called into the hall, so we didn't discuss DC and hodoscope calibrations.
    9. Mark: found used incorrect BPM calibration so is redoing it for optics. We asked if BPM calibration is final and the answer is "yes, though someone can doublecheck" (we will keep this in mind but probably not necessary since it's Mark's work...).
    10. Brad: critical-path items for d2n ramp-up from analysis side are optics (Mark) and field gradient (Jixie). Field gradient work is ongoing and is moving towards the right direction.

  • 06/02/2020 A1n/d2n Analysis
    1. Murchhana presented NGC DiffTIme plots for both SHMS and HMS: Detector_time_window_cuts_Cherenkov-murchhana.pdf. It appears existing cuts are all fine to use at this stage. We fully understood the multiple peak structure for SHMS. For HMS we saw some peaks: the black from raw timing are probably signal reflections (and they should be low in amplitude); there is a magenta peak right to the right side of the main peak that are worth studying further (HW if have time). Mark suggested these could be pedestal-related ref-time shifts, but for SHMS this peak is on the left side and disappears for DIS. While the peak we see here for HMS seem to be present also for DIS. (HW: need posting/documenting the final cut )
    2. Junhao
      1. reported on DC DiffTime, here are final slides DC Time Window Study III.pdf. The cuts are pdc_tdc_min_win=-13200, pdc_tdc_max_win=-10600, hdc_tdc_min_win=-13800, hdc_tdc_max_win=-10800.
      2. looked into DC calibration. The current code calibrates t0 and drift time at the same time. Mark suggest first to modify the code and see if we can fix t0 to previous values (these typically do not change and requires high statistics to calibrate) and calibrate only the drift time.
      3. Also suggest using defocused runs. To replay defocused runs one needs to change the matrix reconstruction file (see hclog 3751655, I believe these refer to directory "go_analysis" (not pro)). Also see hclog 3752641 for Dec '19 calorimeter calibration using these defocused runs (the run numbers are in there). Finally, if one would like to start with the already-replayed defocused ROOTfile, Burcu and Bill chained all the runs together and they are on cdaq. To find them do “go_analysis” (not PRO) and look in the ROOTfiles directory for shms_defocused.root.
      4. Mark showed briefly on how tracks calculated in hcana. I posted his notes at the Study Material list above.
    3. Mingyu reported still working on getting hodoscope timing cuts. -- followup: finalized all hodoscope timing cuts, see phodo_cuts_06_02.param and hhodo_cuts_06_02.param
    4. Melanie was at the meeting but her audio didn't work. She sent in slides on SHMS DIffTime plots, see SHMS_Cal_Cuts_052920.pdf. Still working on Preshower and HMS plots.
    5. Cameron started working on making Calorimeter plots too. See TimeDiff.pdf It appears old cuts are all okay (per Mark).
      1. some channels are shifted, such as HMS layer 2 PMT4+ and layer 3 PMT8+ (layer starts from 1, PMT starts from 0). Brad looked into this and please see meeting minutes for 7/20/2020. We suspect it has always been there, but would be nice to confirm with F2 and other data. Brad made a note to check cable length. Overall not important as long as it is not due to a software mistake.
      2. HMS raw (calculated) spectrum seems to have a 2nd peak on the right. Bill said this should go away once a certain cut is applied. Will ask Bill for the cut.
    6. Xiaochao looked into calorimeter calibrations, and compiled information from Bill and Burcu, see Notes_on_Calo_Cal.pdf. Will pass this on to Cameron and Melanie.
    7. Mark continued working on optics
    8. Jixie will work on 18-deg TOSCA data this week.

  • 05/26/2020 A1n/d2n Analysis
    1. Melanie and Murchhana are both absent today so we will discuss their work next week.
    2. Mark on optics: setting up cut files for the optimization. Optimization of angle and Ytarg is done simultaneously. Delta optimization may be difficult (may not improve). Will consider let Jixie to do SHMS and Mark can move on to HMS optics.
    3. Wrapping up RefTime cuts: for SHMS we will use the values shown on slides SHMS_HMS_Ref_Time_Cuts_050820.pdf: pT1/2: 3400; ROC2: 4200; DCREF: 14400; For HMS we will use the values shown on slides Reference%20Time%20Cut%20HMS%2020_05_05.pdf: hT1/2:1400, ROC1: 3400, DCREF: 20000
    4. Junhao's plots on DC TDC time spectra DC Time Window Study II.pdfwith QandA and HW:
      1. what are differences in ndata vs. nhits vs. array size? Conclusion from slide: they are the same thing
      2. What cut to use in determining "clean" spectra? Answer: use nhit=1. This is because for Hall C DC, a good particle event will trigger only one wire per plane. (unlike Hall A VDC which is 3-5 wires due to particle trajectory at 45-deg w.r.t. planes. For Hall C they are perpendicular). HW: : make plots with nhit=1 cut and make sure the minimum of plotting is 1 or less (for logscale)
      3. Junhao studied electron count vs. DC timing cut but that does not provide an answer to "how tight should the cuts be?". At a later stage we will study DC efficiency for the cut chosen. For now since our main production is low rate DIS, suggest cutting not too tight (on slide 6 probably somewhere between red and green), and HW: : check a sample set of runs throughout the run period.
      4. slide 7: Reason for "dip" to the left of the main peak on SHMS DC DiffTIme: (Brad) possibly due to a random (background) event comes in early, and the DC can't take another hit during deadtime (should be 30-40ns). The exact deadtime can in fact be studied by comparing "dip area" vs. rate or simply "ration of dip area over main npeak" divided by rate. (This can also be a HW )
      5. Why there is no "deadtime dip" on HMS? Rate? Don't know.
    5. Mingyu's plots for hodoscope TDC time spectra HODO time cut 2020_05_26.pdfwith QandA and HW:
      1. After fitting each channel with a Gaussian and plotting +- 5 sigma, it seems existing cuts for some channels are okay but are eitehr too tight or too loose or not symmetric for some other channels.
      2. Brad suggested due to TDC intrinsice 4ns resolution (even if the DiffTime algorithm tries a fancy algorithm to extract the exact time), we should avoid using cuts of 8ns (full width). A min of 12ns (full width) is suggested and 16ns is probably better.
      3. HW : Perhaps Mingyu can use already fitted MEAN for each channel and set the cut at (MEAN+/-8)ns (rounding the MEAN to whole ns value is okay), make the plots and see if this works for all channels. If they do then make plots for a few sample runs throughout the run period.
    6. Xiaochao's report on SHMS NGC timing multi-peaks: SHMS_NGC_DiffTime1.pdfwith QandA and HW:
      1. PMT3's DiffTime is shifted from other 0,1,2 by about 12ns. Possible cable length difference? (Brad may look into this later) -- note: this was fixed at the hardware level on 7/21/2020. See 7/21/2020 meeting minutes.
      2. Cross talk peak (at about +13ns for PMT 0,1): These are background but Mark suggested we don't have to cut it out. (Cutting this requires a tight cut that may affect analysis if the meain peaks shift). Including the cross talk at most adds 1/2 to Npe value and its noise.
      3. Peak at +26ns for PMT 0,1,2 (larger for PMT3): should be included because these are electrons, only the RefTime get shifted due to another event coming in within the same time window.
      4. Peak to the right of main peak: appear to be pions but no need to cut out at this stage.
      5. HW for Murchhana: propose a basic set of NGC timing cuts, make plots for sample of runs throughout the run period.
      6. HW for Murchhana and Maria: Do the same analysis for HMS NGC DiffTime.
    7. Bill's work on F2/EMC analysis: deadtime.pdf and we probably need to do the same for A1n/d2n.
    8. Jixie finished a first draft of the field mapping and TOSCA analysis, see A1ND2N_Field_Mapping_20200526.pdf. Work is still ongoing to check all figures, texts and to post field prediction results for A1n, but the basic structure is very well described. We will discuss this in 1-2 weeks from today and will probably ask Jixie for a short presentation.HW for everyone:: read the note. Meanwhile, Brad mentioned Steve already produced TOSCA files for 18 degrees and HW: Jixie can start looking into it to predice the field gradient map (and current settings) for d2n.
    9. Given that Melanie has starteding doing scaler analysis, sugget Cameron work on Shower/Preshower DiffTime plots and see how that goes. Meanwhile Maria can work with Murchhana on NGC DiffTime (and possibly calibration) to learn PID.

  • 05/19/2020 A1n/d2n Analysis
    1. Maria S. will join us this summer, possibly carrying out PID analysis. Welcome!
    2. Melanie's presentation on SHMS RefTime, separate cuts for elastic (EL-clean) and DIS (3/4): SHMS_Ref_Time_Cuts_051820.pdf. At the last meeting we proposed to use different cuts for different trigger runs. However, when we examine the cuts today, we realized that if we do not plan to use "too tight" cuts for 3/4 DIS runs, then there is no need to use different cuts because EL-clean cuts will work for all type of runs just fine.
    3. Mingyu's presentation on HMS RefTime cut determination: Reference Time Cut HMS 20_05_15.pdf.
    4. HW1: For both SHMS and HMS, choose one set of cuts that work for all runs (ELclean, ELreal, DIS). Cuts for DIS can be loose but okay since there is no background there anyway. (Oops looks like we walked in circles, but hopefully only one circle. need to finalize the cuts and call it done!)
    5. Xiaochao's plots on 3-peak SHMS ROC2 RefTime during elastic running: SHMS_ELclean_Peaks1.pdf. The conclusion is there is probably a flaky cable on the trigger module (most likely L1 Accept since it affects all modules). This is okay as long as our Reftime cuts include all good peaks; Explanation of double peak structure in EL-real trigger: SHMS_ELreal_TwoPeaks.pdf (good thesis material).
    6. Mingyu reported on hodoscope TDC time spectra. Question is whether the existing cuts are too tight (Brad: too tight; Carlos: too tight; XZ: looks like the cuts are already 5sigma so not tight...).HW2:
      1. Fit to each with an Gaussian and quantify whether existing cuts are 3- or 5-sigma or some other value.
      2. make plots for the list of sample runs to make sure these peaks do not drift over time and/or kinematics.
    7. Murchhana's plots on Cherenkov timing cuts: Detector_time_window_cuts_murchhana.pdf and questions questions.png. Discussions:
      1. Variables to plot should be those in Table 4-5 of Carlos analysis note. Calculation of these "Good" variables is nontrivial (not as simple as Eq.6-7), but one should read hcana to understand the exact form of Eq.6-7 that leads to the "Good" variables.
      2. For SHMS ELclean, the cleanest peaks (magenta, Good var with multiplicity 1) appears to have 1 small peak to the right side of the main peak, and there are 3 peaks within the main good peak. We brainstormed some ideas on the cause: the right "shoulder" could be after-pulsing ("ringing of cable"); HW3: The 3 sharp peaks should be studied further (look for correlations with whether there is signal in only one PMT; any position correlation such as hitting between mirrors though one does not expect such visible time difference; or cable cross talks which can be studied if one plot ADC amplitude vs. DiffTime or plot separate ADC amplitudes and ADC pulsetime for 1 vs. 2 hits.)
      3. SHMS ELreal peaks have the same 3-peak structure.
      4. SHMS 3/4 DIS runs appear to have high background (could be simply threshold is too low). Mark noted this is the first time NGC is used in SHMS trigger.
      5. HMS is all good, but some raw cer variables such as “H.cer.adcPulseTime” are missing in the ROOT file. These are not standard variables that are made accessible by the analyzer under normal operations, only in fadc-debug mode. To fill this branch, we have to enable the debug variables for the HMS Cherenkov by adding a line reading “hcer_debug_adc = 1” to the bottom of the PARAM/HMS/GEN/hmsflags.param file. Then this branch will be filled when running with “-m all”.
    8. Junhao's DC TDCtime spectra: DC Time Window Study.pdf, shms_r10614_n200000_dc_twinVSnhit.pdf and shms_r9779_n200000_dc_twinVSnhit.pdfplus QandA:
      1. We didn't know if we should apply a cut on nhit so Junhao studied the DiffTime with different nhit cuts. Turns out this is not useful (or does not make sense). To "clean up" the spectra we only need multiplicity=1 cut (Carlos). We also discussed the requirement of looping over all nhit when filling the spectrum but Junhao is already doing it. To make sure the plots are indeed correctly produced Junhao should HW4: compare codes with Carlos' github.
      2. Last week Carlos mentioned to optimize the DC timing cut, one should study the yield vs. cut (yield can be extracted by applying W cuts for elastic) to make sure the cuts are "not too tight and not too loose". HW5: Need to write a code to extract elastic event yield and plot it vs. a range of DC timing cuts.
    9. Since I can't find this info anywhere else, I write it here: according to Mark, "The ADCTDc DiffTime is units of ns". I think since some timing are from fADC and some from TDCs, the replay script must convert everything to ns before calculating the diff time.
    10. Near-term plan: skipped this -- we may have less uncertainty if we argue instead whether the universe will end with big freeze or big crunch. (But we will meet next Tuesday.)

  • 05/12/2020 A1n/d2n Analysis
    1. Melanie's presentation on SHMS RefTime stability: SHMS_HMS_Ref_Time_Cuts_050820.pdf
    2. Mingyu's presentation on HMS RefTime cut determination: Reference%20Time%20Cut%20HMS%2020_05_05.pdf.
    3. Followups on Ref Timing:
      1. Range for DCREF is too large to see structure. Suggest making range smaller, for example to 5000 total.
      2. HW1: For DIS runs need to choose different cuts because these used the 3/4-trigger and the timing of all ref time signals is different from ELclean runs. (Melanie on SHMS, Mingyu on HMS)
      3. The timing delay between different trigger runs should correspond to the actual signal delay that can be measured on the scope, see two hclog entries: HMS https://logbooks.jlab.org/entry/3752087; and SHMS https://logbooks.jlab.org/entry/3752072. Here is a note to correlate the timing we see in the spectra and the scope observation: Notes on trigger.pdf(Xiaochao)
      4. HW2 (done, see 5/19 notes): Mingyu and Melanie (ideally) should understand all peaks and stuff in each of the reference timing spectra. Double peak can be slow clock resolution (signal fall between two cycles), or two triggers (such as one event triggering both ELclean and ELreal in a ELreal-trigger run), or other hardware problems. These can be understood by looking into all pretrigger or trigger histograms with cuts on one of the observed peaks. Furthermore, one could identify the timing difference observed in the TDC with the scope signals (above). Suggest Mingyu focus on HMS and Melanie focus on SHMS.
    4. Discussion on SHMS ROC2 RefTime 3-peak structure. No solution and only facts: There was 1 peak in Dec, turned into 3 peaks in Jan. Mark suspected an inverted signal somewhere, Simona says no and instead noticed some correlation in NGC signals (NGC rate not propto current, HGC does, etc). The problematic runs were taken with a collimator. No trigger work was done between Dec and Jan. -> done, see 5/19 notes
    5. Discussion on "100ns blocking time" effect on A1n/d2n: -> Xiaochao will join Thursday Hall C commissioning meeting to understand the details.
    6. Summary of Ref Time Q and A's extracted from various email threads: QA_RefTime.docx
    7. Junhao made DC timing signal spectra (for cut study), see hms_r3145_n200000_dc.pdf and shms_r9779_n200000_dc.pdf. However these histograms already had the timing cut applied. Need to remove cuts, remake the plots with nhit cut, and decide on cuts again. Some discussions:
      1. Understanding the shape: These represent the drift time of the ion or electrons to the wire. Left edge is the fastest drift time so it's a sharp peak. Right edge is for slower and slower drifts so has a slow drop and stops at a point where ions/electrons start to predominantly drift to the adjacent wire.
      2. How tight should the cut be? - Carlos: not too tight and not too loose, they have to be just right. The event rate will correlate with the cut and one should study those.
    8. Brad updated on lab status: possibly reopen (partially) soon, then two weeks to start the accelerator and six weeks of running.
    9. Next step: Detector timing work breakdown (same HMS vs. SHMS):
      DC: Junhao, Mingyu
      Hodoscope: Mingyu, Murchhana
      NGC: Murchhana, Melanie
      Preshower and Shower: Melanie, Junhao
    10. HW3: Read Simona's slides on Cherenkov calibration, see wisewordsoncheranalysis.pdf - Melanie, Murchhana.
    11. QandA: how to use VNC to speed up remote analysis? d2n instruction on VNC

  • 05/05/2020 A1n/d2n Analysis
    1. Simona presented a follow-up presentation on the reference timing cuts, see malace_A1n_trainning_session_3.pdf. Some highlights and HWs are listed below:
      1. Suggested values for RefTiming cuts are given as red texts: 14400 for pCDREF1_tdcTimeRaw, 3400 for PT1 and PT2, and 4200 for TRefRoc2. HW1 - done: make the same plots for the latest ELreal runs (after rotation just before the shutdown), as well as 1-2 runs for each kinematics and make sure the same cuts can be applied.
      2. HW2 - done : Do the same for HMS, plot suggested cuts as red lines and send them back (to Simona) for confirmation.
      3. The Time without raw are TimeRaw corrected for reference time and have much better timing resolution. (since it uses a 100ps fast clock rather than a 4ns or 25ns slow clock). (I hope I got this point correct).
      4. The double peak in P1X etc seems to be solved. These came from a software cut in the replay (on shms_TdcTimeWindowMin/Max). After widening the cut (Min from 3000 to 0), the peak at 0 in P1X timing now shows up as a second peak below 3000). Now P1X itself has two 25ns-wide peaks which is expected as a pure effect from the 25ns slow clock resolution. (Note: If not sure what cuts to use, use the widest range possible.)
      5. HW3 - ongoing: try to understand all timing histograms and compare with Carlos' trigger diagram, see docdb-1028. We need to make sure all timing signals are as expected before replaying all data.
    2. Plan to make some progress on getting the scripts ready to plot detector timing signals (step 2 in Carlos' note).
      1. Bill wrote a script showerRef.cpp to plot SHMS calorimeters (shower part), and plots are here: allRows.pdf. Suggest students split up the detectors so each writes an independent script and make plots for a subset of detectors.
      2. We will ask Simona and Carlos for other detector codes they have, though we should write our own code too. Carlos' ref timing script can be found at https://github.com/Yero1990/DEUTERON_ANALYSIS/tree/master/ANALYSIS_SCRIPTS/set_REFTimeCut
    3. Other to do items:
      1. HW0: Read Hall C analysis workshop slides. A list from Brad is given below:
        Starting from this page: https://redmine.jlab.org/projects/podd/wiki/Workshop2018.
        In general, items with a (*) on that page denotes an interactive tutorial component. The page contains a link to bluejean recording at the top.
        Review the following presentations first:
        - Farm Use and Computing Resources Tips and Tricks, and/or this update: Hall A and C computing tips and tricks (and how not to kill the whole farm) -- Brad Sawatzky
        - Overview and Update of the Hall C Analyzer -- Eric Pooser
        - Eric's talk from the 2019 Hall C Winter Collab meeting is also quite useful: Hall C Software, status and outlook
        Then move on to the Tuesday 'Hall C' sessions starting with the git howto:
        - Effective Git use (*) -- Steve Wood
        (Folks really do need to understand how to work with git. They should follow up on the tutorials/howtos Steve mentions in his talk before moving on.)
      2. Links to hcana wiki page: and compiling instruction hcana docs; and tutorial from last year's software meetings: hallc_tutorial/tree/master/Exercise-01; to replay a run with all leaves, do "-m all".

  • 04/28/2020 A1n/d2n Analysis
    1. Everyone is "back from APS April meeting" so we can resume our weekly analysis again.
    2. Before the meeting, Melanie, Mingyu, Junhao and Murchhana sent their timeing plot slides to Simona. Two runs for DIS and two runs for elastic were picked for each of HMS and SHMS: 1-pass elastic runs: HMS: 2608, 2609; SHMS: 9779, 9780; 5-pass DIS production runs: HMS: 3408, 3419; SHMS: 10602, 10614. For slides please see: Mingyu's RefTime slides (all runs), Junhao's RefTime sldies (SHMS 9781 and 10601 only). At the meeting we only had time to go over Melanie's slides, see: Melanie_RefTime_HMS_20200428.pdf and Melanie_RefTime_SHMS_20200428.pdf. QandA are as below:
      1. p1: Why are scalerrate extracted from fitting pTRIG1 different from the report file? - A: they should be the same. -> HW: need to understand how rates are calculated in the report file. (will need more guidance to proceed on this).
      2. p2: DCREF2 is different from others (1 through 10). Could it be double counting? (but odd). -> HW (Brad): will look at time difference between pulses, use data to study this first and later confirm on the scope.
      3. p3: bin size seems to be different for REF2,3,10 here -> HW (Melanie): need to modify script to use fool-proof method, i.e. create histograms with well-defined range and bin size and then fill, do not use T->Draw().
      4. ELclean for elastic runs seems to have some events at 0 multiplicity. Thought this is odd (in principle there shouldn't be any), but there are only 3 entries out of 200k so can ignore for now.
      5. Double peaking in Tref: Simona has found some explanation, see below and more HW for Melanie, Murchhana, Mingyu and Junhao.
    3. Murchhana's slide was focused on questions on how pedestals are treated for FADSs, see MRoy_PedQuestion.pdf. Most of answers below are from Simona:
      1. p1: what are the differences between adcPedRaw and adcPulseTimeRaw? A: adcPedRaw is not important for timing. (this is only important for energy or other calculations where pedestals change the information extracted).
      2. p2: How do we get hFADC_TREF_ROC1_adcPed from hFADC_TREF_ROC1_adcPedRaw? A: the relation depends on how many channels are summed together and the ADC dynamic range. (my note sames "sum of 4 samples" but not much else).
    4. Simona presented her study on the reference time malace_A1n_training_session_2.pdfand in particular looked into the double peaking, see discussions below
      1. First looked into whether these come from EDTM events, see slides 2-5. Decoder seems to have been disabled for EDTM (so no), but -> HW(students, perhaps one of them) should look into how the histograms look if EDTM decoding is enabled, just to confirm we are indeed not studying EDTM events now.
      2. Looked into calorimeter energy deposit and these events look like good electrons, see slide 6.
      3. Double peaking seems to come from events that have 0 value in hodoscope P1X timing. (They have the same 0 value for P1X, P2X, P1Y and P2Y). These events cause the peak on the left. See slides 8-11.
      4. Since these appear to be good events, we should choose cut in DCREF timing to be to the left of both peaks (around 4600 on slide 9).
      5. Did not have time to go over all other timing slides. Will make suggestions for where to place timing cut for other channels next week.

  • 04/14/2020 A1n/d2n Analysis
    1. Ref timing training session 1 from Simona (thanks!): see slides malace_A1n_analysis_training_1.pdf. A few additional notes:
      1. p6: since only one common cut is applied to all channels of DC, need to plot all a top of each other (as shown).
      2. p8-11 are how to define/load all calibrations, cuts, and kinematics to hcana. (the negative sign on p.11 are explained in the file comments.)
      Based on the information presented we will start from follows:
      1. Follow instructions from the slides and make plots for all ref timing (all detectors)
      2. Pick a high rate run from Jan 1-pass elastic run and make the plots. Place a cut on multiplicity (or make 2D plots) to study correlation between timing and multiplicity.
      3. Do a Gaussian fit of scaler rates (as in p.12) and plot mean value vs. run number. Add Trigger type (3/4 or ELreal or ELclean), PS value on top of plot. Maybe repeat with DC rate?
    2. Jixie has prepared a summary of optics runs, we will follow up with Mark J. to make a plan for optics calibration.
    3. Due to the APS April meeting we will not meet next week. Instead we will meet in two weeks, on 4/28. A few of us will follow up with students on ref timing work between now and then.

  • 03/31/2020 A1n/d2n Analysis
    1. A1n completed (partially) its data taking on Friday March 13th. Due to current circumstances of the world we have postponed the end-of-run party until further notice. (no fun drinking beer virtually!)
    2. Otherwise, not sure why people say the world has stopped, we are moving on just fine. We made a master plan for carrying out data analysis and tentatively assigned names. To start we will need to get the detector calibration done though, which we didn't discuss today. See A1d2mtg_notes_20200331_general.pdf for analysis flowchart (and names) and A1d2mtg_notes_20200331_target.pdf for a recap of Monday target analysis meeting (tasks and names).

  • 12/30/2019 Target field updates:
    1. During X'mas break it was also found the KEPKO power supplies have offsets in currents, Jixie generated the predicted fields for the elastic (11.7 deg HB on CC on) setting with the estimated power supply (actual) currents and here are the predictions: gradient_11.7degHB_BxBz_By_2.1483GeV_predict.pdf. The gradients are still quite small and the field direction rotated by about 4 deg. (Jianping's estimate is 5 deg between laser and field direction).

  • 12/21/2019 Target field summary:
    1. During December running of A1n the target seems to experience large AFP loss (1.7%) and masing at the same time. This summary post is aimed at listing all observed target data, results and predictions from field mapping, as well as comparisons to the Tosca model.
    2. Here are hclogs from 12/17 owl shift:
      1. AFP loss was reported in hclog 3755547, at which point the pumping laser was kept at 85-90W longitudinal. Target was at the reference cell position taking data, with SHMS at 11.7 deg and p=-2.1483 GeV/c.
      2. A few hours later masing effect was observed: hclog 3755579; Oven temp=200 degree C. Correction coils were set at I_VL=2.7 A, I_VS=2.2 A using Jixie's predicted optimized values (to minimize both By and gradients), see Jixie's Target Field predictions. His predicted fields for this particular spin/CC setting is https://userweb.jlab.org/~jixie/A1N/FieldRunPlan/1pass/gradient_11.7deg_predict_2.1483GeV.pdf. (z is pointing along beam, x is from left to right viewing downstream, and y is vertical up).
      3. After masing effect was observed, JP and Mingyu suspected the PNMR coil was left on and could cause masing, so they went into the hall and disconnected the coil. See hclog 3755696. Polarization seems to have increased initially but plateaured again.
    3. There was no TOSCA model nor direct field mapping with SHMS at 11.7 deg, but these exist for SHMS at 12.5 deg and -7.5GeV/c (A1n production resonance setting).
      1. Field mapping: For 12.5 deg -7.5GeV/c setting, fields were measured directly for VL=8.7A, VS=6.0A, see gradient_12.5deg_HB7.500GeV.pdf (pages 1 through 4 are measured target fields at 0,180,90 and 270 deg).
      2. Jixie also made predictions for fields from mapping measurement of each individual coil, and made comparison of measured vs. predicted values for the same VL=8.7A, VS=6.0A setting for 0 and 180 deg fields, as well as for VL=6.4, VS=4.2A 0 deg fields. These can be found on pages 1, 3, and page 2 of gradient_12.5deg_HB7.500geV_predict_vs_real_3pages.pdf, respectively. I personally think the predictions works for By to the 0.1-0.2G level, and for Bz to the 0.3-0.4G level. Please note that in these predicted values, the coordinates (x,y,z) have been rotated slightly due to adjustment of the target platform, and resulting in "real values" coordinates (not integers). .
      3. TOSCA model produced by Steve L at the SHMS 12.5 deg, -7.5GeV/c setting was processed by Vladimir, and results are here: talk_05222018.pdf. Pages 4-9 are without correction coils and the remaining slides are WITH correction coils. What are still missing are correction coil currents used in the TOSCA calculation. Thus, at this moment the only comparison we can make between mapping and TOSCA is 12.5 deg, -7.5 GeV/c no correction coil setting.
      4. Jixie has made predictions for 12.5 deg, -7.5 GeV/c setting without any correction coil on, see gradient_12.5deg_HB7.500geV_predict_only.pdf. The 4 pages correspond to field in 0, 180, 90 and 270 deg, respectively. He also measured the real fields with HB on no CC, but only for 90 and 270 degs (due to time constraint). His measured real fields are gradient_12.5deg_7.500GeV_real_NoCC.pdf. Finally, there is the comparison between predicted and measured fields for 90 and 180 deg, see gradient_12.5deg_7.500GeV_real_NoCC_gradient.pdf
      5. At a quick glance, it appears Jixie's predictions for HB on no CC agree fairly well with the measured fields at 90 and 270 deg. On the other hand, comparing to TOSCA model, Jixie's perdictions for 0 deg seem to agree well with page 4 of talk_05222018.pdf, but the predicted (and measured) By field is either to small or shifted compare to page 5 of talk_05222018.pdf.
    4. Over the X'mas break, spin-up and spin-down measurements of cell Brianna using longitudinal laser, 0 deg holding field, with HB and all correction coils OFF showed somewhat expected behavior, with projected saturation polarization around 40% and spin-down constant around 10hrs. Field measurements and calibration of the coil power supply also done. Jixie's prediction and mapping measurement for the helmholtz-coil-only field can be found in gradient_BxBz_0deg.pdf. In this 4-page PDF, page 1 is the predicted field using individual Bx and Bz field data, earth field background subtracted. page 2 is the measurement of main holding field, earth field background subtracted. page 3 is the difference between real measurement and prediction. page 4 is the measurement of the maim holding field including earth background.

  • 11/19/2019 Run preparation weekly meeting:
    1. Participants: Arun Tadepalli, Jianping Chen, Greg Smith, Brad Sawatzky, Jay Benesch, Melanie Ruhfuss, Mingyu Chen, Junhao Chen, Jixie Zhang, Xiaochao Zheng, William Henry, Murchhana Roy, Xinzhan Bai, Wolfgang Korsch.
    2. Greg on RC report: We will not have beam in either A or C until Dec.2nd. Will cancel shifts up to owl Dec.2nd. As for spin dance during Thanksgiving week, Thia made the call to not do it so we can finish target work.
    3. Meeting schedule: We will meet on Tuesday 11/26 our usual time. Then will meet on Friday after Thanksgiving Day for those who will attend. Starting Dec.2nd we will have daily RC meetings (time determined after meeting to be 4:15-5:15pm).
    4. A few topics Brad, Greg, JP et al need to discuss locally: 1) what pass do we start with? Can we have 2-pass beam for commissioning and then drop to 1-pass for 3 calendar days for calibration? (Jay commented doing 2-pass shouldn't be a problem); 2) which target cell do we put first?
    5. Target status: Junhao reported that he will start cleaning the laser room today with Xinzhan. The damanged long fiber is being fixed today by Raytum and will be deliered soon. For 4-1 combiner, each fiber coupler has been tested with no proble (at full power). Still need to test all 4 in the combiner at full power to make sure it works/does not deposit too much heat.
    6. Cell: no update, alkali mix is being delivered from UVa to WM today and cell "Yixin" is being filled this week.
    7. Raster: Mark wasn't at the meeting but by email and hclog it appears there are work to do and progress made on the raster. To-do: set the frequency of rasters to 7.751kHz (currently 8kHz); make sure raster coil packs A and B are identical (now they are not balanced. Note: double coils are needed to bend the 12 GeV beam, thus the raster coils of 6 GeV was copied to produce two); after these problems are solved will look into raster patterns (currently there appear to be a hole at the center).
    8. Detector/DAQ readiness (Brad): Eric is doing all the checks. Ony HMS noble gac Cherenkov is awaiting for opportunistic installation, otherwise things are on track.
    9. Target field (Murchhana): measured target field changes in Bx and Bz for opposite polarity of HLCC (compass cannot measure By). Results on the Bx and Bz field components up to z=+/-12cm are shown in CC-polarity-field-check.pdf, and on the field directions from z=-12cm to +16cm are shown in Field-direction-CC-polarity.pdf.
    10. Definition of field direction: for Murchhana's compass measurement: direction-definition.pdf. For Jixie's 3D probe direction, Bx and By are the same as Murchhana's but Bz is pointing upstream along beamline.
    11. Jixie on field calculation/predictions, see his report He3_Target_Field_Mapping_Status.pdf. We identified some immediate to do items and some problems:
      1. Discussion on background (page 8 bottom plot): It appears every time the 3D probe is turned on, there is a slight variation (below +/-0.04G) in the reading. This should be counted as uncertainties;
      2. Discussion on zeroing (page 8 top): the 3D probe was not zero-ed before the mapping. After all mapping was done, the probe was zeroed in the EEL building. Need to evaluate the correction and the uncertainty due to zeroing.
      3. First to-do is to figure out what was done for the target coil adjustment (page10) (done after Jixie's field mapping and before Murchhana's field compass measurements), is it a 3D rotation? -> then use model to make predictions and to compare with the measured values. (Jixie measured the fields before and after the adjustment).
      4. Second to-do is to make predictions of the difference between HLCC opposite polarity and compare to Murchhana's results. Focus on field directions though field components will of course be involved as the intermediate step. Predictions for fields are already shown on pages 17 and 18, need predictions for the difference in fields and the field directions.
      5. Inconsistencies between the predicted coil settings and the field compass measurement: (1) (page8) In addition to the wrong polarity of HLCC, there was one mistake on another coil; (2) When Jixie calculated the coil settings, the background values were already subtracted. This means the fields were optimized to the background values and not to zeros. Need to evaluate effect of this.
    12. We discussed briefly on the run plan: Although it is more likely we will commission the experiment and set up the beam with 2-pass, the optics can be done either all at 1-pass, or partially at 2-pass (30 deg y and angle) and then at 1-pass (delta scan at small angle of elastic setting). Mark Jones showed previous data with sieve slit with spectrometer at 2 GeV and the holes are clear (no punchthrough), but not sure what energy was used. Need to confirm 2-pass beam also does not cause punchthrough. Xiaochao will work on finalizing the run plan, starting now and possibly at JLab during Thanksgiving break (next week).
    13. Coil settings: JP asked whether we can run elastic and Delta calibrations without any correction coils. The answer at the meeting was yes but not sure about the details and/or exact field values in x and y directions -- need followup.

  • 11/12/2019 Run preparation weekly meeting:
    1. Participants: Arun Tadepalli, Jianping Chen, Dave Gaskell, Greg Smith, Brad Sawatzky, Jay Benesch, Melanie Ruhfuss, Mingyu Chen, Jixie Zhang, Xiaochao Zheng, William Henry, Murchhana Roy, Todd Averett, Vladimir Nelyubin, Wolfgang Korsch.
    2. Jay on accelerator status: there is vacuum problem that needs warmed up and repaired. Not sure how much delay but Hall A is not running during Thanksgiving week either. A few items to discuss:
      1. During Thanksgiving week will do beam polarization measurements in Halls A and B. Is it possible to also send beam to C for polarization? Dave/Brad et al: Will need two shifts at low current (~2uA) CW, 1 for setting up Moller 1 for spin dance. The two shifts will not be back-to-back but does not necessarily be during the day; will need decision made by next Wednesday (11/20) because need to coordinate with cryo on cooling down the Moller magnets -> JP, Brad, Dave and Jay will talk about this offline.
      2. JP pointed out the beamline will be the last to put on (after target is installed). Need to consider effect of sending beam through air, will it cause high radiation that will interfere with target work, etc?
      3. Due to linac vacuum problem not sure what the highest energy will be. In the worst case will be 10.23 GeV with P-sq=0.93.
    3. Todd on cell status, see slides Cell_status_12-Nov-2019.pdf. New cell Austin filled at UVa turns out to be a good (but not very good) cell. Saturation pol is 52% with lifetime about 20 hrs. (XZ)- Gordon mentioned the in-beam polarization will be in the high 40's, maybe 48-49%? Now Todd is waiting for new alkali batch from Huong. New cell "Yixin" ("Ee-Sin") is on the vacuum system at WM and vacuum looks good, now filled with N2, will be filled next week. Mike S. is waiting for Huong to ship more windows.
    4. Jianping on target/Hall status:
      1. Starting target ladder alignment work. Some small problems due to platform not stable. Need to test with real reference cell.
      2. Asked alignment group to install and leave the laser iris system in situ so we can check alignment for every cell without involving the alignment group. The laser iris system will be mounted with aluminum (but outside RF coils).
      3. Carbon target holder (Al) is now done and mounted.
      4. Carbon holes: need to adjust ourselves and may not be trivial.
      5. Optics/fiber connector: Junhao is still working on the heat dump inside the connector, it is still signficant. We discussed whether we can use UVa optics. The main problem is there are only remote control for two optical lines (mostly control of laser waveplates). If we use UVa optics we have to make access everytime we change the target spin direction. However, it can still be a backup plan in case the 4-1 connector does not work and/or burns the long fiber.
      6. Another problem with remote control is the target field coil power supply cannot be put into negative polarity. These are the new PS. Trying to use the older PS unites (about 20 yrs old) which can go to negative.
      7. The problem with not being remotely control the target spin direction is not only limited to the fact that we need to rotate the target spin for data taking, but also that target polarimetry cannot be done in the transverse direction. Thus when we run transverse target the spin needs to be rotated to longitudinal everytime a polarimetry measurement is done.
      8. CANS/laser interlock system: moving forward okay
      9. Laser OSP: An inspection of the laser system will be done this Friday. If passed, can send OSP for signatures.
      10. DOE safety review: a team from "office of enforcement" will be on site on Tuesday-Thursday next week. Wednesday will be on our target system. JLab people already came to target lab yesterday and today to review the status and prepare us, but more preparation needs to be done for next week's DOE review, in particulat the students (Mingyu and Junhao) who will be interviewed.
    5. Raster status: (Brad/Jixie) coils are installed in the hall. Raster can only be controlled from MCC now. Bill G., Eric P. are working with two people from MCC to test the raster and this work is ongoing. An hclog entry 3740501 was made and found two problems: (1) The current readback is not the set value; and (2) the raster appears to be square (not sure if it is an interpretation problem or if it is actually square).
    6. Jianping follows up on the field direction measurement discussed last week: Suggests running at the settings measured by Murchhaha. The vertical field is up to 2 G at this (wrong) setting, but given Ay is 1 order of mag smaller than A_perp which is smaller by a further order of mag than A_par, we can run with nonzero By and use Ay measured from transversity to correct our results. what's important is the uncertainty on Ay which Jixie can get down to 0.2G. Wolfgang will look into Ay results from 6 GeV, including kinematic coverage and how to extrapolate to our conditions.
    7. Murchhana on field direction results: Showed preliminary results on the longitudinal field (no uncertainty yet), see Longitudinal_field_direction.pdf.
    8. Xiaochao will follow up with Jixie and Wolfgang to look over Jixie's field mapping analysis and results. Brad would like to see a report on the results by next Tuesday. Given the time needed to look over data, first thing to report on is all field gradient results and whether they are tolerable by our target. Second is to figure out if Jixie's probe has an offset from the real coordinate (which will affect calcuations). Third is to calculate coil current settings for all running conditions that are not measured by Murchhana (will take longer time and may not be ready within a week).

  • 11/05/2019 Run preparation weekly meeting:
    1. Participants: Gordon Cates, Greg Smith, Brad Sawatzky, Jay Benesch, Jixie Zhang, Xiaochao Zheng, William Henry, Murchhana Roy, Todd Averett, Vladimir Nelyubin, Arun Tadepalli.
    2. Jianping was not at the meeting but he reported target installation status by email. A shortened (but still long) version is as follows:
      1. The oven is mounted and survey/aligned;
      2. the new aluminum carbon foil holder is being machined and should be ready early next week (probably Monday);
      3. target ladder has a few last minute issues/modifications: in addition to the carbon foil holder, added vertical angle adjustments, tested out two mounting plates which are the bottom plate for the oven as well. After mounting on the oven, it was discovered that it's not rigid enough that the bottom side would move by ~20 mils when touched by a small force. An reinforcement was to be added (should already been done on Friday).
      4. Jack's group has been working on the target lifter. They informed us on Thursday evening that there was an issue with the limit switches that the motion system will not be ready for Friday and Monday (could be Tuesday as well). The initially scheduled alignment/survey of the target ladder was postponed to at least Tuesday; The goal is to complete the target ladder alignment/survey this coming week. The final check with actual target cell and reference cell will be done with a laser/iris system.
      5. Discussed with Walter on the work for Hall C tech in the next few weeks: the optics box is mostly there; Walter/Todd have started to work on reference cell/cooling jets system and will continue next week; Will start work on the over gas system. Will start fitting/testing out the "scattering chamber" cover pieces; the nitrogen/helium "hood" for cooling gas pumping out line which connects to the vent line has not detail design yet. Will have a meeting with Bert/Walter on Monday to finalize this;
      6. We had problem with fiber connection. Raytum is helping us. They tested our fiber connector: it worked with both fibers have no coatings, when tried with short-nocoating fiber to short-withcoating fiber, it also worked. But it failed when tested with short-nocoating to long-with coating. They are continuing tests and will also repair the long fiber
      7. Mark Taylor is making the laser interlock box and starting the rest of the interlock system, including the new one on the Hall-C to BSY door. Will helo with interlock on smoking detector.
      8. In addition to help with installations, Junhao is testing the polarization compensation, Mingyu is setting up and performing pressure broadening and thickness n=measurements, Murchhana is analyzing the Compass measurement results, Jixie is analyzing the field mapping results.
      9. For control system, I will leave this to Brad since he and Arun/Junhao are working with Jack's people and EPICS group on this.
    3. Todd and Gordon on cell status, see Cell_status_20191105.pdf. Cell Austin is planned to be filled on Wednesday 11/6. YiXin will be filled at WM but awaiting for new alkali mixture from UVa. UVa would like to test Austin performance first to make sure the alkali quality is good, so earliest time to fill Yixin would be early next week. Also discussed details of the cell production procedure.
    4. Todd wil start the reference cell gas system this week, including cooling jets. Walter will order (or already ordered?) all gasses needed.
    5. Brad on scheduling: at the Hall C meeting yesterday JP reported the target installation is about one week behind schedule. Based on this information, Thia et al. moved Hall A running up by one week and we can assume Hall C will not be running during the Thanksgiving week, i.e. the start time for Hall C is now Dec. 2nd. We will wait until a confirmation from Thia before officially canceling shifts for the first week. We will also ensure that the management knows 1-pass calibration is needed before production. Whether we will use 2-pass to commission the beam/hall alone is still to be discussed.
    6. Jay on accelerator status: "accelerator is unfortunately also blown", at least 1 week behind schedule but starting Nov. 25 is still quite possible (or Nov.22nd for Hall A). South linac has come back. The problem is in the north linac where modules were warmed up and moved around. Will know more by Nov.11th.
    7. Murchhana on field direction measurement, see her report Field-direction2.pdf on the analysis of transverse field directions. (need more time to analyze longitudinal field results and will report next time). A short summary with concerns are below:
      1. At 90 deg setting, the B field direction changes by up to 0.3 deg at z=+/-12.5cm compare to z=0 for nearly all settings. Question: how much the direction deviates beyond +/-12.5cm? Jixie's field mapping/gradient results may give us an estimate.
      2. The uncertainty of the measurement is a combination of the mirror alignment and survey results.
      3. For 3 settings (see slide 11), the HLCC current had wrong polarity and current -1A (opposite of Jixie's calculated +1A) by mistake.
        1. Question 1 here is what setting we should run at? (Jixie) The wrong HLCC setting will cause a ~0.9G field in the vertical direction, at the meeting everyone seems to agree with running at Jixie's, i.e. calculated setting. (Brad: we need By or vertical field bound!). But see meeting minutes on 11/12 above, JP claimed we should run at the measured setting and use transversity data to correct the measured asymmetry.
        2. The second question is then whether we need to repeat the direction measurement with the correct HLCC current setting. The only chance to do so is during target rotation but we budgeted only one week for the rotation. The field direction measurement, since needing survey, will require 3 days minimum.
        3. We decided we need Jixie and Murchhana to complete the analysis and provide an estimate of the field value along with uncertainties if we run at the correct HLCC setting. If the uncertainty is small we will not need to repeat the field measurement. But if it is large then we do.
        4. Xiaochao pointed out the three lines with the wrong HLCC polarity are all for small angle running. We will need to combine the uncertainty due to the field direction being measured with wrong LHCC current, and that due to target rotation.
        5. post-meeting note: We can use a simple field probe to measure the vertical field with HLCC at +1A and -1A and use it as a bound.
      4. In fact, even for cases where all coil settings are correct, we need to consider whether we need to repeat the direction measurement after the target rotation.
    8. Jixie on field mapping analysis, see his report He3_Target_Field_Mapping_Status.pdf. The slides have a lot of details but a short summary with concerns is below:
      1. mapping frame is not surveyed, Jixie will fit data to Tosca model and see if can figure out the probe alignment. Question 1 here is how good we can trust the Tosca? Question 2 is whether a small angle of the probe will be problematic. XZ's quick answer is probably not because (small) probe misalignment angles will cause only second-order effects on the gradient results.
      2. To setup the frame, a steel bubble level was used. Jixie ensured everyone the level was removed during the mapping.
      3. Jixie's caculation for all correction coil (CC) currents is based on optimization of the target chamber (not the pNMR bulb);
      4. Found the background (Earth field) to vary by up to 0.1G in certain directions. This will add to uncertainties. Need to know the uncertainty for each direction and evaluate. (For example, uncertainty in Bz may not matter much for A1n but will matter for d2n). WK commented the change in the background may be due to the order at which the coils are turned on (and the de-Gauss effect).
      5. slide 8: Bx,By,Bz defined by probe direction, Bx is vertical and By is horizontal, and not the real x,y,z direction we commonly use. On the same slide, x,y,z are pinhole index on the frame.
      6. Overall a lot of data need analyzed and predictions for CC current settings made for all kinematic settings. Jixie will continue working on this.
      7. Jixie pointed out the actual measured field differs from his calculated values by 0.1A (roughly 0.15G). This is the same level as the background uncertainty and we may have to treat it as an uncertainty overall. Need numerical values and directions to evaluate impact on physics.
      8. In addition, Jixie is not sure how well calculations work for after target rotation. Do we need repeat both field direction and field mapping during target rotation? Note: this can also be done after the run, though if the setting is very off then the fields during running would be affected.
    9. Brad: need a few target experts to be Radcon-II trained. Jixie mentioned it was already decided, Arun, Murchhana, Junhao, Mingyu and Jianping are the only five need this training.
    10. Post meeting note: RSAD is ready (Brad); Xiaochao finished COO and sent it to Steve W. (he will forward it for posting); Steve is working on the ESAD based on previous polarized 3He experiments' ESAD (GEn and transversity).

  • 10/29/2019 Run preparation bi-weekly meeting:
    1. Participants: Brad Sawatzky, Jianping Chen, Jay Benesch, Jixie Zhang, Xiaochao Zheng, Syvelster Joosten, William Henry, Murchhana Roy, Steve Wood, Todd Averett, Melanie Ruhfuss.
    2. Summary of all trainings needed, for JLab online training see http://www.jlab.org/div_dept/train/webbasedtraining.html
      1. First you must have a JLab user account, see User Liason website for how to register as a user and the trainings required.
      2. Prior to visiting JLab you must register your visit (regardless of whether you are a new or seasoned user, and must be done in advance by at least 7 days): go to JLab access portal https://misportal.jlab.org/jlabAccess/ and register your visit in advance. You will not be allowed through the accelerator gate without a valid visit registration. Also we suggest registering for a year-long visit whenever possible.
      3. Trainings required to take shifts, if not already done as part of user registration:
        • ESH Orientation (SAF 100)
        • Radiation Worker Training (SAF 801), including SAF801C (online class), SAF801T (test, must be done on-site), and SAF801P (Practical, must be done on-site with RadCon, will need registration/appointment in advance). Note you will need to take SAF801P if you let your SAF801 expire too long;
        • Oxygen Deficiency Hazard Training (SAF103, online);
        • Hall C Safety Awareness Walk-Through (SAF112, online and on-site, contact Joe Beaufait beaufai@jlab.org to set this up, typical time is 2pm on Tuesdays).
        • Conduct of Operations (SAF120) -- reading of experimental COO, can be done in the counting house prior to shift taking.
        • Target operator training: arranged by Jianping Chen. I will post details once I hear from him.
      4. Specific for working on the target pivot and with lasers:
        • Ladder Safety SAF307 (online, ~20min);
        • Fall protection (site specific): SAF202B (online, ~20min);
        • Lock/Tag: SAF104 (online)
        • General laser (new DOE training): offered online by DOE's National Training Center (NTC), Instructions below:
          1. Go to the NTC webpage (https://eaccess.ntc.doe.gov/)
          2. Enter your previously issued username and password (issued to anyone here in October 2017 for mandatory National Security Training), retrieve that information if you did not record it in 2017, or create a new account
          3. Once logged in go to "Application Access" and select "NTC LMS"
          4. Select "Course Catalog" then "ELD-Enterprise Learning and Development" then "ESH-Environment, Safety, and Health"
          5. Select ESH-333DE, DOE Laser Worker Training and complete the course
          6. Important: Print out the course completion certificate (Take a screen shot!) and send a copy ( note the online software seems to have problem saving records. Without a screenshot there may not be any proof that you have taken the training and passed the test!) to Tara Tyndall (wilkerso@jlab.org)or Tish Creery (creery@jlab.org) in HR.
          7. Onsite: Target OSP and Laser SOP (LSOP) -- need to be done at JLab locally;
      5. Once we start beam, radiation worker 2 training SAF802C is also recommended in case of working on or around the target pivot.
    3. Xiaochao reported for Todd and Gordon on cell production status. We have five good cells now, one new cell "Columbus" filled and awaiting test, and two more cells "Austin" (at UVa) and "NonameYet"(Mike will make it this week and ship to WM) to be filled, see Cell_status_20191029.pdf
    4. We discussed beam current limit on 3He cell. What if we need 33uA instead of 30uA to optimize FOM? Anser is Ops currently hae a 30uA restriction on 3He cell. If slightly higher (by 10 percent) curents are needed we will nee to make minor update to the paperwork and Ops instruction, but should not a big process.
    5. Hall status/ target installation (JP):
      1. Field mapping and direction measurement (compass work) are done. Finished all survey too. Now alignment group is working on controls. After that will proceed to oven, target ladder and motion system installation (this week). Optical fiber install will be next week (Joe B. will get the fibers and cables ready by this Friday).
      2. Carbon foil holder: The actual material is "ultem 1000" and not the "vespel SP-1" we have simulated. According to Silviu, "vespel and ultem have about the same density and thermal conductivity coefficient, so the simulations results would not change much if I were to change vespel with ultem, the difference is the max working temperature for the two plastics respectively, vespel can take almost 100 K more than ultem; vespel is also more expensive". Also see summary of Silviu's simulation below. It appears "ultem 1000" will melt at 30uA for sure and we need new frame made. Jianping said we will probably use aluminum which has been used before.
      3. Once target ladder is in place, need to check position of the actual cell. May develop a laser/irises system for this.
      4. Fiber/laser status: Raytum is helping to test the fiber combiner (previously suspected the fiber-end coating may cause the fiber to turn). Steve/Raytum will be here on Friday 11/1 to bring a new laser and show how the fiber works. Long fibers: We have 9 so far with 1 broken/burned. Need 8 for the experiment. Will ask Raytum if the burned fiber can be cut/repaired.
      5. CANS system: most problematic as plant engineering had no time to start this work. Now has asked for help from management to place this task (install lock on the hall entrance to/from accelerator tunnel) at a high priority.
    6. Todd: met with Jianping and Walter earlier today and will start working on the reference cell system in the hall.
    7. Melanie on shift help page and checklist, please see checklist instructions at https://hallcweb.jlab.org/wiki/index.php/Shift_Checklist_Help. Will add parity varialbes once we know what we need on the checklist (we probably do not need the full list of Hall A parity checklist, but will need to make sure those not in EPICS or critical variables such as "Hall C IA cell setting" are in our checklist). -- looking good!
    8. Raster (by email): Bill G. will install raster in the hall this week (as planned). Since Mark is away until 11/13, Jixie will check the raster pattern after install and tune the uniformity. Jixie mentioned Pengjia's experience is that changing the x vs. y phase solved the uniformity problem for g2p. Fingers crossed this will work for us!
    9. Xiaochao will work with Steve W. on finalizing the COO and the ESAD. -- followup: COO is done; Steve will work on ESAD, need to add section for the polarized 3He target.
    10. Shifts: Xiaochao will send emails to a1n_d2n and hallc_running lists.
    11. Post-meeting: Jixie will work on putting the runplan into wiki.
    12. Post-meeting: Thermal simulation by Silviu:
      1. On carbon target and holder, see email dated 2019/10/24, summary PDF OpticsTmaxTable_24oct2019.pdf and figures under 20191029/thermal_sim/. In conclusion "For vespel SP-1 (JP please confirm the material we are using) as optics frame material, the max operational temperature, from literature, is 533 K. So, 30 uA beam current rastered or not should be fine on all optics positions, for the O4H (optics 4 high position), the beam current limit could be set to 50 uA, rastered or not."
      2. However if carbon foil holder is "ultem 1000", which can take 100K less than "vespel SP-1", then we can't put 30uA on the carbon foils (frame will melt).
      3. on Be window, see email dated 2019/10/30 and figures under 20191029/thermal_sim_be/. "In conclusion: it seems that the Be windows should be able to take even up to 60 uA unrastered beam with an intrinsic diameter of at least 200 microns." However, only beam heating is simulated here. There can be other factors that require cooling or special care (such as stress and exposure to oxygen). Work is still on going on this topic.

  • 10/15/2019 Run preparation bi-weekly meeting:
    1. Participants: Brad Sawatzky, Jianping Chen, Jay Benesch, Junhao Chen, Jixie Zhang, Xiaochao Zheng, Syvelster Joosten, William Henry, Murchhana Roy, Steve Wood, Todd Averett.
    2. Xiaochao reported for the WM Todd's and UVA Gordon's team on cell production status, see Cell_status_20191015.pdf. Quick summary: Zhou will be a backup cell (40%), new cell BigBrother at UVa and Columbus at WM are ready to be filled.
    3. Jianping on target installation/hall status:
      1. students finished wiring correction coils. Mounted frames for field mapping. Some control for mapping still need to be worked on.
      2. This week, can not access the hall between 3pm and 1am due to safety system tests. (Jay commented) this is for PSS certification that is required by DOE every 6-8 months. The test will end by Friday afternoon for sure.
      3. Expect next week to be field mapping (plan already developed);
      4. The week afterwards (week of Oct.20) will be field direction measurement with the survey group.
      5. There are still quite a lot to do on control and cabling and will be a concern.
      6. Do not plan to use UVa 1-inch optical modules. Huong and Vladimir will come to help (Xiaochao) but need to make sure trainings are completed. This include: (online) ladder, fall protection, lock/tag, general laser (new DOE); and (local) OSP and LOSP, and radiation worker II once the target sees beam.
    4. Todd reported on reference cell, two reference cells are ready. Mingyu is at WM to measure wall thicknesses of Columbus.
    5. ON cooling gas topic (JP): JLab is required to minimize use of 4He so we should plan on using N2 cooling:
      1. Silviu's thermal simulations show 10 l/min of helium or 20 l/min of N2 flow have the same cooling effect and is sufficient (cell temperature in the 600K range), see 3he_jetcooling_case1_sep2019.jpg for helium cooling (note horizontal axis is total flow), and Silviu's email dated 2019/9/30 for N2 ("I did a cursory check considering the cooling agent to be N2, 2 jets, 10 liters/min flow, Tmax ~ 647 K. (In the same conditions with 4He, Tmax is 611 K"). If anyone is wondering the stress from the 600K beam nipple temperature, here is our amendment tech note on the mechanical stress and how it's not a problem given our glass cell windows have been tested up to 20atm: 3He_ERR_Reply_to_Final_Report-Amend-3Aug2019.pdf.
      2. To make sure Silviu's thermal simulation results are reliable, we asked him to simulate 6 GeV, 15uA no raster case (since cells were blown under such conditions). Here are his results: and email dated 2019/10/5 which found the beam nipple temperature reaches 1339.3 K or 1066 C. Given the softening point of GE180 is known to be 1015C, see UVa's glass_properties.html, the simulation shows indeed a raster-off beam would blow the cell by melting the glass.
      3. Also 6 GeV GEn used N2 cooling.
      4. But we need to consider two issues before making a definite answer:
      5. Safety considerations: N2 will have more multiple scattering than 4He and we need to make sure the beam size is not too large on the downstream window. (this problem happened previously and blew the window - need confirm). Need simple calculation for this effect -- XZ to do. If multiple-scatterings are significant we can use 3mm-dia raster size for 1-pass commissioning. Details of calculation: assuming 10cm of N2 in front of cell at 1 atm. (Jay) beam raster angle at pivot, at 5mm square size, is 220 micro-radian half angle. Followup by Xiaochao: spreadsheet on multiple-scattering calculations for 2.05 GeV beam (two sheets are based on two sets of formula of M.S.) Multiple_scattering_1.xlsx. Overall the total extra spreading in radius on the target downstream glass window is 0.01mm or 10-percents larger than 4He (4He is 0.10mm or 0.12mm in radius, or 250-300microrad in half-angle, depending on formula), and is considered not a problem.
      6. Physics consideration: one effect is the two-step process which was a big problem for SAGDH. But we are at higher Ebeam and larger scattering angle so (JP) is reasonably confident this is probably not a problem for A1n/d2n. A simulation would be nice but we may not have the time.
    6. Mark by email: Chris C. expects to have 4 coils ready by Nov. 1. The hardware is ready except the Aligent generator needs to be put on the network. William Lu is working on the software upgrade. -- need email Mark to followup on reliability of plan and time estimated for raster to work.
    7. Syvelster: software meeting is every wednesday now. The goal has not changed: is to have a full set of working scripts for detector calibration/checkout/monitoring by end of month. Higher level analysis scripts will come after.
    8. Melanie by email: Working on help page for the shift checklist, see https://hallcweb.jlab.org/wiki/index.php/Shift_Checklist_Help, and will be adding instructions as Brad adds new checklist items (like parity variables). Has also been adding relevant documents (cell production archive, collaboration meeting minutes) here: https://hallcweb.jlab.org/wiki/index.php/Polarized_3He_target_in_Hall_C, will be adding Murchanna's field direction documents too. -- (XZ) need follow up on shift worker instructions etc.

  • 10/01/2019 Run preparation bi-weekly meeting:
    1. Participants: Brad Sawatzky, Greg Smith, Jianping Chen, Jay Benesch, John Matter, Junhao Chen, Mingyu Chen, Jixie Zhang, Xiaochao Zheng, Syvelster Joosten, Vladimir Nelyubin, Melanie Rehfuss, Murchhana Roy, Steve Wood, Todd Averett.
    2. Jianping on target installation status:
      1. target is now on pivot, Helmholtz coils mounted and aligned, laser platform not stable and is being re-inforced, still need better stablization; Walter's part mostly done and now need more interaction between students and techs. The bottom plate (independent of coils, still stay in position throughout the run) is being aligned too.
      2. Next is to determine if should mount RF+corrections coils, or oven, and in what order? Need field mapping but the previous mapping frame is missing. Design exists and need remaking the frame; Oven may need to stay out when mapping (depends on how it interfere with the mapping setup). Jianping asked if we can do full mapping for only a handful of settings. For A1n it's best to do the worst setting (SHMS highest momentum), and the highest x setting. In Sylvester's magnet setting spreadsheet, the two settings are SHMS setting D concurrently with HMS settings 1 and 2; the highest x setting is SHMS settings B concurrently with HMS setting 4. Brad will need to choose two settings to do for d2n.
      3. After the field mapping will need to mount oven (if not done by then), target ladder, and survey.
      4. There are many complications due to the recent laser incident: (1) The Target OSP and Laser OSP need redone; (2) there is a new safty officer at JLab; and (3) there will be a DOE safty review in November. All these may cause delay.
    3. Todd reported on reference cell, see his slides
    4. Xiaochao reported for the WM Todd's and UVA Gordon's team on cell production status, see Cell_status_20191001.pdf. Cell meeting happened today at 10:30am.
    5. Syvelster: software "kickoff" meeting happened last wednesday. Will have regular meetings at 1:30pm every Wednesday from now on. The goal is to have a full set of working scripts for detector calibration/checkout/monitoring by end of month. Higher level analysis scripts will come after.
    6. Steve W.: mentioned on helicity decoder, when rate is low (below 350Hz) the code is having difficulty picking up helicity information. Xiaochao mentioned high x is very low rate. Brad assured us the problem has a backup solution existing. Steve will continue working on this to make sure we have helicity for low rate running.
    7. Jay: reported not sure yet what momentum we will have for the beam, will be between 10.45 and 10.56 GeV. Jay is pushing for the lower end because it will ensure both A and C have P-sq equals to 1.0, to 3 significiant figures. Higher momentum will still have P squared above 0.95, which is okay but not ideal.

  • 09/17/2019 Run preparation bi-weekly meeting:
    1. Participants: Arun Tadepalli, Brad Sawatzky, Jianping Chen, Junhao Chen, Mingyu Chen, Jixie Zhang, Xiaochao Zheng, Mark Jones, Syvelster Joosten, Vladimir Nelyubin, Jay Benesch, Melanie Rehfuss, Murchhana Roy, Steve Wood, William Henry.
    2. Jianping reported: talked to Walter, target install is a little behind the schedule but will have time to catch up.
    3. Xiaochao: reported the 9/3 meeting with Thia on cell status check, and for both WM Todd's and UVa Gordon's group on cell testing and production, see Cell_status_20190917.pdf.
      1. 9/3 meeting went okay but need to keep current cell production going at the same rate/success probability, will meet end of September to re-assess;
      2. After our own A1n/d2n meeting 3 weeks ago: Wayne turns out short lifetime. Latest cell is Tommy, with preliminary spin-up tests showing above 50% but later tests are affected by large AFP loss. Will confirm performance once the high polarization is reproduced.
      3. Savior found to be a sub-par backup cell, with saturation polarization at 40% (below Flurence)
      4. Update on new cells/retesting/testing (post-meeting amendment from Gordon): Producing and testing the new sphere is now on hold; Brianna hasn't been re-tested yet, but latest data shows it may be capable of producing (45-47)% in-beam polarization; One new cell should be available by end of this week and should go to WM, because UVa's gas system is oversubscribed; Mike may get another cell after this one and that will come to UVa. We are on track for making 3 cells (incl Tommy) by end of September as requested by Thia.
    4. We discussed shift sign-up and RC schedule.
      1. Brad will email IT to get the page setup as soon as the official beam schedule is posted (Jay said it's at the director's office for approval, how long will that take?). Shift sign-up adm will be Brad, Xiaochao, and Jixie (more may be added later). Here is a google doc for the schedule after the discussion today: google doc A1n/d2n 2019-2020 run calendar
      2. We need 20 RC's. So far only about 10 names. We need to invite more people -- XZ will email potential people
      3. First RC will be for beam polarization setup, no shift needed.
      4. Shift requirements: 18 shift "points". The weighting is: Weekday day shifts: 1 point; Weekday swing and Weekend day and swing shifts: 1.5 points; All owl shifts: 2.0 points; Thanksgiving day: 2 points for each of owl, day, and swing. Thesis students, RCs, and spokespeople are not subject to this requirement (for now).
      5. Expert shifts will be unofficial.
    5. Sylvester on software/analysis training: Will be at JLab next week. Kick-off meeting will be Wed 9/25 at 1pm with students. Then will have weekly meetings to keep pace. Higher-level analysis software will be developed after initial software training, detector calibration tools, online checkings, etc.
    6. Murchhana reported starting with Dave G. on Moller work, with Monte-carlo simulation for now; compass work is ongoing.
    7. Jianping brought up the importance of planning for the target field mapping measurement. Ask if Jixie can help (since he has experience with g2p) and answer is yes; Will work with Junhao and Murchhana on this;
    8. Mark Jones on raster: Bill Gunning needs to connect all cables. Once all are connected can power up and do fine-tuning of the raster pattern. Will followup with Bill on whether this can be expedited. Once pattern is obtained, Mingyu (with Jixie) can help to fine-tune.
    9. Steve Wood on helicity reporting and decoding, working with Melanie:
      1. For the past year has been recording helicity signal and have helicity-gated scalers. Has seen (successfully) asymmetry with (e,e'K) data.
      2. Melanie has been working on helicity scalers and data consistent with asymmetry Steve saw from (e,e'K). Will get charge asymmetry next.
      3. Current software works for 30 Hz. CREX may decide to run with 120Hz. Looks like our software decoder will work for that, and need to get it ready. (If CREX ran at 240Hz -- which is not in the current discussion -- then our decoder will definitely not work...)
    10. Jay mentioned CREX may want to add a non-beam HWP so the beam line condition on the laser table is similar between HWP in/out. We need to have the non-BHWP state in our data stream.
    11. Jay reported briefly on cryo work at the accelerator. I (XZ) only noted "shuffling 5 cryo modules..." and "sounds like everything will work once put back together..." (Should the last sentence come with a grin?)
    12. Our next meeting will be in two weeks, on Tuesday 10/1/2019.

  • 08/27/2019 Run preparation bi-weekly meeting:
    1. Participants: Junhao Chen, Mingyu Chen, Jixie Zhang, Xiaochao Zheng, Syvelster Joosten, Vladimir Nelyubin, Jay Benesch, Melanie Rehfuss, Whit Armstrong, Wolfgang Korsch, Murchhana Roy.
    2. Junhao: target has been moved to the machine shop. Not sure what the progress is there.
    3. Xiaochao: reported for both JLab target group and UVa Gordon's group: thanks to the hard work of both we have made some good progress on cell testing and making, see Cell_status_20190827.pdf. A much more involved report on the Fulla test was sent by Junhao by email, see Fulla_Characterization_08_2019.pdf.
      1. Fulla was measured to have Pmax=54% at JLab, after some major change to the optical setup;
      2. There is a remaining question about Savior: Jlab measured cold lifetime at 28 hours but the test at UVa last weekend measured 14 hours.
      3. The new cell Dutch is found to be a good one, preliminary spin up place it at 52% in beam. AFP loss study is still on-going (hence the preliminary).
      4. Another new cell Wayne will be filled later this week, most likely on Saturday.
      5. We will meet Thia on Tuesday 9/3 to discuss target status.
    4. We asked Murchhana to learn Moller analysis and she gracefully agreed. She will be moving to JLab later this week (lot of work on the compass to finish at UKy).
    5. Junhao and Whit/Syvelster discussed the target control. Need: (1) A linux PC (Junhao will ask Brad and Arun for help); and (2) laser control manual -- Junhao will contact Raytum on this.
    6. Junhao reported working on target slow control and moving. Mingyu volunteered for new work, hence see below.
    7. Raster: Mark was not present. Jixie and Mingyu will contact Mark and see if can help out.
    8. No meeting has been scheduled for students' analyzer training due to Brad being on vacation. He will return later this week. Jixie can help out with training locally if needed. Meeting/training should happen next week.
    9. Optics calibration for angle: Jixie did x-section and rate calculation for angle calibration using 1- and 2-pass beams. But (1) the low Eprime setting for SHMS at 1-pass is a concern, will ask Mark about this. (2) May want to use higher Eprime for 2-pass plan, otherwise it's not competative compare with 1-pass. (3) Will need to make sure the coverage is sufficient if using 1-pass beam. Note that we previously planned to use 2-pass for this calibration, but words on beam scheduling does not favor multiple pass changes during the early running time.

  • 08/13/2019 Run preparation bi-weekly meeting:
    1. Participants: Junhao Chen, Jixie Zhang, Xiaochao Zheng, Syvelster Joosten, Todd Averett, Vladimir Nelyubin, Jay Benesch, Mark Jones, Melanie Rehfuss, Whit Armstrong, Arun Tadepalli.
    2. Testing in the target lab is still ongoing (that's the topic of another series of meetings). Mark Jones (and Brad) mentioned the techs are busy taking down the target scattering chamber, the related beam line section, and clean up the platform. Moving the target setup out of the polarized 3He lab can wait!
    3. Analysis training: No update yet except sending training documents to students. Given Brad and Syvelster will both be on short vacation this and next week, and the target lab status, we will start analysis training meeting after the target is being moved into the hall.
    4. Dave Gaskell wasn't present but he sent in 3 hclog entries: 3715247, 3715142, and 3715239. Looks like Moller test was a success and even measured Pbeam to be 33%!
    5. We sent a technote on cell thermal stress to the ERR committee. Along with the recent successful testing of Moller, the ERR committee is happy with our reply but final "run approval" is delayed until September 1st, depending on cell production and test status as we discussed at the 7/24 collab meeting.
    6. Brad will followup on RSAD. Both Xiaochao and Brad will dust off the COO and ESAD and make them ready.
    7. Mark Jones: will meet Chris next week to get the raster work started.
    8. Field gradient calculation: at the 7/24 meeting Gordon expressed he would like to run the gradient calculation for the highest x point setting to make sure it is okay. Syvelster calculated the magnet settings, which can be found at A1n/d2n Magnet Currents Google Doc. We should check under the A1n list, SHMS #B and HMS #4, both at 30 deg. Will send the doc to Vladimir and Steve Lassiter to get the work started.
    9. Junhao and Whit discussed laser control and will continue the discussion offline.
    10. Not discussed at the meeting but for the record: the proposed shift checklist can be found at Google doc Proposed checklist items for A1n/d2n 2019-2020 running. We should semi-finalize this and update the checklist and the instructions.

  • 07/30/2019 Run preparation bi-weekly meeting:
    1. Participants: Wolfgang Korsch, Arun Tadepalli, Mingyu Chen, Junhao Chen, Melanie Rehfuss, Xiaochao Zheng, Todd Averett, Jianping Chen, Murchhana Roy
    2. Manpower discussion: postdoc/senior people: Arun (working on target control with Brad), Jixie (UVa, optics and general analysis), William Henry (Hall C, working on Moller with DG), Syvelster (ANL, leading Hall C analyzer/software work); Possible Temple new postdoc (?)
    3. On Moller (followup with Dave): DG+WH will be the two onsite people for the Moller measurement, but still need one person from collaboration to learn closely how to do the measurement and how to extract polarization online.
    4. Brad discussed software training status:
      1. Please go to June 2018 Joint Hall A/C analysis workshop, download the virtual machine and go through video tutorials. A more detailed instruction by Brad can be found at software_instru_starting.
      2. Need a clear timeline for what fluency students should gain for the next few months.
      3. Will organize weekly software meetings, time/day TBD.
    5. Discussion on ERR status: Jianping met with Chris K. yesterday. ERR cmt willing to close out but one issue remain on the window pressure test: We stated we would test windows up to 25atm (based on 30uA, 5-dia circular raster, and 2 jets at 10 slpm) but actually only testing up to 20.3(300psi). Possible solutions: (1) find GE180 mechanical strength data (probably futile); (2) Get the latest thermal simulation and stress analysis (Xiaochao will email Silviu and Steve L. on this). Note Silviu's simulation for 2 jets at 20slpm each and 1jet at 20slpm both show lower temperature at beam center; (3) changing raster pattern to have a softer edge - not sure if it can be done, but the stress is at the edge for sure; (4) Will thicker glass windows help? (Gordon suggested 175um); -- but this requires simulation and not trivial to actually change the window thickness, so will be the last thing to do; (5) Additional comments: Todd added that in the 20-atm test only about 60% of windows survive (after the meeting Gordon added it's actually only 50%, which is the main reason we are not testing them to higher pressure).
    6. Need a google doc to list all things needed on the shift checklist (target, parity) -- Xiaochao, Brad, Jianping
    7. Wolfgang will call a meeting with Brad, Bert to discuss target field compass test needs, including surveys.
    8. Need to start a wiki page at some point. "At some point" turns out to be one hour later, here it is: https://hallcweb.jlab.org/wiki/index.php/A1nd2n. But any wiki editor should first read this: wiki_instr_general. If unsure, please consult with Brad on whether we should have our own A1n/d2n page on a certain topic, or use existing links.
    9. Need a list of survey -- Brad.
    10. Prosponed to next meeting: (1) Need followup on target field gradient calculation -- need Gordon; (2) Need followup on ref cell pressure needed for run plan -- need JP, Todd

  • 07/11/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. Mingyu: See reports A1N_runplan_draft_19_07_11.pdf, Delta Resonance beam time.ods, and Reference Cell elastic and DIS_19_07_11_12amg.ods.
      1. The main progress is calculation for the Delta transverse asymmetry measurement assuming we need a 3-sigma separation from zero. Estimation of the Delta asymmetry is done following the prescription in XZ's thesis, i.e., inter/extrapolation of previous data, see thesis_A1n_ssp2_Delta3pager.pdf.
    3. Jixie showed some new Hall C data - run 8437 - taken on the carbon target last week with 2.77GeV beam on SHMS, with sieve in. If cutting on only one sieve hole, see carbon cut on 5.2 deg, and cut on 4.5 deg. The ground and the two excited states are clearly visible and separable at 5.2 deg. At 4.5 deg not quite so (excited states are very low and merge with the g.s. peak.
    4. To do's:
      1. It will be useful to know QE asymmetry for both elastic and Delta settings, so online we know what is the actual asymmetry we expect to measure based on how much QE contamination is there. XZ will look into her PWIA code.
      2. JZ: continue working on learning how to analyze Hall C data and optics data run plan: (1) need a prioritized optics calibration list, with a focus on Vz calibration; (2) include different versions of Delta scan (scan density/step vs.beam time); (3) will work with Mark Jones to learn (but should ask M. Jones to be the main point of contact for optics calibration during the run).
      3. Mingyu and Jixie will work together on a run plan document (Latex or WORD) that put things together and include enough information for online data interpretation. For example, runs cannot be shorter than ~2 minutes, need including time for configuration changes, elastic runs can be shorter, and delta can be longer, etc... Everything should fit into 40 cal hours (minus 16 hours of Moller leaves only 24 hours of 1-pass running)
      4. Need to get ready for a run plan presentation for the 24th collab meeting.

  • 06/28/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. Followup on last week's to do list: Now N2 ref cell calculations are done at the 3He elastic setting.
    3. Mingyu: See reports Beam time Estimation for Elastic and Ref Cell part II.pdf, A1NRates_results_19_06_27_larger_bins_12amg.txt, and Reference Cell elastic and DIS 19_06_27_12amg.ods
      1. Finished loading the new code to github. The DIS production rates using the new code is similar to XZ's previous results (MC+getrate.kumac) with up to 20% difference. This may be due to Mingy using non-radiated x-section and XZ's were radiated.
      2. Showed calculations wth 2 scintillator bars on for 3He elastic, 3He ref, and N2 ref cell runs. Two is more than enough to cover 3He elastic peak (see Jixie's plots below).
      3. Note on carbon optics: When we do delta (dp) scans, different scitillator bars need to be turned on/off to suppress QE and background.
      4. Question on H2 elastic run: what's the purpose and how should we keep SC bars on/off? XZ followed up on this with Jianping: The purpose is to see if we can measure the well-know H2 elastic x-section. So we should keep SC bars on at where H2 elastic is (or turn all Sc bars on).
      5. Another followup: carbon elastic runs (for momentum calibration) do not need sieve slits in. This should reduce beam time needed. (?? Is this correct??)
      6. Some discussions on the Delta asymmetry run: (1) SHMS can probably go to below 2 GeV in momentum (we need 1.718 GeV) although the resolution will be poor. But for our purpose it should be okay. (2) Question: What do we need for Delta-A for the Delta resonance run? -- followup: see Xiaochao's thesis 3 pager on Delta: thesis_A1n_ssp2_Delta3pager.pdf.
      7. Some discussion on the Cherenkov: we need noble gas CC for 30-deg and heavy gas CC for 12.5-deg running. We should not talk about taking in/out the CC not in use (to avoid mess with them), even though it will help with resolution (since less material = less multiple scattering).
    4. Jixie presented:
      1. Simulated 3He elastic scattering event distribution for HMS xfp_vs_delta_hms_elas.png and xfp_vs_delta_shms_elas. Given the SC bar width is 8cm, keeping 2 bars on is more than enough to cover the elastic peak. - note still need real data to confirm.
      2. Also showed QE vs. elastic data from Marathon (HRS) for 3He P_3he_E1.1685_HRS17deg.png and carbon P_12C_E1.1685_HRS17deg.png (where carbon excited states are clearly visible). Keeping only 1-2 SC bars in this case will keep QE/el to be 1:1, but not sure if HMS or SHMS resolutions are comparable.
      3. see slides A1NRate_Summary_June2019.pdf. First, XZ's code used method 2 which is precise. Method 1 to get rates from MC is faster (since only require calculating acceptance for each event, do not need x-sec event-by-event), but less precise if x-section varies within a bin.
      4. Found carbon elastic data from Hall C's spring 2018 running (run # 1271-1272 for HMS and 169-1690 for SHMS), but not sure if anyone has already processed the data. Will learn how to cook/analyze Hall C data too. -- followup: XZ talked to Doug during the Hall A/C meeting and he is sure the rootfile has been produced. Will find out whom to contact to get the already replayed file.
    5. Xiaochao mentioned at some point we may need to optimize the 30-deg momentum setting, to see if we can maximize the physics outcome at the highest x bin.

  • 06/20/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. Last week: no slides but Jixie mentioned he (1) could not find the proper Hall C optics data taken with 2 GeV beam energy. Holly left and nobody is really in charge of and/or know where to find the best optics data. Mentioned angle calibration with sieve slits does not require elastic events. Will keep looking; (2) Started working on the single arm simulation code. It is called mc_single_arm which appears to be much updated than the mc_(s)hms_single that Xiaochao was using. THe transport function in mc_(s)hms_single.f however appears to be newer -> need check.
    3. See Mingyu's report today on angle scan of PbPt: 3He elastic Kinematics angle%20scan.pdf. Note that slides 2-5 are sufficient to show that even though the asymmetry increases with theta, the drop in x-section and rate determines the smallest angle is the most advantageous in terms of beam time. Thus we will stick to 11.7 deg for HMS and 8.5 deg for SHMS. (Note: slide 6 and beyond are incorrect as it proposes to use 9 deg for HMS, which is not possible. The remaining slides are calculations for ref cell runs also at this incorrect angle setting.)
    4. Here is Mingyu's updated PDF on elastic calculation with the two beam-time estimate tables as suggested last week, for the correct angles of 11.7 and 8.5 degrees for HMS and SHMS, respectively.
    5. Jixie reported
      1. Previously wrote a code "get rate A1n" to calculated integrated xsec or rate using the simple theta/phi/E' binning method and xsection calculation. Did pass generated events to Q1 entrance but no further transport. Thus the integrated rate is expected to be higher than actual.
      2. Worked on mc_single_arm.f to (1) incorporate sieve slits; (2) produce the acceptance map that can be incorporated into "getRateA1n" code so that the latter is effectively a full simulation. The advantage of this is that one can turn on/off scintillator bars in mc_single_arm.f (the most convenient would be to add variables "SC hit", "SC pos x" and "SC pos y" in the ntuple, which allows turning on/off the bar post-simulation and to plot the hit positions across the SC plane); (3) adapt it to elastic simulation; (but it still does not work for QE yet and maybe no need to); (4) added physics variables to the ntuple such as six xsections (either inelastic or elastic depending on the input setting), xbj, W, Q2 etc.
      3. To do: (1) need clear indication of which bar the elastic events hit (see above); (2) Can we compare the SC hit position between el and QE? (Require QE in mc_single, may not be necessary as the acceptance+getrateA1n code can give us the suppression factor for QE rate; (3) Need real data to confirm scintillator bar hit position/size; (4) Need to redo the calculation by using 3He nuclear elastic momentum (Ref cells should be at the same P0 setting regardless of target gas material). The 12C optics should be done at 12C nuclear elastic (for delta=0 run). (5) Use radiated cross sections. However ran into same problem as Xiaochao/mc_(s)hms_single that the radiated x-section code (Bosted et al) runs very slow. May have to use a look up table method (fix TA, TB to polarized 3He cell values).
    6. Eventually when the single-arm simulation code is fully finished we will need to rerun the DIS production rates and confirm with earlier numbers for the run plan. (Fingers crossed that the rate in the highest x bin remains similar!)
    7. One question came into mind while XZ was taking shifts: Instead of target beam time, can we list target BCM goal for each run/kinematics setting?

  • 05/30/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. See Mingyu's report on runplan calculation: Time estimation Elastic and Reference cell.pdf. Quick summary: Completed time calculation for elastic PbPt and Ref cell runs. For PbPt: (A) HMS PbPt: 2/3 from 3He but within the 3He rate only 1/20 from elastic -> 150Hz pure el 3He rate; (B) SHMS PbPt: 3/5 from 3He and within the 3He rate 1/5 from el -> pure 3He el rate about 500Hz.
    3. We suggested (1) doing an angle scan for the spectrometers to see if a slightly larger angle is better (possible if the asymmetry is larger, and grows faster than how steep the rate drops). (2) Until we figure out the suppression factor from turning on/off scintilator bars, suggest giving two tables for the beam time calculation: one including the full QE rate, and one excluding it. The two values will give us the upper and the lower estimate for the beam time.

  • 05/30/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. See Mingyu's report on runplan calculation: Runplan_A1N(part II).pdf with supporting text output file: A1NRates_results_19_05_29.txt. The focus is elastic scattering window treatment (much better now), etc.
    3. To do's: (1) Now with window and QE rates done, need to calculate beam time needed for elastic PbPt measurements (assuming an initial goal of dA/A=2%), and similarly beam time needed to collect 10K events for each reference cell run.
  • Jixie mentioned to suppress QE we can turn off some of the scintillator pads. Similarly for windows. But we need real data to confirm.

  • 05/24/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. See Mingyu's report on runplan calculation: Runplan_A1N_updated.pdf, with a focus on quasi-elastic and elastic rate calculations for optics and elastic calibrations runs, and reference cell rates for both elastic and DIS settings.
    3. To do's: (1) For elastic, glass window treatment is incorrect, need to use elastic calculation for each element rather than using average Z and average A. (2) Some other adjustments to the ref cell and Delta calculations.

  • 05/16/2019 Runplan/Simulation meeting:
    1. Participants: Mingyu Chen, Jixie Zhang, Xiaochao Zheng
    2. For the past month Mingyu has been learning A1n calculation code (from Xiaochao). Has reproduced the DIS and resonance production x-sections and rates and double-checked XZ's work.
    3. Next we will move on to calibration runplan, which is much more complicated. Here is Xiaochao's draft runplan file to start with: Runplan_A1n_draft.doc.

  • 12/10/2018 A1n/d2n Collaboration meeting:
    1. The main meeting webpage and agenda can be found at:https://www.jlab.org/indico/event/299/
    2. Participants: Melanie Rehfuss, Mingyu Chen, Jay Benesch, Gordon Cates, Jianping Chen, Junhao Chen, Dave Gaskell, Brad Sawatzky, Greg Smith, Nguyen Ton, Steve Wood, Xiaochao Zheng, Steve Lassiter, Thia Keppel.
    3. Remote participants: Todd Averett, Wolfgang Korsch, Melanie Rehfuss, Chris Jantzi, Mark Jones, Sumudu Katugampola, Huong Nguyen, and many others,
    4. see meeting minutes here: Minutes_of_A1nd2n_collab_mtg_20181210.docx

  • 06/15/2018 First A1n/d2n Collaboration meeting:
    1. The main meeting webpage and agenda can be found at:https://www.jlab.org/indico/event/272/
    2. Participants: Todd Averett, Joe Beaufait, Jay Benesch, Michael Berkowitz, Ed Brash, Gordon Cates, Jianping Chen, Junhao Chen, Dave Gaskell, Chris Jantzi, Kai Jin, Mark Jones, Sumudu Katugampola, Victoria Lager, Bob Michaels, Vladimir Nelyubin, Huong Nguyen, Brad Sawatzky, Greg Smith, Nguyen Ton, Steve Wood, Xiaochao Zheng,
    3. Remote participants: Mingyu Chen, Shuo Jia, Wolfgang Korsch, Melanie Rehfuss, and many others,
    4. A few important points recorded at the meeting:
      1. June 15, 2019 may be the day to start installation (exactly one year from today!)
      2. To convert from PAC days to calendar days, the standard factor is x2.
      3. From Dave's talk on beamline: need to make sure the 0.25% increase on Linac energy, which will result in both Hall A and C receiving >99.99% polarization. Otherwise, Hall A may receive only 76%. This needs to be measured at the beginning of Fall '19 run.
      4. We need to be more specific about the maximal tolerable transverse polarization of the beam, and the beam energy ec (i.e. Q2).
      5. If the raster system is ready, should try to install in Jan.'19 and test it out. (Mark Jones)
      6. Detectors: all 3 Cherenkov detectors seem to need work. (HMS C needs repair!), but Brad said so far the manpower is okay. For PID performance we should listen to talks at the upcoming Hall A/C collab mtg.
    5. A few important items that we have overlooked:
      1. Delayed helicity recording/reporting: both hardware and software need to be built from (nearly) scratch. We will ask Chao Gu (Duke) for help (JP);
      2. Beam helicity asymmetry feedback: again both hardware and software need to be built from (near) scratch because everyting was torn down after Qweak. We will try to recruit the Ohio group. Gordon will look into exactly what work needs to be done. We need BCM feedback, BPM feedback, and maybe Lumi?
    6. We will have our next weekly meeting on Tuesday 6/26 to finalize our reply to ERR.

  • 05/29/2018 A1n/d2n (bi-)weekly meeting:
  1. Participants: Jianping Chen, Vladimir Nelyubin, Xiaochao Zheng, Brad Sawaskty, Gordon Cates.
  2. We will have our first collaboration meeting on Friday June 15th. Probably half day will be good enough.
  3. Silviu sent his updates before the meeting along with an email explanation, transcribed below:
    1. Update on solid target, 12 GeV raster on:
      1. opticstargets_model_22may2018.jpg is the 3D CAD model I got from Bert and I used in the CFD simulations.
      2. Tflowprofile_ysec_30uA_he3cell40cm__opticsLOW_RadON20lsec_21may2018.jpg is the temperature profile through the He bag in the midsection through the optics targets for the LOW position of the beam on them (electron beam going through all 7 of them). The electron beam raster is a circle of 5 mm diameter and the current is 30 uA. The 4He jets are on at 20 liters/sec and pointed at the 3He cell windows. Radiative heating/cooling is ON.
      3. Vflowprofile_ysec_30uA_he3cell40cm_opticsLOW_RadON20lsec_21may2018.jpg is similar to above. but it is a profile of 4He flow velocity in the same cross-section through the He bag and optics targets with the beam in the LOW position. This profile explains why the foils adjacent to the longest one are the hottest ones, as there is very little 4He flow around them to convectively cool them.
      4. Tprofile_30uA_he3cell40cm_opticsLOW_RadON20lsec_21may2018.jpg is the temperature profile of the whole optics targets and their Vespel frame with rastered 30 uA beam on the LOW position.
      5. Tprofile_30uA_he3cell40cm_optics4high_21may2018.jpg is similar to above but with the beam on the longest optics target in the high position.
      6. Table with maximum temperatures in the optics targets for jet volume rates of 10 l/s and 20 l/s and beam currents of 30 uA and 60 uA (LOW means beam through all 7 optics targets position and HIGH means beam through the middle foil only position):

        10 liters/sec
        30 uA, HIGH, Optics 402 K, frame 353 K
        30 uA, LOW, Optics 527 K, frame 506 K
        60 uA, HIGH, Optics 503 K, frame 416 K
        60 uA, LOW, Optics 697 K, frame 658 K

        20 liters/sec
        30 uA, HIGH, Optics 394 K, frame 352 K
        30 uA, LOW, Optics 468 K, frame 446 K
        60 uA, HIGH, Optics 467 K, frame 385 K
        60 uA, LOW, Optics 594 K, frame 552 K

      7. In summary, basic conclusion stays the same as last week: We will set max current for the solid target to be 30uA and will use the currently planned material for the holder. New results from this week shows that it's safer to use 20l/s flow rate than 10l/s.
      8. Note that the solid target design is slightly different now because of the change from single- to double-hole carbon targets. The beam position for the multi-foil position is 1/16" lower than this simulation. JP said it's not necessary to repeat the simulation with the new multi-foil beam position. (and this can be seen from the vertical dependence of the Tprofile plot posted above, the temperature change due to this beam position change can't be more than 10-20 deg).

    2. Results for ramping the beam to 30 uA on the 12 GeV 3He cell in 30 s. The cell is in the He bag and the He jets are ON with a volume rate of 10 liters/sec (conservative conditions). The summary plots are as follows:
      1. He3_jetsONbeamON_drho_Tmax_28may2018.jpg shows the 3He density loss in % in the upper plot as the beam is ramped on the target from 0 to 30 uA over 30 sec in 6 steps and in the lower plot are the maximum temperatures in the glass beam nipples in the cell windows (Blue = beam nipple OUT, Red = beam nipple IN)
      2. He3_jetsONbeamON_conservation_28may2018.jpg shows in the upper plot the 3He density behavior over the beam ramp up phase in the beam volume in the cell and the whole volume of the cell; in the middle plot it shows the 3He mass fluctuation over the same period from time step to time step (in perfect conditions this should be zero, so these fluctuations, which are a non-conservation of mass in a closed volume can be taken as a systematic uncertainty of the density loss calculations); and in the lower plot it shows the average temperature in the beam volume in the cell over the same beam ramp up period.
      3. Comments: the maximum temperature in the glass cell beam nipples is estimated to reach about 650 K or a Delta_T_max of about 340 K for a 4He jets rate of 10 liters/sec (the calculation for the transversity cell at 15 uA was Delta_T_max of about 100 K, Tmax ~ 420 K, for a 4He jet rate of 20 liters/sec). The average temperature in the beam nipples is about 80 K lower than the max at 30 uA, so there is a substantial temperature gradient. For one thing the cooling jets rate in this simulations is half the one considered for the transversity cell. The convective heat transfer coefficient from the glass to the 4He jet(s) is less than half for this cell in these conditions than the value for the transversity cell in its conditions, which explains the larger Delta_T_max. I should probably also try with cooling jets volume rate of 20 liters/sec for this cell and maybe try to optimize the jets orientation on the glass windows. When I considered vacuum outside the 3He cell I got a Tmax in the glass windows of 700 K, with jets at 10 liters/sec it seems to be 650 K, better, but not as much as expected. 3He density loss with vacuum outside it was calculated to be about 8% at 30 uA, and with 10 l/s jets is calculated to be 6% (averaged over the beam volume).
      4. Gordon commented the glass annealing temperature is 700C so the absolute value of the temperature shown here is not a problem. We are not sure what the increase in the temperature diff mean (see to do below).
      5. Looks like a ramping rate of 30uA in 30seconds isn't that bad. Temperatures stablize faster than that.
      6. Other profiles:
        • Jets_at_window3_hecell40cm_hebag_apr2018.jpg is the profile of the cooling jets velocity vectors brushing a glass cell window, colored by magnitude of vectors (this profile shows that the jets do not seem to be well oriented, as the flow velocities are very small on the beam nipples, this is part of the explanation for the large temperatures in these nipples at 30 uA).
        • T_ysecprofilebag_he3cell40cm_hebag_beamOFFjetsON_time_23may2018.jpg is the temperature profile through the mid horizontal section through the He bag and the 40 cm long He cell and the glass walls of the 40 cm long cell with beam OFF but jets ON. The He bag wall is not adiabatic, it can radiatively exchange energy with the environment and the environment is considered at 300 K.
        • V_ysecprofilebag_he3cell40cm_hebag_beamOFFjetsON_time_23may2018.jpg is the 4He flow velocity profile through the same cross section as 2. above, it shows that the jets are asymmetric on the glass cell windows.
        • Vyprofilecell_he3cell40cm_hebag_beamOFFjetsON_time_23may2018.jpg is the y projection of the flow velocity (Vy) profile through the vertical mid section of the cell and bag with an upper cut at 3 cm/s and a lower cut at -3 cm/s. This shows (I really like this profile) the 3He gas convection through the glass cell: ascending gas on the side of the heater ball and descending gas on the other vertical leg.
        • Vyprofilebag_he3cell40cm_hebag_beamOFFjetsON_time_23may2018.jpg is the same as above but with no cut on Vy (the previous profile zoomed inside the glass cell, this profile shows the jetting outside it). This one shows how the He jets are split vertically upon hitting the glass windows.
      7. For reference these are the material properties I used in the CFD simulations: Vespel-SP1: thermal conductivity 0.35 W/m.K, specific heat 1130 J/kg.K, density 1.43 g/cm^3; graphite: thermal conductivity 140 W/m.K, specific heat 720 J/kg.K, density 1.63 g/cm^3, total emissivity 0.82, electron energy loss in graphite 2.265 MeV/(g/cm^2), expected beam heating at 30 uA in 0.254 mm thick graphite about 3 W, which goes to about 6 W at 60 uA. 4He jet(s): properties taken from NIST webbook at 1 atm and 300 K and treated as ideal gas under gravity 3He in cell: treated as ideal gas, initial conditions: with the whole cell at 300 K, pressure is considered to be 10 atm in the cell.
    3. To do's: will finish the unrastered beam on optics; optimize the jet orientation on the 12 GeV cell; run a 12 GeV simulation with 20l/s;
    4. Other to do items we suggested at the meeting: transversity simulation with raster off; add thermal stress to all simulations on the glass cell (transversity raster ON and OFF, 20l/s; 12 GeV raster ON 10l/s and 20l/s).
    5. Some questions: how sensitive is the Tprofile to jet orientation? (We don't want this to be too sensitive as it is not easy to fine-tune the jet in the real running situation.)

  • 05/22/2018 A1n/d2n bi-weekly meeting:
  1. Participants: Jianping Chen, Vladimir Nelyubin, Xiaochao Zheng, Dave Gaskell, Jay Benesch.
  2. Vladimir presented the field map calculations for the worst-kinematic condition, see talk_05222018.pdf. These were done both without correction coil, and with the By correction coil only. Correction coils for Bx and Bz have not been added. Main findings and comments:
    1. JP re-emphasized that for transversity the field gradient had to be kept at 30mG/cm or above, otherwise there was masing. This corresponds to a gradient amplitude square of (0.03)^2=1E-3.
    2. p.10, the sharp/large field gradient at z=~-20cm is due to the samll size of the correction coil.
    3. p.15, gradient seems to be too large.
    4. p.18, field gradient at the transfer tube should be as small as possible to allow for pulsed NMR. (JP) needs below 1E-5 for (partial Bx squared +partial By squared)/B sq (which is about 75mG for a 25G holding field). This plots shows slightly above 1E-5 but probably okay for now. Fine-tuning will be done close to running.
    5. Last page: field direction is about 1 deg off, for both long and trans settings.
    6. To do's: We need to add the Bx and Bz correction coils to minimize gradients in these two components. These two pairs of coils are existing and are the same size as the main holding Helmholtz coils. Adding these needs Steve to run his calculation again (Vladimir cannot add the coils).
  3. Silviu reported on the thermal simulations for the carbon foil target (raster on only). Main conclusions:
    1. cooling flow used is 10L/s. T changes by only ~10K for 20L/s so all results shown are for 10L/s only.
    2. for 30uA on single foil, Tmax on foil is 402K, on holder is in the 300's.
    3. 30uA multi foil: similar, with the pair next to the central foil being the most heated due to jet orientation.
    4. 60uA on multi foil: problematic! Tmax on holder is 552K. JP said there are two possible materials for the holder, both are high-density plastics. The currently planned one has a melting point of 250-300c (520-570K) which will melt. It is possible to use a higher melting point plastic (by at least 50K) which used to be used to make the oven (now oven is ceramic), but it's more expensive.
    5. If we can live with 30uA max on carbon then things are okay, we can use the currently planned holder material. -- will followup on this by email.
    6. Silviu will send the table of all Tmax and figures to Xiaochao for book-keeping.
    7. To do: Jianping mentioned the max current should be both 30uA or 60uA for both raster on and off. Silviu will repeat the simulation but with raster off.
    8. Will also continue with 3He simulation for 12 GeV conditions.
  4. The preliminary beam schedule for 2019 has been posted, see 20180511.0_ExpSch.pdf. Xiaochao will email Thia about 1-pass running time.
  5. We need to decide on the date of the first A1n/d2n collaboration meeting. Preferably on a weekday so JLab techs can participate. SHould also invite Hall A and Hall C Collaborations.
  6. Xiaochao will add all proposal authors to the A1n_d2n mailing list.
  7. Xiaochao will look into how to fill out the beam time request.

  • 05/15/2018 A1n/d2n bi-weekly meeting:
  1. Participants: Jianping Chen, Vladimir Nelyubin, Xiaochao Zheng, Zein-Eddine Meziani, Dave Gaskell, Jay Benesch.
  2. We have our own mailing list now! A1n_d2n (the 6 GeV list was named A1n_g2n).
  3. Silviu reported on the target thermal simulations:
    1. Finished simulation for the 6 GeV transversity running condition. Here is a plot Silviu_page1.png to show the beam effect for a 40cm target, 15uA (ignore the title in the plot), beam current ramping rate of 0.5uA/sec (or 15uA/30sec), 4He cooling jet on, and a 3x3mm sq raster. The density drops by 4.4% in the beam volume. Jianping commented that (1) temperatures at various positions of the cell are in agreement within +/- 5 deg with the measured data; (2) the pressure/temperature seems to stablize fast, ramping rate can possibly be doubled to 1uA/sec (for 12 GeV this will be 30uA/30sec).
    2. Next will continue to 12 GeV condition and solid targets.
    3. Will also show thermal gradient and stress, etc. (with help from Steve L.)
  4. Xiaochao calculated the 3He elastic xsec and rate for 2- vs. 1-pass beam (2.1 GeV/pass) at 12.5 deg, and found xsec drops by about 1000 from 106.99 nbarn/str to 0.105 nbarn/str. In addition the asymmetry drops from 5.89% to 1.78%. This makes it impossible to do elastic PbPt at 2-pass. This estimate has been checked by Alexandre Deur (raw xsec 0.130 microbarn/str and asy=5.01% at 1-pass and 5.58E-05 microbarn/str and asy=0.84% at 2-pass) and Zein-Eddine (xsec=0.18E-30 cm2/str at 12.5 deg and 2.1 GeV, and 0.27E-33 cm2/str at 12.0 deg and 4.2 GeV - his code does not work for 12.5 deg for the latter case). Will also check if it is possible/better to detect the 3He (ZEM).
  5. In addition to elastic PbPt and Delta we also need to do all optics calibrations at 1 pass. 2-pass is possible but will be much slower (and if we use 1-pass anyway for elastic we may as well do optics there). We will need to finalize how much beam time is needed for optics calibrations (Mark Jones + Brad). Tentatively we think 2 full days of 1-pass running will be sufficient (2 shifts elastic, 1 shift Delta transverse, 2 shifts detailed optics), plus ~ 1 shift for pass change. For Spring 2020 run the optics may be shortened. The 1-pass beam running needs to be communicated to Thia ASAP.
  6. We then went through the working document replying to ERR final report. Here is the draft by the end of the day: 3He_ERR_Reply_to_Final_Report_May2018.docx. Some items that could not be reconciled at the meeting are summarized below:
    1. Solid target current: do we want to set the max at 30 or 60 uA?
    2. how do we protect the target from possible large beam halo and mis-steering?
      1. option 1: have two carbon hole targets. But current target design does not have enough space to have both the central long carbon with hole and two additional carbon holes at +/-z edges; (see drawings: 67507-00047.pdf and 67507-00087.pdf). A sketch of the target ladder can be found here: target_structure_edit.png. (Jianping) One option is to sacrifice the central single foil (why? how?). Do we need single-foil optics?

        (Jay) If we do have double carbon holes, steering the beam into both may require a shift of tuning. On the other hand this needs to be done only occasionally.

        (JP) double holes: may be hard to see the 2nd hole because of beam multiple-scattering on the first hole.

        (XZ) Can we modify the central foil to have the hole position below the single foil position, and modify one of the upstream foil to be longer to provide the second hole (but not as long as the central foil)? The alternate sketch is here: target_structure_edit_alternate.png.

      2. option 2: Leave the solid target design as is, and rely on a combination of: single carbon hole, ion chamber, upstream collimators, periodic harp scans to check beam size -- to protect the target?

        For the upstream collimators here are Jay's inputs: the three one-meter dipoles provide a 22 mm square collimator of sorts ending 4.3 m upstream of the pivot. An additional collimator could be placed on the superharp girder. In this drawing haysg_JL0059798-MODIFIED_SUPERHARP_GIRDER.pdf one sees ~25 cm (item 10) in the middle of the girder which could have a copper or W90Cu10 collimator placed in the spool piece. The position of this possible collimator will be about -200cm upstream of the pivot. Halo wandering into that would certainly trip target ion chambers.

      3. More FSD?Can we interlock more stuff with the FSD to protect the target?
  7. We hope to hear report from Gordon/Vladimir at the next meeting. At the least should make plots for the field maps at the highest Q2 point (A1n highest x).

  • 05/08/2018 A1n/d2n bi-weekly meeting:
  1. Participants: Jianping Chen, Vladimir Nelyubin, Xiaochao Zheng, Gordon Cates, Zein-Eddine Meziani
  2. Silviu provided his heat simulation by email:
    1. both the Transversity and the 12 GeV cells have been modeled with actual 4He bags outside them. For the 12 GeV cell I'm attaching a temperature profile in a mid vertical cross-section through the He bag and cell, with beam off and 4He jets off (4He is in natural convection in the bag generated by the temperature gradients between the cell and the bag's walls). The cases that I'm running for the transversity cell are: thermal equilibrium (TE) without beam and jets off, TE with jets on and beam off and TE with jets on and beam on. TE with beam off and jets off has been done, the other two are pending. The beam will be ramped with the rate I got from Brad. The same three cases are being pursued for the 12 GeV cell. TE with beam off and jets off and beam off and jets on have been done. I'm currently running cases with beam on and jets off and the last one will be beam on and jets on. It took a long time to model the He bags correctly, but in a week most of these cases should be completed and I will start circulating the results,
    2. We discussed briefly the plot and the temperature profile look reasonable. Looking forward to results in one week!
  3. We discussed the work on field gradient:
    1. There were some confusion about how may correction coils already exist. (JP) There are a total of 4 sets of correction coils: 2 sets for the vertical field (By), and they are oriented at +6deg and -6deg w.r.t. the horizontal plane. Then there is one set for Bz and one for Bx (both fields are horizontal, Bz is along beam and Bx is transverse horizontal). We should use these 4 existing sets. Because the target design work is already done, changing or building new coils will add significant challenge to the design progress and cost, although not completely impossible.
    2. Gordon reported that they received the field map from Steve for the worst case scenerio (presumably the highest Q2 of A1n on SHMS, but this needs to be confirmed), both with and without the two sets of the vertical correction coil. Also found the HMS setting has very little effect on the target. Vladimir mentioned we might need a bigger coil (by 10cm) but see JP's comment above.
    3. Gordon said he may ask Steve to pick two other (second worst or typical) settings to generate the field map, but (according to JP) Steve is doubling his work by picking up the work of Paul Brinza, who is retiring. So adding more settings may be difficult.
    4. Th
  4. Jianping sent the final Target OSP+Laer OSP (LSOP) to Bert for approval. Still working on the 3He target "milestones" requested by the ERR cmt.
  5. Xiaochao mentioned the new student, Mingyu Chen, has been working on the SHMS NGC performance analysis. Still working on the error calculation, may not have final results within the month because of the final exam and Mingyu is going back to China right after. Will continue in the summer, and will proceed to general PID analysis after the NGC is done.
  6. Target at JLab: Junhao is almost done with the kappa0 work at WM and then will be learning target work from Kai Jin at JLab. Mingyu will move to JLab after his China trip in the summer, but given that there are six students, we need to discuss who will be the 2nd student working on the target. Other students need to work on Moller and BPM as hardware tasks.
  7. We discussed the possiblity of having a collaboration meeting in the morning of Saturday June 9th, immediately after the SoLID collaboration meeting (June 7th and 8th).
  8. We need our own mailing list. Xiaochao will look into this. We can simply use "A1n_d2n" (different from the 6 GeV list "A1n_g2n").
  9. Xiaochao will start modifying the draft reply according to the final report (from the preliminary bullet items).
  10. Jianping mentioned the scheduling for next year is done. (What?? Xiaochao is utmostly confused.)

  • 04/10/2018 A1n/d2n bi-weekly meeting:
  1. Participants: Jay Benesch, Dave Gaskell, Silviu Dusa, Brad Sawatzky, Vladimir Nelyubin, Xiaochao Zheng, Gordon Cates, Wolfgang Korsch
  2. Final ERR report is out and has been posted on the ERR page. A preliminary task+reply list can be found at https://docs.google.com/document/d/1CDfb-PXcE0gTRR2NUFHarxp694ItSLnvkSffWk4Dc14/edit?usp=sharing
  3. Siliu showed the first results on beam heating calculation on target, see He3_40cm_30uA_10apr2018.pdf. A few explanations and comments:
    1. Beam has uniform distribution across the size (no rastering yet); target cell is in vacuum (see to do); 30uA beam turn on instantly. Under this condition a cooling flow of 1.5l/min seems to keep the cell at reasonable condition;
    2. slide 3 and 4 show results under the 1.5l/min cooling flow. Effect on density is quite sizable, up to 8% loss at full current.
    3. On slide 4, cell temperature seems to stablize within 1-2 seconds of beam on.

    And to do's:
    1. add 1 atm 4He in the scattering chamber. Convection will help lowering the cell temperature;
    2. Silviu will talk to engineers to see how to deduce the stress on glass from the temperature profile, and estimate the safety factor (vs. breakage);
    3. Will simulate beam ramping for the typical 6 GeV condition. Since we know no glass cell broke with the raster on, this will provide a check point for the calculation. (XZ's recollection is: 12 uA, beam size 4x4mm or similar, ramping rate 1uA/10seconds. Brad will check this too.)
    4. Repeat above but with raster off. Since we know one glass cell broke due to raster being left off by mistake, this will provide a second check point.
  4. Gordon on field gradient: need a table of all kinematics (Brad, XZ), then choose a subset to study/model in detail. Need inputs on maximum Bx and By that can be tolerated by physics (Brad, XZ).
  5. Wolfgang: found parts of the field compass during ERR but need to put them together and see if they work.
  6. Brad and XZ need to determine maximal Moller uncertainty that can be tolerated by physics.

  • 03/19/2018 A1n/d2n ERR:
  1. Please see final talk slides on the ERR wiki page. Talks went well, but there is much work left to do before we can request for beam time. While we wait for the final recommendations from the committee, we already received "closeout bullets" from the cmt chair. We will have bi-weekly meetings every other Tuesday to work on this.

  • 02/15/2018 Meeting to discuss A1n/d2n ERR:
  1. Participants: Jay Benesch, Dave Gaskell, Brad Sawatzky, Vladimir Nelyubin, Xiaochao Zheng, Steve Lassiter, Gordon Cates, Jianping Chen, Jixie Zhang, Zein-Eddine Meziani, Wolfgang Korsch, Todd Averett
  2. Target field work: Steve has generated tables for the elastic and the Delta measurements of A1n and has sent them to Vladimir. Will continue to other kinematics next.
  3. We discussed Xiaochao's draft slides and had a long debate about the raster size. We settled with "up to 2.5mm radius" as the requirement, but this needs more backup material. To-do list:
    1. Confirm the raster size/shape and the current used by 6 GeV A1n (2mm dia circular Zheng's thesis), transversity (5x5mm square Jin's thesis), and d2n (5x5mm square long paper???). Scale by current power density. Input from Jay: the max size using the existing raster system is about 2.5mm radius. If we need larger size then a new raster system is needed which can provide up to 4.5mm radius (but XZ didn't catch the name of the new system and there was a consense that we should avoid the requirement of the new system).
    2. Transverse beam polarization: see study from E94010 and add a backup slide.
    3. Moller measurement: probably need one every 7-10 days.
    4. beam charge asymmetry: Jay's number is 20,000ppm per helicity pair. This should give 20ppm per 2.5min beam ramping but needs to be confirmed. (This was a big issue for the 6 GeV A1n/g2n).
    5. Q2 requirement: Jay's number is Ebeam can be known to E-4 relatively but absolute uncertainty is 3E-3. Angle can usually reach 0.5mrad=0.03 deg for the HMS (and 0.06 deg can probably be reached for the SHMS). XZ will study more carefully the precision need for the Q2.
    6. Brad will prepare to show slides on the optics calibration.
    7. Task list for analysis should be divided to students and carried out in parallel. A total of 3 year towards the first publication is too long (the 3-year is actually 3-year-student but dividing by 3 students = 1 year is probably too short).
  4. Jay's slides: need details on the raster system.
  5. JP reported target work: Todd started working on the ref cell system but we are now building up the inventory (Jixie) and testing existing hardware. THere was some progress on the 5-1 laser combiner. Also found the source files for the TOSP!
  6. Jixie is updating his radiation simulation. Previous one did not use the final lead house location and also used a lower density for the target. This will be fixed within the next week or two.
  7. Brad organized an installation meeting last Tuesday and here are his notes: Minutes from installation meeting on 2/13

  • 02/06/2018 Meeting to discuss A1n/d2n ERR:
  1. Participants: Dave Gaskell, Brad Sawatzky, Vladimir Nelyubin, Xiaochao Zheng, Steve Wood, Gordon Cates, Jianping Chen, Jixie Zhang
  2. Latest news: the ERR has been moved to Monday March 19th!
  3. Our draft agenda (see last week's link) has been accepted by Patrizia. There is a slight change to ERR charge No.8, see updated version: HallC_A1n-g2n_new.pdf
  4. Made a dedicated Hall C experimental webpage for A1n: https://hallcweb.jlab.org/experiments/A1n/. The directory on jlab CUE is /u/group/hallc/www/hallcweb/html/experiments/A1n/. Please let Xiaochao know if the group permission does not work for you.
  5. Made a dedicated wiki page for the ERR, see https://hallcweb.jlab.org/wiki/index.php/A1nd2n. Please help to put the material together.
  6. Xiaochao's draft slides (1st talk): A1nERR_Zheng_talk1_20180206.pdf. A few comments from the meeting: 1) the 7.5 GeV/c momentum setting for the HMS is too high (resonance setting). Need to remove this. (Even the 6.80 GeV/c is pushing it but we can leave it in for now); 2) need table of requirements, give exact numbers; 3) PID: need send numbers to Brad to compare with PID performance we expect to achieve through the Sp2018 run; 4) mention this is the first experiments that require the use of polarized beam in Hall C after the upgrade; 5) Moller time: need typically 1 shift for tune-up hen 4 hours of measurements (actual hours) at each energy; 6) Moller needs dedicated commissioning time (that probably shouldn't come out of our beam time); 7) beamline, spectrometer and target commissioning should be shared between A1n and d2n; 8) need to be clear if we are talking PAC or actual hours. 9) Jixie commented he used 10amg for his radiation simulation, and needs to be corrected to 12amg.
  7. Jay's draft slides: a1n_d2n_23mar2018_Jay_Benesch.pdf -- to be discussed next time.
  8. Update on the raster design and safety: Jianping had a meeting with Mark Jones last Friday. Mark will give one slide to Jay to present at the ERR. There is no need for a new OSP.
  9. Target work -- Jianping talked to Chris about the FSD interlock for the target (lock on the beam current?).
  10. Update on the target field simulation -- Steve L. hasn't had too much time working on this. Vladimir is planning to go to JLab this week (possibly Thursday) to discuss this work.
  11. "New" safety document status:
    1. Brad is keeping an eye on the RSAD, no need to worry.
    2. Moller: Dave's job.
    3. Target TOSP: Jianping reported the latest github source code is missing. Doug will look into this within a week. In the worst case, there are previous versions available (as well as hard copies) from which we can put together a new one, but it will take much longer time.
  12. "Standard" Safety document status -- Jixie download the Hall C standard github directory and Xiaochao looked into it. See below:
    1. Managed to compile the level 4 (Hall C standard OSP), this does not need any modification;
    2. Managed to compile the level 0 (ESAD), need:
      1. add Moller to around section 3.3, see maybe previous example.
      2. add pol3he target (~1 page), around section 3.4, see maybe transversity ESAD for examples. The first GDH experiment's docs can be used too because it was the first in Hall A that used this target.
    3. Made a draft COO. here are discussions at the meeting:
      1. p2: the ESAD will refer to all new OSPs, no need to add the individual ones here;
      2. p3: SAF120 refers to this COO/signature;
      3. p4: graph needs update. Walter Kellner as the Hall C work coordinator and Greg Smith as the Physics liaison;
      4. p9: no need for engineering liaison -- ok
      5. p9: need to make sure we require 2 minimum shift workers;
      6. p15: use "optional 3rd person" or leave this out; (JP) need special procedure for installation, laser controlled period of the hall, also leave raster from this appendix.
    4. ERG: we decided there is no need to modify this. There may be a need to add the information about the target platform before the run, but not right now.

  • 01/30/2018 Meeting to discuss A1n/d2n ERR:
  1. Participants: Thia Keppel, Gordon Cates, Jixie Zhang, Jianping Chen, Zein-Eddine Meziani, Wolfgang Korsch, Xiaochao Zheng
  2. Jay confirmed he can give the talk on item 6 and has already sent us draft slides (to be discussed next time).
  3. Update from Brad (had to miss the meeting as well):
    1. Pavel has completed the Hall radiation assessment. I'm working with him to get it incorporated into a draft RSAD.
    2. JP and I will be meeting early next week with an 'Installation' group (probably Feb 6) to work on installation and infrastructure related items. I expect to run over at least the following items:
      1. Large-scale mechanical/engineering installation requirements: 1) rough target installation; 2) pivot, beamline modifications; 3) target laser platform; ...
      2. Pol He3 target support hardware: 1) control rack locations (+shielding issues); 2)new patch panel(s) between CH and Hall C; 3) field coil power supply placement; 4) gas system requirements; 5) laser fibers
      3. Slow control requirements: 1) Laser controls; 2) Gas system controls
      4. Identify laser safety interlock requirements
  4. Thia instructed us that we need to:
    1. have a dedicated wiki page for the review. This needs to be ready 2 weeks prior to the review for the committee to read all material; Here is the ERR wiki page for PREX/CREX .
    2. Re committee: Patrizia will put together the committee and will let us know in advance who are on it.
    3. Need a physics division liason. Problem is most of staff we work with are also PIs. Dave G. is already liason for PREX/CREX. -- sth to think about.
  5. Jianping updated:
    1. Field gradient: Steve is already working directly with Vladimir. WE suggested to have one of them give weekly updates.
    2. Will have a meeting on Friday with Mark Jones to start the raster design. This must be ready by the ERR. will also talk about writing the OSP for the new raster.
    3. Asked what the angle precision for the target field needs to be. For previous running Wolfgang worked on this. Discussed where the previous equipment could have been in all these past years (???).
  6. Documentations:
    1. from Steve, Commissioning Document about Hall-C current commissioning plans are at: https://hallcweb.jlab.org/safety-docs/Fall17Spring18/ and https://hallcweb.jlab.org/wiki/index.php/Fall_2017_Commissioning_Plan and https://github.com/MarkKJones/fall2017-plans
    2. Thia emphasized we need an OSP for every new equipment we bring in. This means: target (JP), raster (MJ?), and Moller (DG)
    3. Jixie will help to get the source fils for all documents (from Steve probably).
    4. Xiaochao will download all other documents and see how much work is needed to incorporate A1n/d2n.
  7. Finally, here is the preliminary agenda as of today: Agenda_ERR_A1nd2n_draft3.docx
  • 01/18/2018 Meeting to discuss A1n/d2n ERR:

  1. Participants: Todd Averett, Gordon Cates, Dave Gaskell, Jixie Zhang, Jianping Chen, Brad Sawatzky, Zein-Eddine Meziani, Wolfgang Korsch, Xiaochao Zheng
  2. Update: the ERR has been rescheduled to Friday, March 23rd. Note that it is expected we have all the slides done (final) one week prior (that's 3/16), which means we need to discuss all draft slides during the two week period of 2/27-3/2 and 3/5-9.
  3. The focus today is to start finalizing the agenda, here is an updated draft: Agenda_ERR_20180323_draft2.docx
  4. All items on the to-do list from last time still stand.
  5. We worked on assigning names to each of the items on the ERR charge HallC_A1n-g2n.pdf. These are itemized below:
    1. Item 1: overview of the two experiments -> XZ, BS's talks
    2. Item 2: JP, DG's talks
    3. Item 3: JP, needs to work offline with Steve L (Tosca simulation), and with Gordon and Vladimir on the gradient coils. Immediate to-do items include: Steve will send his results to Vladimir once finished. DG reminded that we need a defendable timeline
    4. Item 4: Beamline will be covered by DG's talk, also need to work with Mark Jones on the raster/beam current ramping; Spectrometer/detectors will be covered by BS.
    5. Item 5: This will be divided and covered by separate talks: slow control (BS+ target group also need to talk to EPICS group), radiation - Jixie and Pavel will focus on this and covered by the overview talks, target - JP, beamline/raster - DG, SHMS - BS, shift taking + analysis - XZ, summary of manpower/responsibity - XZ.
    6. For item 5 we need an overall installation talk by Walter Kellner. Need local meeting to discuss installation (Agenda TBD?)
    7. Item 6: need Jay Benesch to talk about this. Need to emphasize ramping/raster/soft+hard+FSD interlock etc.
    8. Item 7: covered by talks possibly of XZ and BS: shielding/radiation (see Jixie's report below)), radcon (Pavel is expected to finish calculation also within a week),
    9. Item 8: documents: JP already started working on LSOP and OSP, awaiting feedback from Bert on what need to modify. Jixie will collect the latest documents on all others (COO, ESAD, RSAD, ERG) and we will see which need to be modified. It is expected that we have a draft ready for each of the required document by the ERR.
  6. Jixie finished the radiation calculation for within the hall, see Jixie_A1ND2N_Radiation_20180110.pdf
  7. For our next meeting we suggest everyone start putting together slides for the respective talks, following the draft agenda posted above.
  8. Future meetings: T 1/30, T 2/6, Th 2/15, Th 2/22, Th 3/1 (slides review part 1), Th 3/6 (slides review part 2), Th 3/15 (finalizing slides).
  • 12/12/2017 Meeting to discuss A1n/d2n ERR:

  1. Participants: Todd Averett, Gordon Cates, Thia Keppel, Jixie Zhang, Jianping Chen, Brad Sawatzky, Zein-Eddine Meziani, Xiaochao Zheng
  2. We continued the discussion on the target preparation:
  3. The 3% requirement on polarimetry means we need to know the kappa0 to 2%. Both Todd and Gordon's group are on this and we will keep tracking the progress.
  4. Todd will look into the reference cell system, including gas control (target ladder already designed by Al).
  5. update on gradients: Steve needs to put the target coil into the calculation, but now he is busy with SHMS problems so not sure when this will be done. The magnet iron of the spectrometers will affect the target field.
  6. the optical fibers need high power tests. These were tested at 30W but need 40-50W. Gordon prefer to see 120W on target. JP mentioned the 110-m long fiber will cause a 10% loss in the transported laser power.
  7. Need update from Gordon on cell making status, need 5 good cells by the start of the run.
  8. Do we need higher beam current for the optics? - XZ will look into this.
    • 12/08/2017 Meeting to discuss A1n/d2n ERR:

    1. Participants: Wolfgang Korsch, Todd Averett, Gordon Cates, Thia Keppel, Jixie Zhang, Jianping Chen, Brad Sawatzky, Xiaochao Zheng
    2. We received the ERR charge, see HallC_A1n-g2n.pdf
    3. Both WK and TA said they will be at JLab for the ERR meeting. GC is a "maybe".
    4. Brad already sent the beam/target conditions to Pavel for the radcon calculation.
    5. Thia told us that the radcon group is only responsible for the external hall radiation. The collaboration needs to calculate the radiation in the Hall and how much shielding is needed for the hall equipment. Jianping will discuss this with Jixie and Brad soon because this is also needed by the engineering design. (Gordon's group had experience with PREX but that was for Hall A).
    6. Jianping talked to Jay about beamline requirements. Looks like apart from the circular raster and the Moller polarimetry all others are quite standard. Thia said it's not trivial to make the circular raster to work (but Jianping said we do need the circular one, not square rasters). This may need the beamline group to put A1n/d2n prep at a higher priority.
    7. We need to draft a schedule for the ERR meeting, and identify tasks with responsible names. Here is a first draft for the agenda: Agenda_ERR_20180302_draft1.docx
    8. Jianping presented a few slides on the polarized 3He target status: polHe3-A1nd2n.ppt. We only had time to discuss the gradient coils (to compensate the SHMS bender field):
      1. Steve L.'s work on the vertical gradient (for this work the target coils are not in yet), see page 4 and 5. These coils were designed for SAGDH. Results on the gradient compensation is nice. The difficulty with the vertical gradient coils is that they need to be winded in-situ do to not knowing the exact geometrical constraint until the target installation is mostly done.
      2. The horizontal field will be distorted and need compentation coils. The compenstation coils are existing (old ones) and are attached to the main Helmholtz coils. Steve will continue working on this.
      3. Steve will also add the target coil to the calculation.
      4. For polarimetry the NMR (top,bottom) and EPR(top only) will be performed from time to time, while pulsed NMR will be done online and more frequent. Gordon reminded that pNMR's requirement on the field gradient is more stringent than AFPs.
    9. Gordon would like to test the metal end-windows. JP said that's okay.
    10. We will continue the discussion on Tuesday 12/12.
  • 10/17/2017 Second meeting to discuss A1n/d2n ERR:

  1. Participants: Jianping Chen, Brad Sawasky, Todd Averett, Jixie Zhang, Xiaochao Zheng
  2. The ERR has been scheduled for Friday, March 2, 2018. Please mark your calendar, plan for childcare (if applicable) and travel to JLab to attend the meeting.
  3. Beamline update: JP talked to Jay. The SHMS bender magnet affects the beamline but this should be taken care of before A1n/d2n.
  4. raster? responsible people are Bill Dunning and Dave Gaskell. We need to make sure circular raster works (no reason why it will not work, since it worked before).
  5. Polarized 3He target status (JP, also information from Gordon collected by Xiaochao on 10/18):
    1. For the five cells ordered, Mike is starting to make a new batch of windows. (The windows will then be tested at UVa, then back to Princeton for cell making)
    2. JP said if we have two good cells by the beginning of the experiment, we will be good enough. Gordon said 8-10 cells are needed by the start of the exp.
    3. For the 3 cells made in 2016, one blew up (due to insufficient amount of glass being used), one was back-filled by mistake at UVa, and the 3rd one has not been fully characterized yet at UVa but will be sent back to JLab for testing. UVa's old laser for the wall thickness measurement is retired/broken. A new laser was bought but the software is not in place yet for the scanning measurement. For the cell that was back-filled, it is planned to re-fill it and test out, for practising purposes.
    4. Gordon suggested we may want to/could (choose the word you prefer) have at least one cell with metal-end windows just to test the metal window design and to see if we can maximize the statistics by crancking up the current.
    5. kappa0 measurement at WM (Todd): Junhao is making good progress and should complete it within the next couple of months. The goal is 2%.
    6. kappa0 measurement at UVa (Gordon): Sumudo Katugampola is the student working on this. He is trying to make the new target system work but will take a while before he can get to the actual kappa0 measurement.
    7. Oven at JLab (JP): the new oven is now installed and being tested. It shows good heat retention. But still need to test the EPR system (EPR is located further away from the cell compared to the old oven).
    8. EPR: need to make sure it works with the new oven and also need to improve EPR's light collection. This will be Kai Jin's last task before he moves out of the target work. Unfortunately Kai is in China now due to family reasons.
    9. NMR: testing pulsed NMR with the new oven now. Another task is to do pNMR with the lock-in instead of on a scope. This is the last task of Nguyen Ton before she moves out of the target work.
    10. It's hoped Junhao will start the target work at JLab after he finished the kappa0 work at WM.
    11. fiber installation: JOe and Jack are almost ready to install the fibers.
    12. laser box, platform, gas system, target ladder, etc: design is all done. Will try to get as much installation done as possible, gradually.
    13. Bender field: Steve L. already finished analyzing the bender field on the target and the effect of the correction coils. No big problem forseen and the correction coils should be sufficient to correct the field gradient to a satisfactory level. The last remaining task is the effect of the bender magnet's steel on the target, which Steve will continue to do but he will be busy with the upcoming run so the earliest time to get this done maybe early months of 2018.
  6. Student update: Xiaochao's student will be Mingyu Chen, moving to JLab in summer 2018. Emlyn Hughes will have a Ph.D. student. And Wolfgang will possibly have a Ph.D. student. We would like to cap the number of students at six which would be: 1 from XZ, 1 from TA, 1 from EH, 1 from ZEM, 1 from WK, leaving at most one from China.
  7. At some point we will need to pair PI/students with each specific task and work with the corresponding JLab staff.
  8. Before next month/meeting, Xiaochao will draft an analysis flow chart, and maybe also a task list with names and the timeline.
  • 08/03/2017 First meeting to discuss A1n/d2n ERR:

  1. Participants: Jianping Chen, Dave Gaskell, Zein-Eddine Meziani, Brad Sawasky, Jixie Zhang, Xiaochao Zheng
  2. The main purpose of this meeting is to prepare for the A1n/d2n ERR. We prefer the ERR to be held no later than February 2018, so that we can request for beam time in summer 2018. If all goes well we hope to run A1n/d2n in Fall 2019, after the Hall C commissioning experiments (10 weeks in 2018?), and use the long summer shutdown for target installation. There are a few short, "standard" expeirments (pentaquark, x>1/EMC) that may run either before or after A1n/d2n.
  3. We expect the formal ERR charges will be given to us about 1 to 2 months prior to the actual review.
  4. We started from reading BONUS12 ERR and the 12 GeV Pentaquark ERR. (For BONUS12 ERR please see https://userweb.jlab.org/~kuhn/BONuS12/Documents/ERR/ Most of the items are common to all experiments and we expect these to be manageable after we receive the ERR charges. Thus we focused on discussing the two main requirements: the polarimeter and the polarized 3He target.
  5. We decided to not require the Hall C Compton. The Compton was completely taken apart after Qweak and it will require significant investment of Dave's time and manpower from the users. Plus, the focus now of Dave is on PREX.
  6. The Moller has refurbnished magnets, etc, and may need new foil(s) and commissioning time, but it's a much simpler system to commission. We will request 2% uncertainty from Moller. Moller will likely require 20% of a graduate student to work on the analysis offline.
  7. We need OSP of Moller, target (LSOP), etc.
  8. The ERR usually requires an analysis flowchart that allows a 3-month floating time. Numbers of students expected are as follows: 1 UVa (Xiaochao), 1 Temple (Zein-Eddine), 1 Todd (Junhao Chen already came to JLab. he previously worked on polarized 3He at LanZhou U.), and 1-2 from China. Students need to start taking shifts (to get training) in 2018.
  9. We need a responsiblity chart for carrying out each task, also with a 3-month floating time.
  10. Beamline items that we can think of now: modification for the 3He target, polarization requirement, beam ramping and rastering requirement, etc. JP and Dave will look into this (starting from 6 GeV documents) and then talk to Jay.
  11. Polarized 3He target status (JP):
    1. need to install optical fiber path from Hall A counting house to Hall C counting house, then into Hall C (JP, Joe)
    2. oven: bought, in target lab already
    3. cell status: protovec #1 and #2 were fully tested for the new pNMR, convection, etc. Three more cells were made in 2016 but two broke right away. The 3rd one is still at UVa waiting for density measurement (UVa laser broke that caused delay);
    4. Funding for a total of 5 cells was requested. These cells can be fully tested at JLab (including density measurements);
    5. Both Todd and Gordon promised new kappa-0 measurements for the EPR calibration for hybrid cells at high temperatures. Existing (very old) data are only for up to 200 C with Rb-only cell. But both got into difficulties.
    6. Magnets: Will use existing ones.
    7. Effect of SHMS bender magnet on the target: need Tosca calculation from Steve Laster. Then we can design the compensation coil setting. Yi did this or transversity (BigBite field) and Alexandre D. did this for SAGDH (Septum).
    8. Lasers: Jianping said no need to worry;
    9. Installation in Hall C: besides laser paths, also need target platform, ref. cell's gas system, and target ladder design.
    To do items:
  • Brad and Xiaochao will talk to THia or Steve next week to get the process started;
  • Brad and Xiaochao will make a beam time table to start the process on radiation dose calculation.